### v1

# Next-Generation Database Interfaces: A Survey of LLM-based Text-to-SQL

Abstract
Generating accurate SQL according to natural language questions (text-to-SQL) is a long-standing problem since it is challenging in user question understanding, database schema comprehension, and SQL generation. Conventional text-to-SQL systems include human engineering and deep neural networks. Subsequently, pre-trained language models (PLMs) have been developed and utilized for text-to-SQL tasks, achieving promising performance. As modern databases become more complex and corresponding user questions more challenging, PLMs with limited comprehension capabilities can lead to incorrect SQL generation. This necessitates more sophisticated and tailored optimization methods, which, in turn, restricts the applications of PLM-based systems. Most recently, large language models (LLMs) have demonstrated significant abilities in natural language understanding as the model scale remains increasing. Therefore, integrating the LLM-based implementation can bring unique opportunities, challenges, and solutions to text-to-SQL research. In this survey, we present a comprehensive review of LLM-based text-to-SQL. Specifically, we propose a brief overview of the current challenges and the evolutionary process of text-to-SQL. Then, we provide a detailed introduction to the datasets and metrics designed to evaluate text-to-SQL systems. After that, we present a systematic analysis of recent advances in LLM-based text-to-SQL. Finally, we discuss the remaining challenges in this field and propose expectations for future directions.
I Introduction
Figure 1: An example for LLM-based text-to-SQL selected from Spider. The user proposes a question, “What cartoons were written by Joseph Kuhr?” the LLM takes the question and the schema of its corresponding database as the input, then generates a SQL query as the output. The SQL query can be executed in the database and retrieve a content “Batman Series” to answer the user question.
Text-to-SQL is a long-standing task in natural language processing research. It aims to convert (translate) natural language questions into database-executable SQL queries. Fig. 1 provides an example of the large language model-based (LLM-based) text-to-SQL; given a user question “What cartoons were written by Joseph Kuhr?”, the LLMs take the question and its corresponding database schema as the input and generate an SQL query as the output, which can be executed in the database to retrieve the content “Batman Series” for answering the user question. The above system builds a natural language interface to the database (NLIDB) with LLMs. Since SQL remains one of the most widely used programming languages, with over half (51.52%) of professional developers using SQL in their work, but only around a third (35.29%) of those systematically trained [^1] https://survey.stackoverflow.co/2023, the NLIDB enables non-skilled users to access structured databases like professional database engineers [1, 2] and also accelerates human-computer interaction [3]. Furthermore, amid the research hotspot of LLMs, text-to-SQL can provide a solution to the prevalent hallucination [4, 5] issue by incorporating realistic content from the database to fill the knowledge gaps of LLMs [6]. The significant value and potential for text-to-SQL have triggered a range of studies on its incorporation and optimization with LLMs [7, 8, 9]; consequently, LLM-based text-to-SQL remains a highly discussed research field within the NLP and database communities.
Previous studies have made notable progress in the implementation of text-to-SQL and have undergone a long evolutionary process. Early efforts were mostly based on well-designed rules and templates [10, 11], specifically suitable for simple database scenarios. In recent years, with the heavy labor costs [12] brought by rule-based methods and the growing complexity of database environments [13, 14, 15], designing a rule or template for each scenario has become increasingly difficult and impractical. The development of deep neural networks has advanced the progress of text-to-SQL [16, 17], which can automatically learn a mapping from the user question to its corresponding SQL [18, 19]. Subsequently, pre-trained language models (PLMs) with strong semantic parsing capacity have become the new paradigm for text-to-SQL [20], taking its performance to a new level [21, 22, 23]. Incrementally, PLM-based research focused on table content encoding [24, 19, 25] and pre-training [20, 11] has further advanced this field. Recently, the LLM-based approaches implementing text-to-SQL through in-context learning (ICL) [7] and supervised fine-tuning (SFT) [9] paradigm, reaching state-of-the-art accuracy with the well-designed framework and stronger comprehension capability compared to PLMs.
The overall implementation details of LLM-based text-to-SQL can be divided into 3 aspects: 1. Question understanding: The NL question is a semantic representation of the user’s intention, which the corresponding generated SQL query is expected to align with; 2. Schema comprehension: The schema provides the table and column structure of the database, and the text-to-SQL system is required to identify the target components in the database that match the user question; 3. SQL generation: This involves incorporating the above parsing and then writing correct syntax to generate executable SQL queries that can retrieve the desired answer. The LLMs have proven to perform a good vanilla implementation [26, 27], benefiting from the more powerful semantic parsing enabled by the richer training corpus compared to the PLMs [28, 29]. Further studies on enhancing the LLMs for question understanding [7, 8], schema comprehension [30, 31], and SQL generation [32] are being increasingly released.
Figure 2: A sketch of the evolutionary process for text-to-SQL research from the perspective of implementation paradigm. Each stage is presented with some implementation techniques and representative works. The timestamps for the stages are not exactly accurate; we set each timestamp according to the release time of the representative works of each paradigm, with a margin of error of about one year before and after. The format is inspired from [29]
[…]
LLM-based Text-to-SQL
Datasets III-A) Original Datasets & Post-annotated Datasets Cross- domain BIRD [33], DuSQL [34], CoSQL [35], Spider [13], WikiSQL [14], KaggleDBQA [36], ADVETA [37] Spider-$S$ [38], Spider-CG [38], Spider-DK [39], Spider-SYN [40], Spider-Realistic [41], CSpider [42] SParC [43] Knowledge- augmented BIRD [33], SQUALL [44], Spider-DK [39] Cross- lingual DuSQL [34], CSpider [42] Context- dependent CoSQL [35], Spider-$S$ [38], Spider-CG [38], SparC [43] Robustness ADVETA [37], Spider-SYN [40], Spider-Realistic [41] Evaluation Metrics III-B) Content- Matching based Component Matching (CM) [13], Exact Matching (EM) [13] Execution based Execution Accuracy (EX) [13], Valid Efficiency Score (VES) [33] Methods IV) In-context Learning Paradigm IV-A) Zero-shot [26], [33], [27], [45], [46], [47], [48], [49], [8], [50] Few-shot [33], [7], [51], [52], [53], [54], [49], [8], [32], [55] Decomposition Coder-Reviewer [56], DIN-SQL [7], QDecomp [51], C3 [30], MAC-SQL [57], DEA-SQL [58], SGU-SQL [32], MetaSQL [59], PET-SQL [60], PURPLE [61] Prompt Optimization DESEM+P [62], StructGPT [63], SD+SA+Voting [52]RAG+SP&DRC [64], C3 [30], DAIL-SQL [8] ODIS [54], ACT-SQL [49], FUSED [65], DELLM [31] Reasoning Enhancement COT [51, 33, 8, 32], QDecomp [51], Least-to-Most [51], SQL-PaLM [53], ACT-SQL [49], POT [55] SQL-CRAFT [55], FUXI [66] Execution Refinement MBR-Exec [67], Coder-Reviewer [56], LEVER [68], SELF-DEBUGGING [48], DESEM+P [62] DIN-SQL [7], SD+SA+Voting [52], SQL-PaLM [53], RAG+SP&DRC [64], C3 [30], MAC-SQL [57] DELLM [31], SQL-CRAFT [55], FUXI [66], PET-SQL [60], PURPLE [61] Fine-tuning Paradigm IV-B) Supervised Fine-tuning [45], [8], [50], [53] Enhanced Architecture CLLMs [69] Pre-training CodeS [9] Data Augmentation DAIL-SQL [8], Symbol-LLM [50], CodeS [9], StructLM [70] Decomposition DTS-SQL [71]
Figure 3: Taxonomy tree of the research in LLM-based text-to-SQL. The display order in each node is organized by the released time. The format is adapted from [72]
I-A Challenges in Text-to-SQL
Despite the significant progress made in text-to-SQL research, several challenges remain that hinder the development of robust and generalized text-to-SQL systems [73]. In this survey, we aim to catch up with the recent advances and provide a comprehensive review of the current state-of-the-art (SOTA) models and approaches in LLM-based text-to-SQL. We begin by introducing the fundamental concepts and challenges associated with text-to-SQL, highlighting the importance of this task in various domains. We then delve into the evolution of LLMs and their application to text-to-SQL, discussing the key advancements and breakthroughs in this field. After the overview, we provide a detailed introduction to the recent advances of text-to-SQL incorporating LLMs. Specifically, the body of our survey covers a range of contents related to LLM-based text-to-SQL, including:
  • Datasets and Benchmarks: We provide an overview of the commonly used datasets and benchmarks for evaluating LLM-based Text-to-SQL systems. We discuss their characteristics, complexity, and the challenges they pose for text-to-SQL system development and evaluation.
  • Evaluation Metrics: We present the evaluation metrics used to assess the performance of LLM-based Text-to-SQL systems, including accuracy, exactness, and execution correctness. We discuss the advantages and limitations of each metric and their relevance to real-world applications.
  • Methods and Models: We explore the different methods and models employed for LLM-based text-to-SQL, including in-context learning and fine-tuning-based paradigms. We discuss their implementation details, strengths, and adaptations specific to the text-to-SQL task.
  • Expectations and Future Directions: We discuss the current challenges and limitations of LLM-based Text-to-SQL, such as real-world robustness, computational efficiency, data privacy and extensions. We also outline potential future research directions and opportunities for improvement.
We hope this survey will provide a clear overview of recent studies and inspire future studies. A taxonomy tree is shown in Fig. 3
II Overview
Text-to-SQL is a task that aims to convert natural language-type questions into corresponding SQL queries that can be executed on a relational database. Formally, given a user question $\mathcal{Q}$ (also known as a user query, natural language, NL, NL question, etc.) and a database schema $\mathcal{D}$, the goal is to generate an SQL query $Y$ that can be executed on the database to obtain the desired answer. Text-to-SQL has the potential to democratize access to data by allowing users to interact with databases using natural language without the need for specialized knowledge of SQL. This can benefit various domains, such as business intelligence, customer support, and scientific research, by enabling non-technical users to easily retrieve information from databases and facilitating more efficient data analysis.
II-1 Linguistic Complexity and Ambiguity
Natural language questions often contain complex linguistic structures, such as nested clauses, coreferences, and ellipses, which make it challenging to map them accurately to SQL queries. Additionally, natural language is inherently ambiguous, with multiple possible interpretations for a given question. Resolving these ambiguities and understanding the intent behind the question requires deep language understanding and the ability to incorporate context and domain knowledge.
II-2 Schema Understanding and Representation
To generate accurate SQL queries, text-to-SQL systems need to have a comprehensive understanding of the database schema, including table names, column names, and relationships between tables. However, database schemas can be complex and vary significantly across different domains. Representing and encoding the schema information in a way that can be effectively utilized by the text-to-SQL model is a challenging task.
II-3 Rare and Complex SQL Operations
Some SQL queries involve rare or complex operations, such as nested subqueries, outer joins, and window functions. These operations are less frequent in the training data and pose challenges for text-to-SQL models to generate accurately. Designing models that can handle a wide range of SQL operations, including rare and complex ones, is an important consideration.
II-4 Cross-Domain Generalization
Text-to-SQL models often struggle to generalize across different database schemas and domains. Models trained on a specific domain may not perform well on questions from a different domain due to differences in vocabulary, schema structure, and question patterns. Developing models that can effectively adapt to new domains with minimal fine-tuning or domain-specific training data is an ongoing challenge.
II-A Evolutionary Process
The field of text-to-SQL has witnessed significant advancements over the years, evolving from rule-based approaches to deep learning-based methods and, more recently, to the integration of pre-trained language models (PLMs) and large language models (LLMs), as shown in Fig. 2.
II-A1 Rule-based Approaches
Early text-to-SQL systems relied heavily on rule-based approaches [10, 12, 11], where manually crafted rules and heuristics were used to map natural language questions to SQL queries. These approaches often involved extensive feature engineering and domain-specific knowledge. While rule-based methods achieved success in specific domains, they lacked the flexibility and generalization capabilities needed to handle diverse and complex questions.
TABLE I: The statistics and analysis of well-known datasets of text-to-SQL ordered by release time. The original dataset indicates that the dataset is designed with a corresponding database, while post-annotated datasets involve annotating new components within existing datasets and databases rather than releasing a new database.
Original Dataset Release Time #Example #DB #Table/DB #Row/DB Characteristics
BIRD [33] May-2023 12,751 95 7.3 549K Cross-domain, Knowledge-augmented
KaggleDBQA [36] Jun-2021 272 8 2.3 280K Cross-domain
DuSQL [34] Nov-2020 23,797 200 4.1 - Cross-domain, Cross-lingual
SQUALL [44] Oct-2020 11,468 1,679 1 - Knowledge-augmented
CoSQL [35] Sep-2019 15,598 200 - - Cross-domain, Context-dependent
Spider [13] Sep-2018 10,181 200 5.1 2K Cross-domain
WikiSQL [14] Aug-2017 80,654 26,521 1 17 Cross-domain
Post-annotated Dataset Release Time Base Dataset Special Setting Characteristics
ADVETA [37] Dec-2022 Spider, etc. Adversarial table perturbation Robustness
Spider-SS&CG [38] May-2022 Spider Splitting example into sub-examples Context-dependent
Spider-DK [39] Sep-2021 Spider Adding domain knowledge Knowledge-augmented
Spider-SYN [40] Jun-2021 Spider Manual synonym replacement Robustness
Spider-Realistic [41] Oct-2020 Spider Removing column names in question Robustness
CSpider [42] Sep-2019 Spider Chinese version of Spider Cross-lingual
SParC [43] Jun-2019 Spider Annotate conversational contents Context-dependent
II-A2 Deep Learning-based Methods
With the rise of deep learning, sequence-to-sequence models, such as LSTMs and transformers, were adapted to generate SQL queries from natural language input [74, 19]. Typically, RYANSQL [19] introduced techniques like intermediate representations and sketch-based slot filling to handle complex questions and improve cross-domain generalization. Recently, researchers introduced graph neural networks (GNNs) for text-to-SQL tasks by leveraging schema dependency graphs to capture the relationships between database elements [18, 75].
II-A3 PLM-based Approaches
Pre-trained language models (PLMs) have emerged as a powerful solution for text-to-SQL, leveraging the vast amounts of linguistic knowledge and understanding captured during pre-training. The early adoption of PLMs in text-to-SQL primarily focused on fine-tuning pre-existing PLMs, such as BERT [24] and RoBERTa [76], on text-to-SQL datasets. These PLMs, which were pre-trained on large amounts of text data, captured rich semantic representations and language understanding capabilities. By fine-tuning them on text-to-SQL tasks, researchers aimed to leverage the knowledge and linguistic understanding of PLMs to generate accurate SQL queries [74, 20]. Another line of research focuses on incorporating schema information into PLMs to improve their understanding of database structures and generate more accurate SQL queries. Schema-aware PLMs are designed to capture the relationships and constraints present in the database schema [21].
II-A4 LLM-based Implementation
Large language models (LLMs), such as GPT series [77, 78], have gained significant attention in recent years due to their ability to generate coherent and fluent text. Researchers have started exploring the potential of LLMs for text-to-SQL by leveraging their vast knowledge and generative capabilities [26, 8]. These approaches often involve prompt engineering to guide the proprietary LLMs in the generation process or fine-tuning the open-source LLMs on text-to-SQL datasets.
The integration of LLMs in text-to-SQL is still an emerging area, and there is significant potential for further exploration and improvement. Researchers are investigating ways to better leverage the knowledge and reasoning capabilities of LLMs, incorporate domain-specific knowledge, and develop more efficient fine-tuning strategies. As the field continues to evolve, we can expect to see more advanced and powerful LLM-based approaches that push the boundaries of text-to-SQL performance and generalization.
III Benchmarks & evalution
In this section, we introduce the benchmarks for text-to-SQL, encompassing well-known datasets and evaluation metrics.
III-A Datasets
As shown in Table I, we classify the datasets into ”Original Datasets” and ”Post annotated Datasets” based on whether they were released with the original dataset and databases or created by adapting existing datasets and databases with special processing. We also highlight their release times. For the original datasets, we provide a detailed analysis, including the number of examples, databases, tables per database, and rows per database. For the post-annotated datasets, we identify their base dataset and describe the special processing applied to them. To illustrate the potential opportunities of each dataset, we categorize their characteristics into Cross-domain, Knowledge-augmented, Context-dependent, Robustness, and Cross-lingual, which we will discuss in detail below.
III-A1 Cross-domain Dataset.
This refers to datasets where the background information of different databases comes from various domains. Since real-world text-to-SQL tasks often involve databases from multiple domains, most original text-to-SQL datasets [14, 13, 35, 34, 36, 33] and post annotated datasets [43, 42, 39, 40, 41, 37, 38] are in the cross-domain setting to fit well with the needs of cross-domain applications.
III-A2 Knowledge-augmented Dataset
Interest in incorporating knowledge into text-to-SQL tasks has increased significantly in recent years. BIRD [33] employs domain experts to annotate each text-to-SQL sample with external knowledge, categorized into Numeric Reasoning Knowledge, Domain Knowledge, Synonym Knowledge, and Value Illustration. Similarly, Spider-DK [39] defines and adds five types of domain knowledge for a human-curated version of the Spider dataset: SELECT Columns Mentioned by Omission, Simple Inference Required, Synonyms Substitution in Cell Value Word, One Non-Cell Value Word Generate a Condition, and Easy to Conflict with Other Domains. Both studies found that human-annotated knowledge significantly improves SQL generation accuracy for samples requiring external domain knowledge. Additionally, SQUALL [44] manually annotates alignments between the words in NL questions and the entities in SQL, providing finer-grained supervision than other datasets.
III-A3 Context-dependent Dataset
SParC [43] and CoSQL [35] explore context-dependent SQL generation by constructing a conversational database querying system. Unlike traditional text-to-SQL datasets that only have a single (question, SQL) pair for one example, SParC decomposes the (question, SQL) examples in the Spider dataset into multiple (sub-question, SQL) pairs to construct a simulated and meaningful interaction, including inter-related sub-questions that aid SQL generation, and unrelated sub-questions that enhance data diversity. CoSQL, in comparison, involves conversational interactions in natural language, simulating real-life scenarios to increase complexity and diversity. Additionally, Spider-SS&CG [38] splits the NL question in the Spider dataset into multiple sub-questions and sub-SQLs, demonstrating that training on these sub-examples can improve a text-to-SQL model’s generalization ability on out-of-distribution samples.
III-A4 Robustness Dataset
Evaluating the accuracy of text-to-SQL models with attacked or perturbed database contents (e.g., schema and tables) is crucial for assessing robustness. Spider-Realistic [41] removes explicit schema-related words from the NL questions, while Spider-SYN [40] replaces them with manually selected synonyms. ADVETA [37] introduces Adversarial Table Perturbation (ATP), which perturbs tables by replacing original column names with misleading alternatives and inserting new columns with high semantic associations but low semantic equivalency. These perturbations lead to significant drops in accuracy, as a text-to-SQL model with low robustness may be misled by incorrect matches between tokens in NL questions and database entities.
III-A5 Cross-lingual Dataset
SQL keywords, function names, table names, and column names are typically written in English, posing challenges for applications in other languages. CSpider [42] translates the Spider dataset into Chinese, identifying new challenges in word segmentation and cross-lingual matching between Chinese questions and English database contents. DuSQL [34] introduces a practical text-to-SQL dataset with Chinese questions and database contents provided in both English and Chinese.
III-B Evaluation Metrics
We introduce four widely used evaluation metrics for the text-to-SQL task as follows: Component Matching and Exact Matching, which are based on SQL content matching, and Execution Accuracy and Valid Efficiency Score, which are based on execution results.
III-B1 Content Matching-based Metrics
SQL content matching metrics focus on comparing the generated SQL query with the ground truth SQL query based on their structural and syntactic similarities.
  • Component Matching (CM) [13] evaluates the model performance by measuring the exact match between predicted and ground truth SQL components—SELECT, WHERE, GROUP BY, ORDER BY, and KEYWORDS—using the F1 score. Each component is decomposed into sets of sub-components and compared for an exact match, accounting for SQL components without order constraints.
  • Exact Matching (EM) [13] measures the percentage of examples whose predicted SQL query is equivalent to the ground truth SQL query. A predicted SQL is considered correct only if all its components, as described in Component Matching, match exactly with those of the ground truth query.
III-B2 Execution-based Metrics
Execution result metrics assess the correctness of the generated SQL query by comparing the results obtained from executing the query on the target database with the expected results.
  • Execution Accuracy (EX) [13] measures the correctness of a predicted SQL query by executing it in the corresponding database and comparing the executed results with the ground truth queries.
  • Valid Efficiency Score (VES) [33] is defined to measure the efficiency of valid SQL queries, which are the predicted SQL queries whose executed results exactly match the ground truth query. Thus, VES evaluates both the efficiency and accuracy of predicted SQL queries. For an evaluation dataset with $N$ examples, VES can be computed by:
    $\text{VES}=\frac{1}{N}\sum_{n=1}^{N}\mathbbm{1}(V_{n},\hat{V}_{n})\cdot\textbf{R}(Y_{n},\hat{Y}_{n}),$ (1)
    where $\hat{Y}_{n}$ and $\hat{V}_{n}$ are the predicted SQL query and its executed results. $Y_{n}$ and $V_{n}$ are those of the corresponding ground truth SQL query. $\mathbbm{1}(V_{n},\hat{V}_{n})$ is an indicator function, where:
    $\mathbbm{1}(V_{n},\hat{V}_{n})=\begin{cases}1,V_{n}=\hat{V}_{n}\\ 0,V_{n}\neq\hat{V}_{n}\end{cases}$ (2)
    Then, $\textbf{R}(Y_{n},\hat{Y}_{n})=\sqrt{E(Y_{n})/E(\hat{Y}_{n})}$ denotes the relative execution efficiency of the predicted SQL query in comparison to ground-truth query, where $E(\cdot)$ is the execution time of each SQL in the database. [33] ensures the stability of this metric by computing the average of $\textbf{R}(Y_{n},\hat{Y}_{n})$ over 100 runs for each example.
Most of the recent LLM-based text-to-SQL studies focus on these four datasets: Spider [13], Spider-Realistic [41], Spider-SYN [40], and BIRD [33]; and these three evaluation methods: EM, EX, and VES, we will primarily focus on them in the following introduction.
IV Methods
The implementation of LLM-based applications mostly relies on in-context learning (prompt engineering) [79] and fine-tuning [80] since the powerful proprietary models and well-architected open-source models are being released in large quantities [81, 82, 83]. The specific methods can generally be divided into two paradigms, in-context learning (ICL) and fine-tuning (FT). In this survey, we primarily focus on these text-to-SQL paradigms and will discuss them accordingly.
IV-A In-context Learning
Through extensive and widely recognized research, prompt engineering has been proven to play a decisive role in the performance of LLMs [28], also impacting the text-to-SQL task in different prompt styles [8, 46]. Necessarily, developing text-to-SQL methods in the in-context learning (ICL) paradigm is valuable for achieving promising improvement. The implementation of LLM-based text-to-SQL process to generate executable SQL query $Y$ is formulated as:
$Y=f(\mathcal{Q},\mathcal{S},\mathcal{I}\mid\theta),$ (3)
where $\mathcal{Q}$ represents the user question. $\mathcal{S}$ is the database schema/content, which can be decomposed as $\mathcal{S}=\langle\mathcal{C},\mathcal{T},\mathcal{K}\rangle$, where $\mathcal{C}=\{c_{1},c_{2},$...\} and tables $\mathcal{T}=\{t_{1},t_{2},$...\} represent the collection of various columns and tables, $\mathcal{K}$ is the potentially external knowledge (e.g. foreign key relationships [49], schema linking [30] and domain knowledge [33, 31]). $\mathcal{I}$ represents the instruction for the text-to-SQL task, which performs indicative guidance to guide the LLMs for generating an accurate SQL query. $f(\cdot\mid\theta)$ is a LLM with parameter $\theta$. In the in-context learning (ICL) paradigm, we utilize an off-the-shelf text-to-SQL model (i.e., parameter $\theta$ of the model is frozen) for implementation. Various well-designed methods have been adopted in the ICL paradigm for text-to-SQL. We group them into five categories $\mathbf{C}_{0:4}$, including: $\mathbf{C}_{0}$-Trivial Prompt, $\mathbf{C}_{1}$-Decomposition, $\mathbf{C}_{2}$-Prompt Optimization, $\mathbf{C}_{3}$-Reasoning Enhancement, and $\mathbf{C}_{4}$-Execution Refinement, the details are shown in Tab. II.
$\mathbf{C}_{0}$-Trivial Prompt
Trained through massive data, LLMs have a strong overall proficiency in different downstream tasks with zero-shot and few-shot prompting [84, 80, 85], which is widely recognized and achieves promising results. In our survey, we categorized the above prompting approaches without the well-designed framework as trivial prompts. As introduce above, the formulated process of LLM-based text-to-SQL Eq. 3 can also represent zero-shot prompting, the input $\mathcal{P}_{0}$ can be obtained by concatenate $\mathcal{I}$, $\mathcal{S}$ and $\mathcal{Q}$:
$\mathcal{P}_{0}=\mathcal{I}\oplus\mathcal{S}\oplus\mathcal{Q}.$ (4)
To regulate the prompting process, we set the OpenAI demonstration [^2] The prompt style that follows the official document from OpenAI platform: https://platform.openai.com/examples/default-sql-translate as the standard (trivial) prompt [30] for text-to-SQL.
TABLE II: Typical methods used for in-context learning (ICL) in LLM-based text-to-SQL. The full table of existing methods with categorization $\mathbf{C}_{1:4}$ and more details are listed in Table III.
Methods Adopted by Applied LLMs
$\mathbf{C}_{0}$-Trivial Prompt Zero-shot [26] ChatGPT
Few-shot [8] ChatGPT
$\mathbf{C}_{1}$-Decomposition DIN-SQL [7] GPT-4
$\mathbf{C}_{2}$-Prompt Optimization DAIL-SQL [8] GPT-4
$\mathbf{C}_{3}$-Reasoning Enhancement ACT-SQL [49] GPT-4
$\mathbf{C}_{4}$-Execution Refinement LEVER [68] Codex
TABLE III: Well-designed methods used in in-context learning (ICL) paradigm for LLM-based text-to-SQL ordered by release time. The methods are grouped in four categories based on their implementation perspective: $\mathbf{C}_{1}$-Decomposition, $\mathbf{C}_{2}$-Context Augmentation, $\mathbf{C}_{3}$-Reasoning Enhancement, $\mathbf{C}_{4}$-Execution Refinement. The method in multiple categories will be introduced respectively. *There are multiple applied LLMs in the corresponding method; we present the selected LLM with representative performance. COT method are reported in multiple venues: NeurIPS’23 [33], EMNLP’23 [51], VLDB’24 [8], arXiv’24 [32]
Methods Applied LLMs Benchmark Metrics $\mathbf{C}_{1}$ $\mathbf{C}_{2}$ $\mathbf{C}_{3}$ $\mathbf{C}_{4}$ Release Time Publication Venue
MBR-Exec [67] Codex [13] EX Apr-2022 EMNLP’22
Coder-Reviewer [56] Codex [13] EX Nov-2022 ICML’23
LEVER [68] Codex [13] EX Feb-2023 ICML’23
SELF-DEBUGGING [48] StarCoder* [13] EX Apr-2023 ICLR’24
DESEM+P [62] ChatGPT [13, 40] EX Apr-2023 PRICAI’23
DIN-SQL [7] GPT-4* [33, 13] EX, EM, VES Apr-2023 NeurIPS’23
COT [51, 33, 8, 32] GPT-4 [13, 41, 33] EX, VES May-2023 Multiple Venues
StructGPT [63] ChatGPT* [13, 40, 41] EX May-2023 EMNLP’23
SD+SA+Voting [52] ChatGPT* [13, 40, 41] EX May-2023 EMNLP’23 Findings
QDecomp [51] Codex [13, 41] EX May-2023 EMNLP’23
Least-to-Most [51] Codex [13] EX May-2023 EMNLP’23
SQL-PaLM [53] PaLM-2 [13] EX May-2023 arXiv’23
RAG+SP&DRC [64] ChatGPT [13] EX Jul-2023 ICONIP’23
C3 [30] ChatGPT [13] EX Jul-2023 arXiv’23
DAIL-SQL [8] GPT-4* [13, 33, 41] EX, EM, VES Aug-2023 VLDB’24
ODIS [54] Codex* [13] EX Oct-2023 EMNLP’23 Findings
ACT-SQL [49] GPT-4* [13, 40] EX, EM Oct-2023 EMNLP’23 Findings
MAC-SQL [57] GPT-4* [13, 33] EX, EM, VES Dec-2023 arXiv’23
DEA-SQL [58] GPT-4 [13] EX Feb-2024 ACL’24 Findings
FUSED [65] ChatGPT* [13] EX Feb-2024 arXiv’24
DELLM [31] GPT-4* [13, 33] EX, VES Feb-2024 ACL’24 Findings
SGU-SQL [32] GPT-4* [13, 33] EX, EM Feb-2024 arXiv’24
POT [55] GPT-4* [13, 33] EX Feb-2024 arXiv’24
SQL-CRAFT [55] GPT-4* [13, 33] EX Feb-2024 arXiv’24
FUXI [66] GPT-4* [33] EX Feb-2024 arXiv’24
MetaSQL [59] GPT-4* [13] EX, EM Feb-2024 ICDE’24
PET-SQL [60] GPT-4 [13] EX Mar-2024 arXiv’24
PURPLE [61] GPT-4* [13, 40, 41] EX, EM Mar-2024 ICDE’24
Zero-shot: Many research works [26, 27, 46]l utilize zero-shot prompting, focusing mainly on the influence of the style of prompt construction and the zero-shot performance of various LLMs for text-to-SQL. As an empirical evaluation, [26] evaluates the baseline text-to-SQL capabilities of different early-developed LLMs [86, 77, 87], which also presents the results for different prompt styles. The results indicate that prompt design is critical for performance, with error analysis, [26] propose more database content can harm the overall performance. Since ChatGPT emerged with impressive capabilities in conversational scenarios and code generation, [27] assesses its performance of text-to-SQL. With zero-shot settings, the results demonstrate that ChatGPT has a promising text-to-SQL ability compared to the state-of-the-art (SOTA) models. For fair comparability, [47] reveal effective prompt construction for LLM-based text-to-SQL; they study different styles of prompt construction and make conclusions of zero-shot prompt design based on the comparisons.
Primary key and foreign key carry contiguous knowledge of different tables, [49] study their impact by adding these keys on different designed database prompt style on zero-shot prompting results. A benchmark evaluation [8] also studies the influence of foreign keys, with five different prompt representation styles, each style can be considered as the permutation and combination of the instruction, rule implication, and foreign key. Apart from the foreign key, this study also explores zero-shot prompting combined with “no explanation” rule and the rule implication “Let’s think step by step”. Empowered by the annotated knowledge of human experts, [33] follow the standard prompting and obtain improvement by combining the provided oracle knowledge. With the explosion of open-source LLMs, according to the results of similar evaluation, these models are also capable of zero-shot text-to-SQL task [45, 50, 46], especially code generation models [48, 46]. For zero-shot prompting optimization, [46] raises a challenge for designing an effective prompt template for LLMs; the former prompt construction lacks structure uniformity, which makes it hard to find out a concrete element within a prompt constructing template influences the performance of LLMs. They address this challenge by investigating a more unified series of prompt templates warping with different prefixes, infixes, and postfixes.
Few-shot: The technique of few-shot prompting is widely used in both practical applications and well-designed research, which has been proven efficient for eliciting better performance of LLMs [28, 88]. The input prompt of the few-shot approach LLM-based text-to-SQL can be formulated as an extension of Eq. 3:
$\mathcal{P}_{n}=\{\mathcal{F}_{1},\mathcal{F}_{2},\ldots,\mathcal{F}_{n}\}\oplus\mathcal{P}_{0},$ (5)
where the $\mathcal{P}_{n}$ represent the input prompt for $n$-shot learning, $n$ is the provided instances (examples) number; $\mathcal{F}$ denote the few-shot instance, which can be decomposed as $\mathcal{F}_{i}=(\mathcal{S}_{i},\mathcal{Q}_{i},Y_{i})$, $i$ is the serial number of instances. The study of few-shot prompting focuses on the number of representations and few-shot instance selection.
As pilot experiments, few-shot prompting for text-to-SQL are evaluated in multiple datasets with various LLMs [7, 32], achieve better performance rather than zero-shot approaches. [33] provides a 1-shot example combining with CoT [89] prompting, the decomposed reasoning steps of the given example trigger the text-to-SQL model for generating accurate SQL. [55] study the effect of the number of few-shot examples. [52] focus on the sampling strategies by studying the similarity and the diversity between different demonstrations, setting random sampling as the baseline, and evaluating different strategies and their combination for comparison. Furthermore, above the similarity selection, [8] evaluated masked question similarity selection and the upper limit of similarity approaches with various numbers of few-shot examples. A study of difficult-level samples selection [51] compared the performance of few-shot Codex, with random selection and difficulty-based selection for few-shot instances on difficulty categorized dataset [13, 41]. Three difficulty-based selection strategies are devised based on the number of selected samples at different difficulty levels. [49] utilize a hybrid strategy for selecting samples, which combines static examples and similarity-based dynamic examples for few-shot prompting. In their evaluations, they also test the impact of different input schema styles and different static and dynamic exemplar numbers.
The impact of cross-domain few-shot examples is also being studied [54]. When incorporating in-domain and out-of-domain with different numbers of examples, the in-domain demonstration always outperforms zero-shot and out-of-domain examples and gets better as the number of examples rises. To explore the detailed construction of input prompt, [53] compare the concise and verbose prompt design approaches. The former style splits the schema, the column names, and the primary and foreign keys by bar, and the latter organizes them as natural language descriptions.
$\mathbf{C}_{1}$-Decomposition
As an intuitive solution, decomposing a challenging user question into simpler sub-questions and using multi-step reasoning for implementation can reduce the complexity of the full text-to-SQL task [51, 7]. Dealing with less complexity, LLMs have the potential to have better performance. The decomposition approaches for LLM-based text-to-SQL are categorized into two paradigms: (1) sub-task decomposing, provides additional parsing to assist the final SQL generation by decomposing the overall text-to-SQL task into smaller effective sub-tasks (e.g., schema linking, domain classification). (2) sub-question decomposing, divides the user question into sub-questions to reduce the question’s complexity and difficulty, then generates the sub-SQL by solving these questions to deduce the final SQL query.
DIN-SQL [7] proposed a decomposed in-context learning method consisting of four modules: schema linking, classification & decomposition, SQL generation, and self-correction. DIN-SQL first finishes the schema linking between the user question and the target database; the following module decomposes the user question into correlated sub-questions and does a difficulty classification. Based on the above information, the SQL generation module generates a corresponding SQL, and the self-correction module identifies and corrects the potential errors in the predicted SQL. This approach comprehensively considers the decomposition of both sub-tasks and sub-questions. Coder-Reviewer [56] framework proposed a re-ranking method, combining Coder models for the generation and Reviewer models to evaluate the likelihood of the instruction. Refer to the Chain-of-Thought [89] and Least-to-Most prompting [90], QDecomp [51] introduce question decomposition prompting, which follows the question reduction stage in least-to-most prompting and instruct the LLM to decompose the original complex question as the intermediate reasoning steps. C3 [30] consists of three key components: clear prompting, calibration bias prompting, and consistency; these components are accomplished by assigning ChatGPT with different tasks. Firstly, the clear prompting component generates the schema linking and the distilled question-relevant schema as a clear prompt. Then, a multi-turn dialogue about text-to-SQL hints is utilized as a calibration bias prompt, which combines with the clear prompt to guide the SQL generation. The generated SQL queries are selected by consistency and execution-based voting to get the final SQL. MAC-SQL [57] presents a multi-agent collaborating framework; the text-to-SQL process is finished as the collaboration of the agents: Selector, Decomposer, and Refiner. The Selector preserves relevant tables for user questions; the Decomposer breaks down user questions into sub-questions and provides solutions; finally, the Refiner validates and refines the defective SQL. DEA-SQL [58] introduces a workflow paradigm aiming to enhance the attention and problem-solving scope of LLM-based text-to-SQL through decomposition. This method decomposes the overall task, enabling the SQL generation module to have the corresponding prerequisite (information determination, question classification) and subsequent (self-correction, active learning) sub-tasks. Through the workflow paradigm, an accurate SQL query is generated. SGU-SQL [32] is a structure-to-SQL framework, leveraging the inherent structure information to assist SQL generation. Specifically, the framework constructs a graph structure for the user question and the corresponding database respectively, then uses the encoded graphs to construct structure linking [91, 92]. A meta operator decomposes the user question with a grammar tree and finally designs the input prompt with meta-operation in SQL. MetaSQL [59] introduces a three-stage approach for SQL generation, which consists of decomposition, generation, and rank. The decomposition stage uses semantic decomposition and metadata composition to process the user question. Taking the previously processed data as input, a text-to-SQL model using metadata-conditioned generation to generate some candidate SQL queries. Finally, a two-stage ranking pipeline is applied to get a global-optimal SQL query. PET-SQL [60] proposed a prompt-enhanced two-stage framework. Firstly, an elaborated prompt instructs the LLMs to generate preliminary SQL (PreSQL) where some few-shot demonstrations are selected based on similarity. Then, schema linking is found based on PreSQL and combined to prompt the LLMs to generate the Final SQL (FinSQL). Finally, multiple LLMs are utilized to generate a FinSQL, ensuring consistency based on the execution results.
$\mathbf{C}_{2}$-Prompt Optimization
As previously introduced, few-shot learning is widely studied for prompting LLMs [77]. For LLM-based text-to-SQL with in-context learning, trivial few-shot approaches obtained promising results [33, 7, 8], further optimization of few-shot prompting has the potential to achieve higher performance. Since the performance of SQL generation in off-the-shelf LLMs largely depends on the quality of the corresponding input prompt [93], many decisive factors that can influence the quality of the prompt have become focuses of the research [8] (e.g., quality and quantity in the few-shot organization, the similarity between user questions and few-shots instances, external knowledge/hints). The process of prompt quality improvement is actually the prompt’s optimization, including few-shot sampling strategies, schema augmentation, and external knowledge generation.
DESEM [62] is a prompt engineering framework with de-semanticization and skeleton retrieval. The framework first employs domain-specific words masking module to remove the semantic tokens in questions that preserve the question’s intentions. And then utilizes an adjustable prompting module that retrieves the few-shot examples with identical question intentions and incorporates schema-relevance filtering to guide the LLM’s SQL generation. The QDecomp [51] framework introduces the InterCOL mechanism to incrementally incorporate the decomposed sub-questions with correlative table and column names. With difficulty-based selection, the few-shot examples for QDecomp are difficult-level sampled. Besides similarity-diversity sampling, [52] proposed SD+SA+Voting (Similarity-Diversity+Schema Augmentation+Voting) sampling strategy. They first employ semantic similarity and $k$-Means cluster diversity for sampling few-shot examples and then enhance the prompt with schema knowledge (semantic or structure augmentation). C3 [30] framework comprises a clear prompting component, which takes the question and schema as the LLMs input, generates a clear prompt that includes a schema that removes the redundant information irrelevant to the user question and a schema linking, and also a calibration component providing hints. The LLMs take their composition as context-augmented prompts for SQL generation. A retrieval-augmented framework is introduced with sample-aware prompting [64], which simplifies the original question and extracts the question skeleton from the simplified question, then finishes the sample retrieval in the repository according to skeleton similarities. The retrieved samples are combined with the original question for few-shot prompting. ODIS [54] introduces the selection of a sample with out-of-domain demonstrations and in-domain synthetic data, which retrieves few-shot demonstrations from hybrid sources to augment the prompt representations. DAIL-SQL [8] proposed a novel approach to address the issues in few-shot sampling and organization, presenting a better balance between the quality and quantity of few-shot examples. DAIL Selection first masks domain-specific words in user and few-shot example questions, then ranks the candidate examples based on the embedded Euclidean distance. Meanwhile, the similarity between the pre-predicted SQL queries is calculated. Finally, the selection mechanism obtains the similarity-sorted candidates according to the pre-set criteria. The few-shot examples are guaranteed good similarity with both questions and SQL queries with this method. ACT-SQL [49] proposed dynamic examples in few-shot prompting, which is selected according to similarity score. FUSED [65] are presented to build a high-diversity demonstrations pool through human-free multiple-iteration synthesis to improve the diversity of the few-shot demonstrations. The pipeline of FUSED samples the demonstrations to be fused by clustering, then fuse the sampled demonstrations to construct the pool to enhance few-shot learning. Knowledge-to-SQL [31] framework aims to build a Data Expert LLM (DELLM) to provide knowledge for SQL generation. The DELLM is trained by supervised fine-tuning using human expert annotations [33] and further refined by preference learning with the database’s feedback. DELLM generates four categories of knowledge, the well-designed methods (e.g. DAIL-SQL [8], MAC-SQL [57]) incorporating the generated knowledge to achieve better performance for LLM-based text-to-SQL with in-context learning.
$\mathbf{C}_{3}$-Reasoning Enhancement
LLMs have exhibited promising capabilities in tasks involving commonsense reasoning, symbolic reasoning, and arithmetic reasoning [94]. Since for the text-to-SQL tasks, numeric and synonym reasoning frequently occur in realistic scenarios [33, 41], the prompting strategies for the LLMs reasoning possess the potential to enhance their SQL generation capabilities. Recent studies primarily focus on incorporating well-designed reasoning-enhanced methods for text-to-SQL adaptation, improving LLMs to address the challenge about complex questions that require multi-step reasoning and the issue of self-consistency [95] in SQL generation.
Chain-of-Thoughts (CoT) prompting technique [89] involves a comprehensive reasoning process that guides LLMs towards accurate deduction, eliciting reasoning in LLMs. The study of LLM-based text-to-SQL utilizes CoT prompting as rule implication [8], which setting the instruction “Let’s think step by step” in prompt construction [33, 51, 8, 32]. However, the straightforward (original) CoT strategy has not demonstrated the potential in text-to-SQL tasks that it has in other reasoning tasks; studying CoT for adaptations is still an ongoing research [51]. Since CoT prompting always uses static examples with human annotation for demonstrations, which requires empirical judgment for the effective selection of few-shot examples, and manual annotating is also an essential need. As a solution, ACT-SQL [49] proposed a method to generate CoT examples automatically. Specifically, given a question, ACT-SQL truncates a set of slices of the question and then enumerates every column appearing in the corresponding SQL query. Each column will be linked with its most relevant slice through the similarity function and appended to the CoT prompt. Through systematical study for enhancing LLMs SQL generation incorporating CoT prompting, QDecomp [51] presents a novel framework to address the challenge for CoT how to come up with the reasoning steps to predict the SQL query. The framework utilizes every slice of the SQL query to construct a logical step in CoT reasoning, then employs natural language templates to articulate each slice of the SQL query and arranges them in the logical execution order. Least-to-Most [90] is another prompting technique that decomposes questions into sub-questions and then sequentially solves them. As iterative prompting, pilot experiments [51] demonstrate that it may be unnecessary for text-to-SQL parsing. Using detailed reasoning steps tends to have more error propagation issues. As a variant of CoT, Program-of-Thoughts (PoT) prompting strategy [96] are proposed to enhance arithmetic reasoning for LLMs. Through evaluation [55], PoT enhances the LLM for SQL generation, especially in complicated datasets [33]. SQL-CRAFT [55] are proposed to enhance LLM-based SQL generation, which incorporates PoT prompting for Python-enhanced reasoning. PoT strategy requires the model to simultaneously generate the Python code and the SQL queries, enforcing the model to incorporate Python code in its reasoning process. Self-Consistency [95] is a prompting strategy improving reasoning in LLMs, which leverages the intuition that a complex reasoning problem typically admits multiple different ways of thinking, leading to its unique correct answer. In the text-to-SQL task, self-consistency is adapted to sampling a set of different SQL and voting for consistent SQL via execution feedback [30, 53]. Similarly, the SD+SA+Voting [52] framework eliminates those with execution errors identified by the deterministic database management system (DBMS) and opts for the prediction that garners the majority vote. Furthermore, motivated by recent research on extending the capabilities of LLMs with tools, FUXI [66] are proposed to enhance LLMs SQL generation through effectively invoking crafted tools.
$\mathbf{C}_{4}$-Execution Refinement
To design criteria for accurate SQL generation, whether a generated SQL can be successfully executed and elicit a correct answer for the user question is always the priority [13]. As a complex programming task, generating the correct SQL in one go becomes challenging. Intuitively, considering the execution feedback/results in SQL generation assists the alignment to the corresponding database environment, which allows the LLMs to gather the potential executed errors and results to refine the generated SQL or hold a majority vote [30]. The execution-aware methods in text-to-SQL incorporate the execution feedback in two main approaches: 1) Incorporating the feedback through second round prompting for regeneration, for every SQL query generated in the initial response, it will be executed in the corresponding database, thus obtaining feedback from the database. This feedback might be an error, or it might yield results that will be appended to the second round prompt. Through in-context learning of this feedback, LLMs are able to refine or regenerate the original SQL, thereby enhancing accuracy. 2) Utilize execution-based selection strategies for generated SQL, sample multiple generated SQL queries from LLM), and execute each in the database. Based on the results of each SQL query, use selection strategies (e.g., self-consistency, majority vote [60]) to define a query from the SQL set that satisfies the criteria as the final predicted SQL.
MRC-EXEC [67] introduced a natural language to code (NL2Code) translation framework with execution, which executes each sampled SQL query and selects the example with the minimal execution result–based Bayes risk [97]. LEVER [68] proposed an approach to verify NL2Code with execution, utilizing a generation and execution module to collect sampled SQL set and their execution results, respectively, then using a learned verifier to output the probability of the correctness. Similarly, the SELF-DEBUGGING [48] framework is presented to teach LLMs to debug their predicted SQL via few-shot demonstrations. The model is able to refine its mistakes by investigating the execution results and explaining the generated SQL in natural language without human interventions
As previously introduced, to incorporate the well-designed framework with execution feedback, two-stage implications are widely-used: 1. sampling a set of SQL queries. 2. majority vote (self-consistency). Specifically, the C3 [30] framework removes the errors and identifies the most consistent SQL; The retrieval-augmented framework [64] introduced a dynamic revision chain, combining fine-grained execution messages with database content to prompt the LLMs to convert the generated SQL query into natural language explanation; the LLMs are requested to identify the semantic gaps and revise their own generated SQL. Although schema-filtering approaches enhance SQL generation, the generated SQL could be unexecutable. DESEM [62] incorporates a fallback revision to address the issue; it revises and regenerates the SQL base on different kinds of errors and sets termination criteria to avoid the loop. DIN-SQL [7] designed a generic and gentle prompt in their self-correction module; the generic prompt requests the LLMs to identify and correct the error, and the gentle prompt asks the model to check the potential issue. The multi-agent framework MAC-SQL [57] comprises a refiner agent, which is able to detect and automatically rectify SQL errors, taking SQLite error and exception class to regenerate fixed SQL. Since different questions may require different numbers of revisions, SQL-CRAFT [55] framework introduced interactive correction with an automated control determination process to avoid over-correction or insufficient correction. FUXI [66] considers the error feedback in tool-based reasoning for SQL generation. The Knowledge-to-SQL [31] introduces a preference learning framework incorporating the database execution feedback with a direct preference optimization [98] for refining the proposed DELLM. PET-SQL [60] proposed cross consistency, which comprises two variants: 1) naive voting: instruct multiple LLMs to generate the SQL query, then utilizing the majority vote for the final SQL base on different execution results; 2) fine-grained voting: refine the naive voting based on the difficulty level to mitigate the voting bias.
IV-B Fine-tuning
TABLE IV: Well-designed methods used in fine-tuning (FT) for LLM-based text-to-SQL. The methods in each category are ordered by release time. *The methods are utilized in multiple open-source LLMs; we select a representative model to present.
Category Adopted by Applied LLMs Dataset EX EM VES Release Time Publication Venue
Enhanced Architecture CLLMs [69] Deepseek* [13] Mar-2024 ICML’24
Pre-training CodeS [9] StarCoder [13, 33] Feb-2024 SIGMOD’24
Data Augmentation DAIL-SQL [71] LLaMA* [13, 41] Aug-2023 VLDB’24
Symbol-LLM [50] CodeLLaMA [13] Nov-2023 ACL’24
CodeS [9] StarCoder [13, 33] Feb-2024 SIGMOD’24
StructLM [70] CodeLLaMA [13] Feb-2024 arXiv’24
Decomposition DTS-SQL [71] Mistral* [13, 40] Feb-2024 arXiv’24
Since supervised fine-tuning (SFT) is the mainstream approach in the LLMs training [29], for open-source LLMs (e.g., LLaMA-2 [82], Gemma [99]), the most straightforward method to enable the model to adapt a specific domain quickly is to use collected domain label to perform SFT on the model. The SFT phase is typically the preliminary phase of the well-designed training framework [98, 100], as well as the fine-tuning of text-to-SQL. The auto-regressive generation process of SQL query $Y$ can be formulated as follows:
$P_{\pi}(Y\mid\mathcal{P})=\prod_{k=1}^{n}P_{\pi}(y_{k}\mid\mathcal{P},Y_{1:k-1}),$ (6)
where $Y=\{y_{1},y_{2},\ldots,y_{n}\}$ is an SQL query of length $n$, $y_{k}$ is the corresponding $k^{th}$ token of the SQL query, $Y_{1:k-1}$ is the prefix sequence of $Y$ ahead the token $y_{k}$. $P_{\pi}(y_{k}\mid\cdot)$ is a conditional probability of a LLM $\pi$ for generating the $k^{th}$ token of $Y$ base on the input prompt $\mathcal{P}$ and the prefix sequence.
Given a basic open-source model $\pi^{0}$, the goal of SFT is obtain a model $\pi^{SFT}$ through minimizing the cross-entropy loss:
$\mathcal{L}_{SFT}=-\sum_{k=1}^{n}\operatorname{log}P_{\pi^{0}}(\hat{y}_{k}=y_{k}\mid\mathcal{P},Y_{1:k-1}),$ (7)
where $\hat{y}_{k}$ is the $k^{th}$ token of the generated SQL query $\hat{Y}$, and $Y$ is the corresponding ground-truth label.
The SFT approach for text-to-SQL has been widely adopted in text-to-SQL research for various open-source LLMs [9, 46, 8]. Compared to in-context learning (ICL) approaches, fine-tuning paradigms are more inclined to be at a starting point in LLM-based text-to-SQL. Currently, several studies exploring a better fine-tuning method have been released. We categorize the well-designed fine-tuning methods in different groups based on their mechanisms, as shown in Tab. IV.
Enhanced Architecture
The widely-used generative pre-trained transformer (GPT) framework utilizes decoder-only transformer architecture and conventional auto-regressive decoding for text generation. Recent studies on the efficiency of LLMs have revealed a common challenge: when generating long sequences with the auto-regressive paradigm, the need to incorporate the attention mechanism results in high latency for LLMs [101, 102]. In LLM-based text-to-SQL, the speed of generating SQL queries is significantly slower compared to traditional language modeling [21, 28], which has become a challenge in constructing high-efficiency local NLIDB.
As one of the solutions, CLLMs [69] are designed to address the above challenge with an enhanced model architecture and achieve a speedup for SQL generation.
Data Augmentation
During the fine-tuning process, the most straightforward factor affecting the model’s performance is the quality of the training labels [103]. The fine-tuning under the low quality or lack of the training labels is “making bricks without straw”, using high-quality or augmented data for fine-tuning always surpasses the meticulous design of fine-tuning methods on low-quality or raw data [104, 29]. Data-augmented fine-tuning in text-to-SQL made substantial progress, focusing on enhancing the data quality during the SFT process.
DAIL-SQL [8] are designed as an in-context learning framework, utilizing a sampling strategy for better few-shot instances. Incorporating the sampled instances in the SFT process improves the performance of open-source LLMs. Symbol-LLM [50] propose injection and infusion stage for data augmented instruction tuning. CodeS [9] augmented the training data with bi-directional generation with the help of ChatGPT. StructLM [70] are trained on multiple struct knowledge tasks for improving overall capability.
Pre-training
Pre-training is a fundamental phase of the complete fine-tuning process, aimed at acquiring text generation capabilities through auto-regressive training on extensive data [105]. Conventionally, the current powerful proprietary LLMs (e.g., ChatGPT [106], GPT-4 [78], Claude [107]) are pre-trained on hybrid corpus, which mostly benefit from the dialogue scenario that exhibits text generation capability [77]. The code-specific LLMs (e.g., CodeLLaMA [108], StarCoder [109]) are pre-trained on code data [87], and the mixture of various programming languages enables the LLMs to generate code to meet with the user’s instruction [110]. As a sub-task of code generation, the main challenge of SQL-specific pre-training technique is that the SQL/Database-related content occupies only a small portion of the entire pre-training corpus. Then, as a result, the open-source LLMs with comparatively limited comprehensive capacity (compared to ChatGPT, GPT-4) do not acquire a promising understanding of how to convert NL questions to SQL during their pre-training process.
The pre-training phase of the Codes [9] model consists of three stages of incremental pre-training. Starting from a basic code-specific LLM [109], CodeS are further pre-trained on a hybrid training corpus, including SQL-related data, NL-to-Code data, and NL-related data. The text-to-SQL understanding and performance are significantly improved.
Decomposition
Decomposing a task into multiple steps or using multiple models to solve the task is an intuitive solution for addressing a complex scenario, as we previously introduced in Sec. IV-A, ICL paradigm. The proprietary models utilized in ICL-based methods have a massive number of parameters that are not at the same parameter level as the open-source models used in fine-tuning methods. These models inherently possess the capability to perform assigned sub-tasks well (through mechanisms such as few-shot learning) [30, 57]. Thus, to replicate the success of this paradigm in ICL methods, it is necessary to reasonably assign corresponding sub-tasks to open-source models (such as generating external knowledge, schema linking, and distilling the schema) for sub-task-specific fine-tuning and constructing the corresponding data for fine-tuning, thereby assisting in the final SQL generation.
DTS-SQL [71] proposed a two-stage decomposed text-to-SQL fine-tuning framework and designed a schema-linking pre-generation task ahead of the final SQL generation.
V Expectations
Despite the significant advancements made in text-to-SQL research, there are still several challenges that need to be addressed. In this section, we discuss the remaining challenges that we expect to overcome in future work.
V-A Robustness in Real-world Applications
The text-to-SQL implemented by LLMs is expected to perform generalization and robustness across complex scenarios in real-world applications. Despite recent advances having made substantial progress in robustness-specific datasets [37, 41], its performance still falls short of practical application [33]. There are still challenges that are expected to be overcome in future studies. From the user aspect, there is a phenomenon that the user is not always a clear question proposer, which means the user questions might not have the exact database value and also can be varied from the standard datasets, the synonyms, typos, and vague expressions could be included [40]. For instance, the models are trained on clear indicative questions with concrete expressions in the fine-tuning paradigm. Since the model has not learned the mapping of realistic questions to the corresponding database, this leads to a knowledge gap when applied to real-world scenarios [33]. As reported in the corresponding evaluations of the dataset with synonym and incomplete instruction [26, 51], the SQL queries generated by ChatGPT contain around 40% incorrect execution, which is 10% lower than the original evaluation [51]. Simultaneously, the fine-tuning with local text-to-SQL datasets may contain non-standardized samples and labels. As an example, the name of the table or column is not always an accurate representation of its content, which yields an inconsistency within the training data construction and may lead to a semantic gap between the database schema and the user question. To address this challenge, aligning the LLMs with intention bias and designing the training strategy towards noisy scenarios will benefit the recent advances. At the same time, the data size in real-world applications is relatively smaller than the research-oriented benchmark. Since extending a large amount of the data by human annotation incurs high labor costs, designing data-augmentation methods to obtain more question-SQL pairs will support the LLM in data scarcity. Also, the adaptation study of fine-tuned open-source LLM to the local small-size dataset can be potentially beneficial. Furthermore, the extensions on multi-lingual [42, 111] and multi-modal scenarios [112] should be studied comprehensively in future research, which will benefit more language groups and help build more general database interfaces.
V-B Computational Efficiency
The computational efficiency is determined by the inference speed and the cost of computational resources, which is worth considering in both application and research work [69, 49]. With the increasing complexity of databases in up-to-date text-to-SQL benchmarks [15, 33], databases will carry more information (including more tables and columns), and the token length of the database schema will correspondingly increase, raising a series of challenges. Dealing with an ultra-complex database, taking the corresponding schema as input may encounter the challenge that the cost of calling proprietary LLMs will significantly increase, potentially exceeding the model’s maximum token length, especially with the implementation of open-source models that have shorter context lengths. Meanwhile, another obvious challenge is that most works use the full schema as model input, which introduces significant redundancy [57]. Providing LLMs with a precise question-related filtered schema directly from the user end to reduce cost and redundancy is a potential solution to improve computational efficiency [30]. Designing an accurate method for schema filtering remains a future direction. Although the in-context learning paradigm achieves promising accuracy, as a computational efficiency concern, the well-designed methods with the multi-stage framework or extended context increasing the number of API calls to enhance performance has simultaneously led to a substantial rise in costs [7]. As reported in related approaches [49], a trade-off between performance and computational efficiency should be considered carefully, and designing a comparable (even better) in-context learning method with less API cost will be a practical implementation and is still under exploration. Compared to PLM-based methods, the inference speed of LLM-based methods is observably slower [28, 21]. Accelerating inference by shortening the input length and reducing the number of stages in implementation would be intuitive for the in-context learning paradigm. For local LLMs, from a starting point [69], more speedup strategies can be studied in enhancing the model’s architecture in future exploration.
V-C Data Privacy and Interpretability
As a part of the LLMs’ study, LLM-based text-to-SQL also faces some general challenges present in LLM research [113, 4, 114]. Potential improvements from the text-to-SQL perspective are also expected to be seen in these challenges, thereby extensively benefiting the study of LLMs. As previously discussed in Sec. IV-A, the in-context learning paradigm predominates the number and performance in recent studies, with the majority of work using proprietary models for implementation [7, 8]. A straightforward challenge is proposed regarding data privacy, as calling proprietary APIs to handle local databases with confidentiality can pose a risk of data leakage. Using a local fine-tuning paradigm can partially address this issue. Still, the current performance of vanilla fine-tuning is not ideal [8], and advanced fine-tuning framework potentially relies on proprietary LLMs for data augmentation [9]. Based on the current status, more tailored frameworks in the local fine-tuning paradigm for text-to-SQL deserve widespread attention. Overall, the development of deep learning continually faces challenges regarding interpretability [114, 115]. As a long-standing challenge, considerable work has already been studied to address this issue [116, 117]. However, in text-to-SQL research, the interpretability of LLM-based implementation is still not being discussed, whether in the in-context learning or fine-tuning paradigm. The approaches with a decomposition phase explain the text-to-SQL implementation process from the perspective of step-by-step generation [7, 51]. Building on this, combining advanced study in interpretability [118, 119] to enhance text-to-SQL performance and interpreting the local model architecture from the database knowledge aspect remain future directions.
V-D Extensions
As a sub-field of LLMs and natural language understanding research, many studies in these fields have been adopted for text-to-SQL tasks, advancing its development [89, 95]. However, text-to-SQL research can also be extended to the larger scope studies of these fields at meanwhile. For instance, SQL generation is a part of code generation. The well-designed approaches in code generation also obtain promising performance in text-to-SQL [68, 48], performing generalization across various programming languages. The potential extension of some tailored text-to-SQL frameworks to NL-to-code studies can also be discussed. For instance, frameworks integrating execution output in NL-to-code can also achieve solid performance in SQL generation [7]. An attempt to extend execution-aware approaches in text-to-SQL with other advancing modules [31, 30] for code generation is worth discussing. From another perspective, we previously discussed that text-to-SQL can enhance LLM-based question-answering (QA) by providing factual information. The database can store relational knowledge as structural information, and the structure-based QA can potentially benefit from text-to-SQL (e.g., knowledge-based question-answering, KBQA [120]). Construct the factual knowledge with database structure, and then incorporate the text-to-SQL system to achieve information retrieval, which can potentially assist further QA with more accurate factual knowledge [121]. More extensions of text-to-SQL studies are expected in future work.

### v2

@@
 Abstract
-Generating accurate SQL according to natural language questions (text-to-SQL) is a long-standing problem since it is challenging in user question understanding, database schema comprehension, and SQL generation.
-Conventional text-to-SQL systems include human engineering and deep neural networks.
+Generating accurate SQL according to natural language questions (text-to-SQL) is a long-standing challenge due to the complexities involved in user question understanding, database schema comprehension, and SQL generation.
+Conventional text-to-SQL systems, comprising human engineering and deep neural networks, have made substantial progress.
 Subsequently, pre-trained language models (PLMs) have been developed and utilized for text-to-SQL tasks, achieving promising performance.
-As modern databases become more complex and corresponding user questions more challenging, PLMs with limited comprehension capabilities can lead to incorrect SQL generation.
-This necessitates more sophisticated and tailored optimization methods, which, in turn, restricts the applications of PLM-based systems.
-Most recently, large language models (LLMs) have demonstrated significant abilities in natural language understanding as the model scale remains increasing.
-Therefore, integrating the LLM-based implementation can bring unique opportunities, challenges, and solutions to text-to-SQL research.
+As modern databases become more complex, the corresponding user questions also grow more challenging, leading PLMs with limited comprehension capabilities to produce incorrect SQL.
+This necessitates more sophisticated and tailored optimization methods for PLMs, which, in turn, restricts the applications of PLM-based systems.
+Most recently, large language models (LLMs) have demonstrated significant capabilities in natural language understanding as the model scale remains increasing.
+Therefore, integrating the LLM-based implementation can bring unique opportunities, improvements, and solutions to text-to-SQL research.
 In this survey, we present a comprehensive review of LLM-based text-to-SQL.
-Specifically, we propose a brief overview of the current challenges and the evolutionary process of text-to-SQL.
+Specifically, we propose a brief overview of the technical challenges and the evolutionary process of text-to-SQL.
 Then, we provide a detailed introduction to the datasets and metrics designed to evaluate text-to-SQL systems.
 After that, we present a systematic analysis of recent advances in LLM-based text-to-SQL.
-Finally, we discuss the remaining challenges in this field and propose expectations for future directions.
+Finally, we discuss the remaining challenges in this field and propose expectations for future research directions.
 I Introduction
-Figure 1: An example for LLM-based text-to-SQL selected from Spider.
-The user proposes a question, “What cartoons were written by Joseph Kuhr?” the LLM takes the question and the schema of its corresponding database as the input, then generates a SQL query as the output.
-The SQL query can be executed in the database and retrieve a content “Batman Series” to answer the user question.
+Figure 1: An example for LLM-based text-to-SQL selected from the BIRD dataset.
+A user proposes a question about football leagues.
+The LLM takes the question and the schema of its corresponding database as the input and then generates an SQL query as the output.
+The SQL query can be executed in the database and retrieve the content “The 5 leagues with the highest matches” to answer the user question.
 Text-to-SQL is a long-standing task in natural language processing research.
@@
 Fig.
-1 provides an example of the large language model-based (LLM-based) text-to-SQL; given a user question “What cartoons were written by Joseph Kuhr?”, the LLMs take the question and its corresponding database schema as the input and generate an SQL query as the output, which can be executed in the database to retrieve the content “Batman Series” for answering the user question.
+1 provides an example of a large language model-based (LLM-based) text-to-SQL system.
+Given a user question such as “Could you tell me the names of the 5 leagues with the highest matches of all time and how many matches were played in the said league?”, the LLM takes the question and its corresponding database schema as input and then generates an SQL query as output.
+This SQL query can be executed in the database to retrieve the relevant content to answer the user’s question.
 The above system builds a natural language interface to the database (NLIDB) with LLMs.
-Since SQL remains one of the most widely used programming languages, with over half (51.52%) of professional developers using SQL in their work, but only around a third (35.29%) of those systematically trained [^1] https://survey.stackoverflow.co/2023, the NLIDB enables non-skilled users to access structured databases like professional database engineers [1, 2] and also accelerates human-computer interaction [3].
-Furthermore, amid the research hotspot of LLMs, text-to-SQL can provide a solution to the prevalent hallucination [4, 5] issue by incorporating realistic content from the database to fill the knowledge gaps of LLMs [6].
-The significant value and potential for text-to-SQL have triggered a range of studies on its incorporation and optimization with LLMs [7, 8, 9]; consequently, LLM-based text-to-SQL remains a highly discussed research field within the NLP and database communities.
+Since SQL remains one of the most widely used programming languages, with over half (51.52%) of professional developers using SQL in their work, it is notable that only around a third (35.29%) of those developers are systematically trained [^1] https://survey.stackoverflow.co/2023, the NLIDB enables non-skilled users to access structured databases like professional database engineers [1, 2] and also accelerates human-computer interaction [3].
+Furthermore, amid the research hotspot of LLMs, text-to-SQL can provide a potential solution to the prevalent hallucination [4, 5] issue by incorporating realistic content from the database to fill the knowledge gaps of LLMs [6].
+The significant value and potential for text-to-SQL have triggered a range of studies on its integration and optimization with LLMs [7, 8, 9, 10]; consequently, LLM-based text-to-SQL remains a highly discussed research field within the NLP and database communities.
 Previous studies have made notable progress in the implementation of text-to-SQL and have undergone a long evolutionary process.
-Early efforts were mostly based on well-designed rules and templates [10, 11], specifically suitable for simple database scenarios.
+Early efforts were mostly based on well-designed rules and templates [11], specifically suitable for simple database scenarios.
 In recent years, with the heavy labor costs [12] brought by rule-based methods and the growing complexity of database environments [13, 14, 15], designing a rule or template for each scenario has become increasingly difficult and impractical.
 The development of deep neural networks has advanced the progress of text-to-SQL [16, 17], which can automatically learn a mapping from the user question to its corresponding SQL [18, 19].
-Subsequently, pre-trained language models (PLMs) with strong semantic parsing capacity have become the new paradigm for text-to-SQL [20], taking its performance to a new level [21, 22, 23].
-Incrementally, PLM-based research focused on table content encoding [24, 19, 25] and pre-training [20, 11] has further advanced this field.
-Recently, the LLM-based approaches implementing text-to-SQL through in-context learning (ICL) [7] and supervised fine-tuning (SFT) [9] paradigm, reaching state-of-the-art accuracy with the well-designed framework and stronger comprehension capability compared to PLMs.
+Subsequently, pre-trained language models (PLMs) with strong semantic parsing capacity have become the new paradigm for text-to-SQL systems [20], boosting their performance to a new level [21, 22, 23].
+Incremental research on PLM-based optimization, such as table content encoding [24, 19, 25] and pre-training [20, 26], has further advanced this field.
+Most recently, the LLM-based approaches implementing text-to-SQL through in-context learning (ICL) [8] and fine-tuning (FT) [10] paradigm, reaching state-of-the-art accuracy with the well-designed framework and stronger comprehension capability compared to PLMs.
 The overall implementation details of LLM-based text-to-SQL can be divided into 3 aspects: 1.
 Question understanding: The NL question is a semantic representation of the user’s intention, which the corresponding generated SQL query is expected to align with; 2.
-Schema comprehension: The schema provides the table and column structure of the database, and the text-to-SQL system is required to identify the target components in the database that match the user question; 3.
-SQL generation: This involves incorporating the above parsing and then writing correct syntax to generate executable SQL queries that can retrieve the desired answer.
-The LLMs have proven to perform a good vanilla implementation [26, 27], benefiting from the more powerful semantic parsing enabled by the richer training corpus compared to the PLMs [28, 29].
+Schema comprehension: The schema provides the table and column structure of the database, and the text-to-SQL system is required to identify the target components that match the user question; 3.
+SQL generation: This involves incorporating the above parsing and then predicting the correct syntax to generate executable SQL queries that can retrieve the required answer.
+The LLMs have proven to perform a good vanilla implementation [7, 27], benefiting from the more powerful semantic parsing capacity enabled by the richer training corpus [28, 29].
 Further studies on enhancing the LLMs for question understanding [8, 9], schema comprehension [30, 31], and SQL generation [32] are being increasingly released.
 Figure 2: A sketch of the evolutionary process for text-to-SQL research from the perspective of implementation paradigm.
-Each stage is presented with some implementation techniques and representative works.
-The timestamps for the stages are not exactly accurate; we set each timestamp according to the release time of the representative works of each paradigm, with a margin of error of about one year before and after.
-The format is inspired from [29]
+Each stage is presented with two representative implementation techniques.
+The timestamps for the stages are not exactly accurate; we set each timestamp according to the release time of the representative works of each implementation paradigm, with a margin of error of about one year before and after.
+The format is inspired from [29].
 […]
@@
 Datasets III-A) Original Datasets & Post-annotated Datasets Cross- domain BIRD [33], DuSQL [34], CoSQL [35], Spider [13], WikiSQL [14], KaggleDBQA [36], ADVETA [37] Spider-$S$ [38], Spider-CG [38], Spider-DK [39], Spider-SYN [40], Spider-Realistic [41], CSpider [42] SParC [43] Knowledge- augmented BIRD [33], SQUALL [44], Spider-DK [39] Cross- lingual DuSQL [34], CSpider [42] Context- dependent CoSQL [35], Spider-$S$ [38], Spider-CG [38], SparC [43] Robustness ADVETA [37], Spider-SYN [40], Spider-Realistic [41] Evaluation Metrics III-B) Content- Matching based Component Matching (CM) [13], Exact Matching (EM) [13] Execution based Execution Accuracy (EX) [13], Valid Efficiency Score (VES) [33] Methods IV) In-context Learning Paradigm IV-A) Zero-shot [7], [33], [27], [45], [46], [47], [48], [49], [9], [50] Few-shot [33], [8], [51], [52], [53], [54], [49], [9], [32], [55] Decomposition Coder-Reviewer [56], DIN-SQL [8], QDecomp [51], C3 [30], MAC-SQL [57], DEA-SQL [58], SGU-SQL [32], MetaSQL [59], PET-SQL [60], PURPLE [61] Prompt Optimization DESEM+P [62], StructGPT [63], SD+SA+Voting [52]RAG+SP&DRC [64], C3 [30], DAIL-SQL [9] ODIS [54], ACT-SQL [49], FUSED [65], DELLM [31] Reasoning Enhancement CoT [51, 33, 9, 32], QDecomp [51], Least-to-Most [51], SQL-PaLM [53], ACT-SQL [49], POT [55] SQL-CRAFT [55], FUXI [66] Execution Refinement MBR-Exec [67], Coder-Reviewer [56], LEVER [68], SELF-DEBUGGING [48], DESEM+P [62] DIN-SQL [8], SD+SA+Voting [52], SQL-PaLM [53], RAG+SP&DRC [64], C3 [30], MAC-SQL [57] DELLM [31], SQL-CRAFT [55], FUXI [66], PET-SQL [60], PURPLE [61] Fine-tuning Paradigm IV-B) Supervised Fine-tuning [45], [9], [50], [53] Enhanced Architecture CLLMs [69] Pre-training CodeS [10] Data Augmentation DAIL-SQL [9], Symbol-LLM [50], CodeS [10], StructLM [70] Decomposition DTS-SQL [71]
-Figure 3: Taxonomy tree of the research in LLM-based text-to-SQL. The display order in each node is organized by the released time. The format is adapted from [72]
-I-A Challenges in Text-to-SQL
+Figure 3: Taxonomy tree of the research in LLM-based text-to-SQL. The display order in each node is organized by the released time. The format is adapted from [72].
 Despite the significant progress made in text-to-SQL research, several challenges remain that hinder the development of robust and generalized text-to-SQL systems [73].
-In this survey, we aim to catch up with the recent advances and provide a comprehensive review of the current state-of-the-art (SOTA) models and approaches in LLM-based text-to-SQL.
+In this survey, we aim to catch up with the recent advances and provide a comprehensive review of the current state-of-the-art models and approaches in LLM-based text-to-SQL.
 We begin by introducing the fundamental concepts and challenges associated with text-to-SQL, highlighting the importance of this task in various domains.
-We then delve into the evolution of LLMs and their application to text-to-SQL, discussing the key advancements and breakthroughs in this field.
-After the overview, we provide a detailed introduction to the recent advances of text-to-SQL incorporating LLMs.
+We then delve into the evolution of the implementation paradigm for text-to-SQL systems, discussing the key advancements and breakthroughs in this field.
+After the overview, we provide a detailed introduction and analysis of the recent advances in text-to-SQL integrating LLMs.
 Specifically, the body of our survey covers a range of contents related to LLM-based text-to-SQL, including:
-  • Datasets and Benchmarks: We provide an overview of the commonly used datasets and benchmarks for evaluating LLM-based Text-to-SQL systems.
-We discuss their characteristics, complexity, and the challenges they pose for text-to-SQL system development and evaluation.
-  • Evaluation Metrics: We present the evaluation metrics used to assess the performance of LLM-based Text-to-SQL systems, including accuracy, exactness, and execution correctness.
-We discuss the advantages and limitations of each metric and their relevance to real-world applications.
-  • Methods and Models: We explore the different methods and models employed for LLM-based text-to-SQL, including in-context learning and fine-tuning-based paradigms.
-We discuss their implementation details, strengths, and adaptations specific to the text-to-SQL task.
-  • Expectations and Future Directions: We discuss the current challenges and limitations of LLM-based Text-to-SQL, such as real-world robustness, computational efficiency, data privacy and extensions.
-We also outline potential future research directions and opportunities for improvement.
-We hope this survey will provide a clear overview of recent studies and inspire future studies. A taxonomy tree is shown in Fig. 3
+  • Datasets and Benchmarks: We provide a detailed introduction to the commonly used datasets and benchmarks for evaluating LLM-based text-to-SQL systems.
+We discuss their characteristics, complexity, and the challenges they pose for text-to-SQL development and evaluation.
+  • Evaluation Metrics: We present the evaluation metrics used to assess the performance of LLM-based text-to-SQL systems, including both content matching-based and execution-based paradigms.
+We then briefly introduce the characteristics of each metric.
+  • Methods and Models: We present a systematic analysis of the different methods and models employed for LLM-based text-to-SQL, including in-context learning and fine-tuning-based paradigms.
+We discuss their implementation details, strengths, and adaptations specific to the text-to-SQL task from various implementation perspectives.
+  • Expectations and Future Directions: We discuss the remaining challenges and limitations of LLM-based text-to-SQL, such as real-world robustness, computational efficiency, data privacy, and extensions.
+We also outline potential future research directions and opportunities for improvement and optimization.
+We hope this survey provides a clear overview of recent studies and inspires future research. Fig. 3 shows a taxonomy tree that summarizes the structure and contents of our survey.
 II Overview
-Text-to-SQL is a task that aims to convert natural language-type questions into corresponding SQL queries that can be executed on a relational database.
-Formally, given a user question $\mathcal{Q}$ (also known as a user query, natural language, NL, NL question, etc.) and a database schema $\mathcal{D}$, the goal is to generate an SQL query $Y$ that can be executed on the database to obtain the desired answer.
-Text-to-SQL has the potential to democratize access to data by allowing users to interact with databases using natural language without the need for specialized knowledge of SQL.
-This can benefit various domains, such as business intelligence, customer support, and scientific research, by enabling non-technical users to easily retrieve information from databases and facilitating more efficient data analysis.
-II-1 Linguistic Complexity and Ambiguity
-Natural language questions often contain complex linguistic structures, such as nested clauses, coreferences, and ellipses, which make it challenging to map them accurately to SQL queries.
-Additionally, natural language is inherently ambiguous, with multiple possible interpretations for a given question.
-Resolving these ambiguities and understanding the intent behind the question requires deep language understanding and the ability to incorporate context and domain knowledge.
-II-2 Schema Understanding and Representation
-To generate accurate SQL queries, text-to-SQL systems need to have a comprehensive understanding of the database schema, including table names, column names, and relationships between tables.
-However, database schemas can be complex and vary significantly across different domains.
+Text-to-SQL is a task that aims to convert natural language questions into corresponding SQL queries that can be executed in a relational database.
+Formally, given a user question $\mathcal{Q}$ (also known as a user query, natural language question, etc.) and a database schema $\mathcal{S}$, the goal of the task is to generate an SQL query $Y$ that retrieves the required content from the database to answer the user question.
+Text-to-SQL has the potential to democratize access to data by allowing users to interact with databases using natural language without the need for specialized knowledge of SQL programming [74].
+This can benefit various domains, such as business intelligence, customer support, and scientific research, by enabling non-skilled users to easily retrieve target content from databases and facilitating more efficient data analysis.
+II-A Challenges in Text-to-SQL
+The technical challenges for text-to-SQL implementations can be summarized as follows:
+II-A1 Linguistic Complexity and Ambiguity
+Natural language questions often contain complex linguistic representations, such as nested clauses, coreferences, and ellipses, which make it challenging to map them accurately to the corresponding part of SQL queries [41].
+Additionally, natural language is inherently ambiguous, with multiple possible representations for a given user question [75, 76].
+Resolving these ambiguities and understanding the intent behind the user question requires deep natural language understanding and the capability to incorporate context and domain knowledge [33].
+II-A2 Schema Understanding and Representation
+To generate accurate SQL queries, text-to-SQL systems need to have a comprehensive understanding of the database schema, including table names, column names, and relationships between various tables.
+However, the database schema can be complex and vary significantly across different domains [13].
 Representing and encoding the schema information in a way that can be effectively utilized by the text-to-SQL model is a challenging task.
-II-3 Rare and Complex SQL Operations
-Some SQL queries involve rare or complex operations, such as nested subqueries, outer joins, and window functions.
-These operations are less frequent in the training data and pose challenges for text-to-SQL models to generate accurately.
-Designing models that can handle a wide range of SQL operations, including rare and complex ones, is an important consideration.
-II-4 Cross-Domain Generalization
-Text-to-SQL models often struggle to generalize across different database schemas and domains.
-Models trained on a specific domain may not perform well on questions from a different domain due to differences in vocabulary, schema structure, and question patterns.
-Developing models that can effectively adapt to new domains with minimal fine-tuning or domain-specific training data is an ongoing challenge.
-II-A Evolutionary Process
-The field of text-to-SQL has witnessed significant advancements over the years, evolving from rule-based approaches to deep learning-based methods and, more recently, to the integration of pre-trained language models (PLMs) and large language models (LLMs), as shown in Fig.
+II-A3 Rare and Complex SQL Operations
+Some SQL queries involve rare or complex operations and syntax in challenging scenarios, such as nested sub-queries, outer joins, and window functions.
+These operations are less frequent in the training data and pose challenges for text-to-SQL systems to generate accurately.
+Designing models that can be generalized to a wide range of SQL operations, including rare and complex scenarios, is an essential consideration.
+II-A4 Cross-Domain Generalization
+Text-to-SQL systems often struggle to generalize across various database scenarios and domains.
+Models trained on a specific domain may not perform well on the proposed questions from other domains due to the variety in vocabulary, database schema structure, and question patterns.
+Developing systems that can effectively generalized to new domains with minimal domain-specific training data or fine-tuning adaptation is a significant challenge [77].
+II-B Evolutionary Process
+The research field of text-to-SQL has witnessed significant advancements over the years in the NLP community, having evolved from rule-based methods to deep learning-based approaches and, more recently, to integrating pre-trained language models (PLMs) and large language models (LLMs), a sketch of the evolutionary process is shown in Fig.
 2.
-II-A1 Rule-based Approaches
-Early text-to-SQL systems relied heavily on rule-based approaches [10, 12, 11], where manually crafted rules and heuristics were used to map natural language questions to SQL queries.
+II-B1 Rule-based Methods
+Early text-to-SQL systems relied heavily on rule-based methods [11, 12, 26], where manually crafted rules and heuristics were used to map natural language questions to SQL queries.
 These approaches often involved extensive feature engineering and domain-specific knowledge.
-While rule-based methods achieved success in specific domains, they lacked the flexibility and generalization capabilities needed to handle diverse and complex questions.
+While rule-based methods achieved success in specific simple domains, they lacked the flexibility and generalization capabilities needed to handle diverse and complex questions.
+II-B2 Deep Learning-based Approaches
+With the rise of deep neural networks, sequence-to-sequence models and encoder-decoder structures, such as LSTMs [78] and transformers [17], were adapted to generate SQL queries from natural language input [79, 19].
+Typically, RYANSQL [19] introduced techniques like intermediate representations and sketch-based slot filling to handle complex questions and improve cross-domain generalization.
+Recently, researchers introduced graph neural networks (GNNs) for text-to-SQL tasks by leveraging schema dependency graphs to capture the relationships between database elements [18, 80].
+II-B3 PLM-based Implementation
+Pre-trained language models (PLMs) have emerged as a powerful solution for text-to-SQL, leveraging the vast amounts of linguistic knowledge and semantic understanding captured during the pre-training process.
+The early adoption of PLMs in text-to-SQL primarily focused on fine-tuning off-the-shelf PLMs, such as BERT [24] and RoBERTa [81], on standard text-to-SQL datasets [13, 14].
+These PLMs, pre-trained on large amounts of training corpus, captured rich semantic representations and language understanding capabilities.
+By fine-tuning them on text-to-SQL tasks, researchers aimed to leverage the semantic and linguistic understanding of PLMs to generate accurate SQL queries [79, 20, 82].
+Another line of research focuses on incorporating schema information into PLMs to improve their understanding of database structures and enable them to generate more executable SQL queries.
+Schema-aware PLMs are designed to capture the relationships and constraints present in the database structure [21].
+II-B4 LLM-based Implementation
+Large language models (LLMs), such as GPT series [83, 84, 85], have gained significant attention in recent years due to their capability to generate coherent and fluent text.
+Researchers have started exploring the potential of LLMs for text-to-SQL by leveraging their extensive knowledge reserves and superior generation capabilities [7, 9].
+These approaches often involve prompt engineering to guide the proprietary LLMs in SQL generation [47] or fine-tuning the open-source LLMs on text-to-SQL datasets [9].
+The integration of LLMs in text-to-SQL is still an emerging research area with significant potential for further exploration and improvement.
+Researchers are investigating approaches to better leverage the knowledge and reasoning capabilities of LLMs, incorporate domain-specific knowledge [33, 31], and develop more efficient fine-tuning strategies [10].
+As the field continues to evolve, we anticipate the development of more advanced and superior LLM-based implementations that will elevate the performance and generalization capabilities of text-to-SQL to new heights.
+III Benchmarks & evalution
+In this section, we introduce the benchmarks for text-to-SQL, encompassing well-known datasets and evaluation metrics.
 TABLE I: The statistics and analysis of well-known datasets of text-to-SQL ordered by release time.
@@
 WikiSQL [14] Aug-2017 80,654 26,521 1 17 Cross-domain
-Post-annotated Dataset Release Time Base Dataset Special Setting Characteristics
+Post-annotated Dataset Release Time Source Dataset Special Setting Characteristics
 ADVETA [37] Dec-2022 Spider, etc. Adversarial table perturbation Robustness
@@
 SParC [43] Jun-2019 Spider Annotate conversational contents Context-dependent
-II-A2 Deep Learning-based Methods
-With the rise of deep learning, sequence-to-sequence models, such as LSTMs and transformers, were adapted to generate SQL queries from natural language input [74, 19].
-Typically, RYANSQL [19] introduced techniques like intermediate representations and sketch-based slot filling to handle complex questions and improve cross-domain generalization.
-Recently, researchers introduced graph neural networks (GNNs) for text-to-SQL tasks by leveraging schema dependency graphs to capture the relationships between database elements [18, 75].
-II-A3 PLM-based Approaches
-Pre-trained language models (PLMs) have emerged as a powerful solution for text-to-SQL, leveraging the vast amounts of linguistic knowledge and understanding captured during pre-training.
-The early adoption of PLMs in text-to-SQL primarily focused on fine-tuning pre-existing PLMs, such as BERT [24] and RoBERTa [76], on text-to-SQL datasets.
-These PLMs, which were pre-trained on large amounts of text data, captured rich semantic representations and language understanding capabilities.
-By fine-tuning them on text-to-SQL tasks, researchers aimed to leverage the knowledge and linguistic understanding of PLMs to generate accurate SQL queries [74, 20].
-Another line of research focuses on incorporating schema information into PLMs to improve their understanding of database structures and generate more accurate SQL queries.
-Schema-aware PLMs are designed to capture the relationships and constraints present in the database schema [21].
-II-A4 LLM-based Implementation
-Large language models (LLMs), such as GPT series [77, 78], have gained significant attention in recent years due to their ability to generate coherent and fluent text.
-Researchers have started exploring the potential of LLMs for text-to-SQL by leveraging their vast knowledge and generative capabilities [26, 8].
-These approaches often involve prompt engineering to guide the proprietary LLMs in the generation process or fine-tuning the open-source LLMs on text-to-SQL datasets.
-The integration of LLMs in text-to-SQL is still an emerging area, and there is significant potential for further exploration and improvement.
-Researchers are investigating ways to better leverage the knowledge and reasoning capabilities of LLMs, incorporate domain-specific knowledge, and develop more efficient fine-tuning strategies.
-As the field continues to evolve, we can expect to see more advanced and powerful LLM-based approaches that push the boundaries of text-to-SQL performance and generalization.
-III Benchmarks & evalution
-In this section, we introduce the benchmarks for text-to-SQL, encompassing well-known datasets and evaluation metrics.
 III-A Datasets
-As shown in Table I, we classify the datasets into ”Original Datasets” and ”Post annotated Datasets” based on whether they were released with the original dataset and databases or created by adapting existing datasets and databases with special processing.
-We also highlight their release times.
-For the original datasets, we provide a detailed analysis, including the number of examples, databases, tables per database, and rows per database.
-For the post-annotated datasets, we identify their base dataset and describe the special processing applied to them.
-To illustrate the potential opportunities of each dataset, we categorize their characteristics into Cross-domain, Knowledge-augmented, Context-dependent, Robustness, and Cross-lingual, which we will discuss in detail below.
-III-A1 Cross-domain Dataset.
+As shown in Tab.
+I, we classify the datasets into ’Original Datasets’ and ’Post-annotated Datasets’ based on whether they were released with the original dataset and databases or created by adapting existing datasets and databases with special settings.
+For the original datasets, we provide a detailed analysis, including the number of examples, the number of databases, the number of tables per database, and the number of rows per database.
+For the post-annotated datasets, we identify their source dataset and describe the special setting applied to them.
+To illustrate the potential opportunities of each dataset, we annotate them based on their characteristics.
+The annotations are listed in the rightmost column of Tab.
+I, which we will discuss in detail below.
+III-A1 Cross-domain Dataset
 This refers to datasets where the background information of different databases comes from various domains.
-Since real-world text-to-SQL tasks often involve databases from multiple domains, most original text-to-SQL datasets [14, 13, 35, 34, 36, 33] and post annotated datasets [43, 42, 39, 40, 41, 37, 38] are in the cross-domain setting to fit well with the needs of cross-domain applications.
+Since real-world text-to-SQL applications often involve databases from multiple domains, most original text-to-SQL datasets [14, 13, 35, 34, 36, 33] and post-annotated datasets [43, 42, 39, 40, 41, 37, 38] are in the cross-domain setting to fit well with the requirements of cross-domain applications.
 III-A2 Knowledge-augmented Dataset
-Interest in incorporating knowledge into text-to-SQL tasks has increased significantly in recent years.
-BIRD [33] employs domain experts to annotate each text-to-SQL sample with external knowledge, categorized into Numeric Reasoning Knowledge, Domain Knowledge, Synonym Knowledge, and Value Illustration.
-Similarly, Spider-DK [39] defines and adds five types of domain knowledge for a human-curated version of the Spider dataset: SELECT Columns Mentioned by Omission, Simple Inference Required, Synonyms Substitution in Cell Value Word, One Non-Cell Value Word Generate a Condition, and Easy to Conflict with Other Domains.
-Both studies found that human-annotated knowledge significantly improves SQL generation accuracy for samples requiring external domain knowledge.
+Interest in incorporating domain-specific knowledge into text-to-SQL tasks has increased significantly in recent years.
+BIRD [33] employs human database experts to annotate each text-to-SQL sample with external knowledge, categorized into numeric reasoning knowledge, domain knowledge, synonym knowledge, and value illustration.
+Similarly, Spider-DK [39] defines and adds five types of domain knowledge for a human-curated version of the Spider dataset [13]: SELECT columns mentioned by omission, simple inference required, synonyms substitution in cell value word, one non-cell value word generate a condition, and easy to conflict with other domains.
+Both studies found that human-annotated knowledge significantly improves SQL generation performance for samples requiring external domain knowledge.
 Additionally, SQUALL [44] manually annotates alignments between the words in NL questions and the entities in SQL, providing finer-grained supervision than other datasets.
@@
 SParC [43] and CoSQL [35] explore context-dependent SQL generation by constructing a conversational database querying system.
-Unlike traditional text-to-SQL datasets that only have a single (question, SQL) pair for one example, SParC decomposes the (question, SQL) examples in the Spider dataset into multiple (sub-question, SQL) pairs to construct a simulated and meaningful interaction, including inter-related sub-questions that aid SQL generation, and unrelated sub-questions that enhance data diversity.
-CoSQL, in comparison, involves conversational interactions in natural language, simulating real-life scenarios to increase complexity and diversity.
-Additionally, Spider-SS&CG [38] splits the NL question in the Spider dataset into multiple sub-questions and sub-SQLs, demonstrating that training on these sub-examples can improve a text-to-SQL model’s generalization ability on out-of-distribution samples.
+Unlike traditional text-to-SQL datasets that only have a single question-SQL pair for one example, SParC decomposes the question-SQL examples in the Spider dataset into multiple sub-question-SQL pairs to construct a simulated and meaningful interaction, including inter-related sub-questions that aid SQL generation, and unrelated sub-questions that enhance data diversity.
+CoSQL, in comparison, involves conversational interactions in natural language, simulating real-world scenarios to increase complexity and diversity.
+Additionally, Spider-SS&CG [38] splits the NL question in the Spider dataset [13] into multiple sub-questions and sub-SQLs, demonstrating that training on these sub-examples can improve a text-to-SQL system’s generalization ability on out-of-distribution samples.
 III-A4 Robustness Dataset
-Evaluating the accuracy of text-to-SQL models with attacked or perturbed database contents (e.g., schema and tables) is crucial for assessing robustness.
+Evaluating the accuracy of text-to-SQL systems with polluted or perturbed database contents (e.g., schema and tables) is crucial for assessing robustness.
 Spider-Realistic [41] removes explicit schema-related words from the NL questions, while Spider-SYN [40] replaces them with manually selected synonyms.
 ADVETA [37] introduces adversarial table perturbation (ATP), which perturbs tables by replacing original column names with misleading alternatives and inserting new columns with high semantic associations but low semantic equivalency.
-These perturbations lead to significant drops in accuracy, as a text-to-SQL model with low robustness may be misled by incorrect matches between tokens in NL questions and database entities.
+These perturbations lead to significant drops in accuracy, as a text-to-SQL system with low robustness may be misled by incorrect matches between tokens in NL questions and database entities.
 III-A5 Cross-lingual Dataset
@@
 III-B1 Content Matching-based Metrics
-SQL content matching metrics focus on comparing the generated SQL query with the ground truth SQL query based on their structural and syntactic similarities.
-  • Component Matching (CM) [13] evaluates the model performance by measuring the exact match between predicted and ground truth SQL components—SELECT, WHERE, GROUP BY, ORDER BY, and KEYWORDS—using the F1 score.
+SQL content matching metrics focus on comparing the predicted SQL query with the ground truth SQL query based on their structural and syntactic similarities.
+  • Component Matching (CM) [13] evaluates the performance of text-to-SQL system by measuring the exact match between predicted and ground truth SQL components—SELECT, WHERE, GROUP BY, ORDER BY, and KEYWORDS—using the F1 score.
 Each component is decomposed into sets of sub-components and compared for an exact match, accounting for SQL components without order constraints.
-  • Exact Matching (EM) [13] measures the percentage of examples whose predicted SQL query is equivalent to the ground truth SQL query.
-A predicted SQL is considered correct only if all its components, as described in Component Matching, match exactly with those of the ground truth query.
+  • Exact Matching (EM) [13] measures the percentage of examples whose predicted SQL query is identical to the ground truth SQL query.
+A predicted SQL is considered correct only if all its components, as described in CM, match exactly with those of the ground truth query.
 III-B2 Execution-based Metrics
 Execution result metrics assess the correctness of the generated SQL query by comparing the results obtained from executing the query on the target database with the expected results.
-  • Execution Accuracy (EX) [13] measures the correctness of a predicted SQL query by executing it in the corresponding database and comparing the executed results with the ground truth queries.
-  • Valid Efficiency Score (VES) [33] is defined to measure the efficiency of valid SQL queries, which are the predicted SQL queries whose executed results exactly match the ground truth query.
-Thus, VES evaluates both the efficiency and accuracy of predicted SQL queries.
-For an evaluation dataset with $N$ examples, VES can be computed by:
+  • Execution Accuracy (EX) [13] measures the correctness of a predicted SQL query by executing it in the corresponding database and comparing the executed results with the results obtained by the ground truth query.
+  • Valid Efficiency Score (VES) [33] is defined to measure the efficiency of valid SQL queries.
+A valid SQL query is a predicted SQL query whose executed results exactly match the ground truth results.
+Specifically, VES evaluates both the efficiency and accuracy of predicted SQL queries.
+For a text dataset with $N$ examples, VES can be computed by:
     $\text{VES}=\frac{1}{N}\sum_{n=1}^{N}\mathbbm{1}(V_{n},\hat{V}_{n})\cdot\textbf{R}(Y_{n},\hat{Y}_{n}),$ (1)
-    where $\hat{Y}_{n}$ and $\hat{V}_{n}$ are the predicted SQL query and its executed results.
-$Y_{n}$ and $V_{n}$ are those of the corresponding ground truth SQL query.
+    where $\hat{Y}_{n}$ and $\hat{V}_{n}$ are the predicted SQL query and its executed results and $Y_{n}$ and $V_{n}$ are the ground truth SQL query and its corresponding executed results, respectively.
 $\mathbbm{1}(V_{n},\hat{V}_{n})$ is an indicator function, where:
@@
     Then, $\textbf{R}(Y_{n},\hat{Y}_{n})=\sqrt{E(Y_{n})/E(\hat{Y}_{n})}$ denotes the relative execution efficiency of the predicted SQL query in comparison to ground-truth query, where $E(\cdot)$ is the execution time of each SQL in the database.
-[33] ensures the stability of this metric by computing the average of $\textbf{R}(Y_{n},\hat{Y}_{n})$ over 100 runs for each example.
-Most of the recent LLM-based text-to-SQL studies focus on these four datasets: Spider [13], Spider-Realistic [41], Spider-SYN [40], and BIRD [33]; and these three evaluation methods: EM, EX, and VES, we will primarily focus on them in the following introduction.
+BIRD benchmark [33] ensures the stability of this metric by computing the average of $\textbf{R}(Y_{n},\hat{Y}_{n})$ over 100 runs for each example.
+Most of the recent LLM-based text-to-SQL studies focus on these four datasets: Spider [13], Spider-Realistic [41], Spider-SYN [40], and BIRD [33]; and these three evaluation methods: EM, EX, and VES, we will primarily focus on them in the following analysis.
 IV Methods
-The implementation of LLM-based applications mostly relies on in-context learning (prompt engineering) [79] and fine-tuning [80] since the powerful proprietary models and well-architected open-source models are being released in large quantities [81, 82, 83].
-The specific methods can generally be divided into two paradigms, in-context learning (ICL) and fine-tuning (FT).
-In this survey, we primarily focus on these text-to-SQL paradigms and will discuss them accordingly.
+The implementation of current LLM-based applications mostly relies on in-context learning (ICL) (prompt engineering) [86, 87, 88] and fine-tuning (FT) [89, 90] paradigms since the powerful proprietary models and well-architected open-source models are being released in large quantities [85, 45, 91, 92, 93, 94].
+LLM-based text-to-SQL systems follow these paradigms for implementation.
+In this survey, we will discuss them accordingly.
 IV-A In-context Learning
-Through extensive and widely recognized research, prompt engineering has been proven to play a decisive role in the performance of LLMs [28], also impacting the text-to-SQL task in different prompt styles [8, 46].
+Through extensive and widely recognized research, prompt engineering has been proven to play a decisive role in the performance of LLMs [95, 28], also impacting the SQL generation in different prompt styles [9, 46].
 Necessarily, developing text-to-SQL methods in the in-context learning (ICL) paradigm is valuable for achieving promising improvement.
-The implementation of LLM-based text-to-SQL process to generate executable SQL query $Y$ is formulated as:
+The implementation of LLM-based text-to-SQL process to generate executable SQL query $Y$ can be formulated as:
 $Y=f(\mathcal{Q},\mathcal{S},\mathcal{I}\mid\theta),$ (3)
@@
 foreign key relationships [49], schema linking [30] and domain knowledge [33, 31]).
-$\mathcal{I}$ represents the instruction for the text-to-SQL task, which performs indicative guidance to guide the LLMs for generating an accurate SQL query.
+$\mathcal{I}$ represents the instruction for the text-to-SQL task, which provides indicative guidance to trigger the LLMs for generating an accurate SQL query.
 $f(\cdot\mid\theta)$ is a LLM with parameter $\theta$.
-In the in-context learning (ICL) paradigm, we utilize an off-the-shelf text-to-SQL model (i.e., parameter $\theta$ of the model is frozen) for implementation.
-Various well-designed methods have been adopted in the ICL paradigm for text-to-SQL.
-We group them into five categories $\mathbf{C}_{0:4}$, including: $\mathbf{C}_{0}$-Trivial Prompt, $\mathbf{C}_{1}$-Decomposition, $\mathbf{C}_{2}$-Prompt Optimization, $\mathbf{C}_{3}$-Reasoning Enhancement, and $\mathbf{C}_{4}$-Execution Refinement, the details are shown in Tab.
+In the in-context learning (ICL) paradigm, we utilize an off-the-shelf text-to-SQL model (i.e., parameter $\theta$ of the model is frozen) for generating the predicted SQL query.
+Various well-designed methods within the ICL paradigm have been adopted for LLM-based text-to-SQL tasks.
+We group them into five categories $\mathbf{C}_{0:4}$, including $\mathbf{C}_{0}$-Trivial Prompt, $\mathbf{C}_{1}$-Decomposition, $\mathbf{C}_{2}$-Prompt Optimization, $\mathbf{C}_{3}$-Reasoning Enhancement, and $\mathbf{C}_{4}$-Execution Refinement, the representative methods of each category are given in Tab.
 II.
 $\mathbf{C}_{0}$-Trivial Prompt
-Trained through massive data, LLMs have a strong overall proficiency in different downstream tasks with zero-shot and few-shot prompting [84, 80, 85], which is widely recognized and achieves promising results.
-In our survey, we categorized the above prompting approaches without the well-designed framework as trivial prompts.
-As introduce above, the formulated process of LLM-based text-to-SQL Eq.
-3 can also represent zero-shot prompting, the input $\mathcal{P}_{0}$ can be obtained by concatenate $\mathcal{I}$, $\mathcal{S}$ and $\mathcal{Q}$:
+Trained through massive data, LLMs have a strong overall proficiency in different downstream tasks with zero-shot and few-shot prompting [96, 89, 97], which is widely recognized and used in real-world applications.
+In our survey, we categorized the above prompting approaches without the well-designed framework as trivial prompts (vanilla prompt engineering).
+As introduced above, Eq.
+3 formulated the process of LLM-based text-to-SQL, which can also represent zero-shot prompting.
+The overall input $\mathcal{P}_{0}$ can be obtained by concatenating $\mathcal{I}$, $\mathcal{S}$ and $\mathcal{Q}$:
 $\mathcal{P}_{0}=\mathcal{I}\oplus\mathcal{S}\oplus\mathcal{Q}.$ (4)
-To regulate the prompting process, we set the OpenAI demonstration [^2] The prompt style that follows the official document from OpenAI platform: https://platform.openai.com/examples/default-sql-translate as the standard (trivial) prompt [30] for text-to-SQL.
+To regulate the prompting process, the OpenAI demonstration [^2] The prompt style that follows the official document from OpenAI platform: https://platform.openai.com/examples/default-sql-translate is set as the standard (trivial) prompt [30] for text-to-SQL.
 TABLE II: Typical methods used for in-context learning (ICL) in LLM-based text-to-SQL.
-The full table of existing methods with categorization $\mathbf{C}_{1:4}$ and more details are listed in Table III.
+The full table of existing methods with categorization $\mathbf{C}_{1:4}$ and more details are listed in Tab.
+III.
 Methods Adopted by Applied LLMs
@@
 TABLE III: Well-designed methods used in in-context learning (ICL) paradigm for LLM-based text-to-SQL ordered by release time.
-The methods are grouped in four categories based on their implementation perspective: $\mathbf{C}_{1}$-Decomposition, $\mathbf{C}_{2}$-Context Augmentation, $\mathbf{C}_{3}$-Reasoning Enhancement, $\mathbf{C}_{4}$-Execution Refinement.
+The methods are grouped in four categories based on their implementation perspective: $\mathbf{C}_{1}$-Decomposition, $\mathbf{C}_{2}$-Prompt Optimization, $\mathbf{C}_{3}$-Reasoning Enhancement, $\mathbf{C}_{4}$-Execution Refinement.
 The method in multiple categories will be introduced respectively.
 *There are multiple applied LLMs in the corresponding method; we present the selected LLM with representative performance.
-COT method are reported in multiple venues: NeurIPS’23 [33], EMNLP’23 [51], VLDB’24 [8], arXiv’24 [32]
+CoT method are reported in multiple venues: NeurIPS’23 [33], EMNLP’23 [51], VLDB’24 [9], arXiv’24 [32].
 Methods Applied LLMs Benchmark Metrics $\mathbf{C}_{1}$ $\mathbf{C}_{2}$ $\mathbf{C}_{3}$ $\mathbf{C}_{4}$ Release Time Publication Venue
@@
 PURPLE [61] GPT-4* [13, 40, 41] EX, EM Mar-2024 ICDE’24
-Zero-shot: Many research works [26, 27, 46]l utilize zero-shot prompting, focusing mainly on the influence of the style of prompt construction and the zero-shot performance of various LLMs for text-to-SQL.
-As an empirical evaluation, [26] evaluates the baseline text-to-SQL capabilities of different early-developed LLMs [86, 77, 87], which also presents the results for different prompt styles.
-The results indicate that prompt design is critical for performance, with error analysis, [26] propose more database content can harm the overall performance.
-Since ChatGPT emerged with impressive capabilities in conversational scenarios and code generation, [27] assesses its performance of text-to-SQL.
-With zero-shot settings, the results demonstrate that ChatGPT has a promising text-to-SQL ability compared to the state-of-the-art (SOTA) models.
+Zero-shot: Many research works [7, 27, 46] utilize zero-shot prompting, studying mainly on the influence of the style of prompt construction and the zero-shot performance of various LLMs for text-to-SQL.
+As an empirical evaluation, [7] evaluates the baseline text-to-SQL capabilities of different early-developed LLMs [98, 84, 99] and the performance for different prompting styles.
+The results indicate that prompt design is critical for the performance, with error analysis, [7] propose more database content can harm the overall accuracy.
+Since ChatGPT emerged with impressive capabilities in conversational scenarios and code generation [100], [27] assesses its performance of text-to-SQL.
+With zero-shot settings, the results demonstrate that ChatGPT has a promising text-to-SQL performance compared to the state-of-the-art PLM-based systems.
 For fair comparability, [47] reveal effective prompt construction for LLM-based text-to-SQL; they study different styles of prompt construction and make conclusions of zero-shot prompt design based on the comparisons.
-Primary key and foreign key carry contiguous knowledge of different tables, [49] study their impact by adding these keys on different designed database prompt style on zero-shot prompting results.
+Primary keys and foreign keys carry contiguous knowledge of different tables.
+[49] studies their impact by incorporating these keys into various prompt styles with different database content to analyze zero-shot prompting results.
 A benchmark evaluation [9] also studies the influence of foreign keys, with five different prompt representation styles, each style can be considered as the permutation and combination of the instruction, rule implication, and foreign key.
-Apart from the foreign key, this study also explores zero-shot prompting combined with “no explanation” rule and the rule implication “Let’s think step by step”.
-Empowered by the annotated knowledge of human experts, [33] follow the standard prompting and obtain improvement by combining the provided oracle knowledge.
+Apart from the foreign key, this study also explores zero-shot prompting combined with the rule implication “no explanation” to collect concise outputs.
+Empowered by the annotated external knowledge of human experts, [33] follows the standard prompting and achieves improvement by incorporating the provided annotated oracle knowledge.
 With the explosion of open-source LLMs, according to the results of similar evaluation, these models are also capable of zero-shot text-to-SQL task [45, 50, 46], especially code generation models [48, 46].
@@
 Few-shot: The technique of few-shot prompting is widely used in both practical applications and well-designed research, which has been proven efficient for eliciting better performance of LLMs [28, 101].
-The input prompt of the few-shot approach LLM-based text-to-SQL can be formulated as an extension of Eq.
+The overall input prompt of the few-shot approach LLM-based text-to-SQL can be formulated as an extension of Eq.
 3:

### v3

@@
 Abstract
-Generating accurate SQL according to natural language questions (text-to-SQL) is a long-standing challenge due to the complexities involved in user question understanding, database schema comprehension, and SQL generation.
+Generating accurate SQL from natural language questions (text-to-SQL) is a long-standing challenge due to the complexities in user question understanding, database schema comprehension, and SQL generation.
 Conventional text-to-SQL systems, comprising human engineering and deep neural networks, have made substantial progress.
 Subsequently, pre-trained language models (PLMs) have been developed and utilized for text-to-SQL tasks, achieving promising performance.
-As modern databases become more complex, the corresponding user questions also grow more challenging, leading PLMs with limited comprehension capabilities to produce incorrect SQL.
-This necessitates more sophisticated and tailored optimization methods for PLMs, which, in turn, restricts the applications of PLM-based systems.
-Most recently, large language models (LLMs) have demonstrated significant capabilities in natural language understanding as the model scale remains increasing.
-Therefore, integrating the LLM-based implementation can bring unique opportunities, improvements, and solutions to text-to-SQL research.
+As modern databases become more complex, the corresponding user questions also grow more challenging, causing PLMs with parameter constraints to produce incorrect SQL.
+This necessitates more sophisticated and tailored optimization methods, which, in turn, restricts the applications of PLM-based systems.
+Recently, large language models (LLMs) have demonstrated significant capabilities in natural language understanding as the model scale increases.
+Therefore, integrating LLM-based implementation can bring unique opportunities, improvements, and solutions to text-to-SQL research.
 In this survey, we present a comprehensive review of LLM-based text-to-SQL.
@@
 Finally, we discuss the remaining challenges in this field and propose expectations for future research directions.
+Index Terms:
+text-to-SQL, large language models, database, natural language understanding
 I Introduction
@@
 The significant value and potential for text-to-SQL have triggered a range of studies on its integration and optimization with LLMs [7, 8, 9, 10]; consequently, LLM-based text-to-SQL remains a highly discussed research field within the NLP and database communities.
-Previous studies have made notable progress in the implementation of text-to-SQL and have undergone a long evolutionary process.
+Previous studies have made notable progress in implementing text-to-SQL and have undergone a long evolutionary process.
 Early efforts were mostly based on well-designed rules and templates [11], specifically suitable for simple database scenarios.
@@
 The format is inspired from [29].
-[…]
-LLM-based Text-to-SQL
-Datasets III-A) Original Datasets & Post-annotated Datasets Cross- domain BIRD [33], DuSQL [34], CoSQL [35], Spider [13], WikiSQL [14], KaggleDBQA [36], ADVETA [37] Spider-$S$ [38], Spider-CG [38], Spider-DK [39], Spider-SYN [40], Spider-Realistic [41], CSpider [42] SParC [43] Knowledge- augmented BIRD [33], SQUALL [44], Spider-DK [39] Cross- lingual DuSQL [34], CSpider [42] Context- dependent CoSQL [35], Spider-$S$ [38], Spider-CG [38], SparC [43] Robustness ADVETA [37], Spider-SYN [40], Spider-Realistic [41] Evaluation Metrics III-B) Content- Matching based Component Matching (CM) [13], Exact Matching (EM) [13] Execution based Execution Accuracy (EX) [13], Valid Efficiency Score (VES) [33] Methods IV) In-context Learning Paradigm IV-A) Zero-shot [7], [33], [27], [45], [46], [47], [48], [49], [9], [50] Few-shot [33], [8], [51], [52], [53], [54], [49], [9], [32], [55] Decomposition Coder-Reviewer [56], DIN-SQL [8], QDecomp [51], C3 [30], MAC-SQL [57], DEA-SQL [58], SGU-SQL [32], MetaSQL [59], PET-SQL [60], PURPLE [61] Prompt Optimization DESEM+P [62], StructGPT [63], SD+SA+Voting [52]RAG+SP&DRC [64], C3 [30], DAIL-SQL [9] ODIS [54], ACT-SQL [49], FUSED [65], DELLM [31] Reasoning Enhancement CoT [51, 33, 9, 32], QDecomp [51], Least-to-Most [51], SQL-PaLM [53], ACT-SQL [49], POT [55] SQL-CRAFT [55], FUXI [66] Execution Refinement MBR-Exec [67], Coder-Reviewer [56], LEVER [68], SELF-DEBUGGING [48], DESEM+P [62] DIN-SQL [8], SD+SA+Voting [52], SQL-PaLM [53], RAG+SP&DRC [64], C3 [30], MAC-SQL [57] DELLM [31], SQL-CRAFT [55], FUXI [66], PET-SQL [60], PURPLE [61] Fine-tuning Paradigm IV-B) Supervised Fine-tuning [45], [9], [50], [53] Enhanced Architecture CLLMs [69] Pre-training CodeS [10] Data Augmentation DAIL-SQL [9], Symbol-LLM [50], CodeS [10], StructLM [70] Decomposition DTS-SQL [71]
+{forest}
+forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-draw, rounded corners, align=left, minimum width=1em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=0pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , [LLM-based Text-to-SQL, leaf-head, ver [ Datasets
+III-A), leaf-datasets, text width=5em [ Original
+Datasets
+Post-annotated
+Datasets, leaf-datasets, text width=6em [ Cross-
+domain, leaf-datasets, text width=5.5em [BIRD [33], DuSQL [34], CoSQL [35], Spider [13], WikiSQL [14], KaggleDBQA [36], ADVETA [37],
+Spider-SS [38], Spider-CG [38], Spider-DK [39], Spider-SYN [40], Spider-Realistic [41], CSpider [42],
+SParC [43], modelnode-datasets, text width=43.5em] ] [ Knowledge-
+augmented, leaf-datasets, text width=5.5em [BIRD [33], SQUALL [44], Spider-DK [39], modelnode-datasets, text width=43.5em] ] [ Cross-
+lingual, leaf-datasets, text width=5.5em [DuSQL [34], CSpider [42], modelnode-datasets, text width=43.5em] ] [ Context-
+dependent, leaf-datasets, text width=5.5em [CoSQL [35], Spider-SS [38], Spider-CG [38], SparC [43], modelnode-datasets, text width=43.5em] ] [ Robustness, leaf-datasets, text width=5.5em [ADVETA [37], Spider-SYN [40], Spider-Realistic [41], modelnode-datasets, text width=43.5em] ] ] ] [ Evaluation
+Metrics
+III-B), leaf-metrics,text width=5em [ Content-
+Matching
+based, leaf-metrics, text width=6em [Component Matching (CM) [13], Exact Matching (EM) [13], modelnode-metrics, text width=50.7em] ] [ Execution
+based, leaf-metrics, text width=6em [Execution Accuracy (EX) [13], Valid Efficiency Score (VES) [33], modelnode-metrics, text width=50.7em] ] ] [ Methods
+IV), leaf-methods,text width=5em [ In-context
+Learning
+Paradigm
+IV-A), leaf-methods, text width=6em [ Trivial
+Prompt, leaf-methods, text width=6.5em [ Zero-shot, modelnode-methods, text width=4.5em [[7], [33], [27], [45], [46], [47], [48], [49], [9], [50], modelnode-methods,text width=36.2em]] [ Few-shot, modelnode-methods, text width=4.5em [[33], [8], [51], [52], [53], [54], [49], [9], [32], [55] , modelnode-methods,text width=36.2em]] ] [ Decomposition, leaf-methods, text width=6.5em [Coder-Reviewer [56], DIN-SQL [8], QDecomp [51], C3 [30], MAC-SQL [57], DEA-SQL [58],
+SGU-SQL [32], MetaSQL [59], PET-SQL [60], PURPLE [61], modelnode-methods,text width=42.5em] ] [ Prompt
+Optimization, leaf-methods, text width=6.5em [DESEM+P [62], StructGPT [63], SD+SA+Voting [52], RAG+SP&DRC [64], C3 [30], DAIL-SQL [9],
+ODIS [54], ACT-SQL [49], FUSED [65], DELLM [31], modelnode-methods,text width=42.5em] ] [ Reasoning
+Enhancement, leaf-methods, text width=6.5em[CoT [51, 33, 9, 32], QDecomp [51], Least-to-Most [51], SQL-PaLM [53], ACT-SQL [49], POT [55],
+SQL-CRAFT [55], FUXI [66], modelnode-methods,text width=42.5em] ] [ Execution
+Refinement, leaf-methods, text width=6.5em[MBR-Exec [67], Coder-Reviewer [56], LEVER [68], SELF-DEBUGGING [48], DESEM+P [62],
+DIN-SQL [8], SD+SA+Voting [52], SQL-PaLM [53], RAG+SP&DRC [64], C3 [30], MAC-SQL [57],
+DELLM [31], SQL-CRAFT [55], FUXI [66], PET-SQL [60], PURPLE [61], modelnode-methods,text width=42.5em] ] ] [ Fine-tuning
+Paradigm
+IV-B), leaf-methods,text width=6em [ Vanilla
+Fine-tuning, leaf-methods, text width=6.5em [[45], [9], [50], [53], modelnode-methods,text width=42.5em] ] [ Enhanced
+Architecture, leaf-methods, text width=6.5em [CLLMs [69], modelnode-methods,text width=42.5em] ] [ Pre-training, leaf-methods, text width=6.5em [CodeS [10], modelnode-methods,text width=42.5em] ] [ Data
+Augmentation, leaf-methods, text width=6.5em [DAIL-SQL [9], Symbol-LLM [50], CodeS [10], StructLM [70], modelnode-methods,text width=42.5em] ] [ Decomposition, leaf-methods, text width=6.5em [DTS-SQL [71], modelnode-methods,text width=42.5em] ] ] ] ]
 Figure 3: Taxonomy tree of the research in LLM-based text-to-SQL. The display order in each node is organized by the released time. The format is adapted from [72].
 Despite the significant progress made in text-to-SQL research, several challenges remain that hinder the development of robust and generalized text-to-SQL systems [73].
+Related works in recent years have surveyed the text-to-SQL systems in deep learning approaches and provided insights into the previous deep neural network and PLM-based research [74, 2, 29].
 In this survey, we aim to catch up with the recent advances and provide a comprehensive review of the current state-of-the-art models and approaches in LLM-based text-to-SQL.
@@
 CoT method are reported in multiple venues: NeurIPS’23 [33], EMNLP’23 [51], VLDB’24 [9], arXiv’24 [32].
-Methods Applied LLMs Benchmark Metrics $\mathbf{C}_{1}$ $\mathbf{C}_{2}$ $\mathbf{C}_{3}$ $\mathbf{C}_{4}$ Release Time Publication Venue
+Methods Applied LLMs Dataset Metrics $\mathbf{C}_{1}$ $\mathbf{C}_{2}$ $\mathbf{C}_{3}$ $\mathbf{C}_{4}$ Release Time Publication Venue
 MBR-Exec [67] Codex [13] EX Apr-2022 EMNLP’22
@@
 The study of few-shot prompting focuses on the number of representations and few-shot instance selection.
-As pilot experiments, few-shot prompting for text-to-SQL are evaluated in multiple datasets with various LLMs [8, 32], achieve better performance rather than zero-shot approaches.
-[33] provides a 1-shot example combining with CoT [102] prompting, the decomposed reasoning steps of the given example trigger the text-to-SQL model for generating accurate SQL.
+As empirical studies, few-shot prompting for text-to-SQL are evaluated in multiple datasets with various LLMs [8, 32], exhibit a solid performance during the comparison with zero-shot prompting.
+[33] provides a 1-shot detailed example to trigger the text-to-SQL model for generating accurate SQL.
 [55] study the effect of the number of few-shot examples.
 [52] focus on the sampling strategies by studying the similarity and the diversity between different demonstrations, setting random sampling as the baseline, and evaluating different strategies and their combination for comparison.
-Furthermore, above the similarity selection, [9] evaluated masked question similarity selection and the upper limit of similarity approaches with various numbers of few-shot examples.
-A study of difficult-level samples selection [51] compared the performance of few-shot Codex, with random selection and difficulty-based selection for few-shot instances on difficulty categorized dataset [13, 41].
+Furthermore, above the similarity-based selection, [9] evaluated masked question similarity selection and the upper limit of similarity approaches with various numbers of few-shot examples.
+A study of difficult-level samples selection [51] compared the performance of few-shot Codex [100], with random selection and difficulty-based selection for few-shot instances on difficulty categorized dataset [13, 41].
 Three difficulty-based selection strategies are devised based on the number of selected samples at different difficulty levels.
 [49] utilize a hybrid strategy for selecting samples, which combines static examples and similarity-based dynamic examples for few-shot prompting.
-In their evaluations, they also test the impact of different input schema styles and different static and dynamic exemplar numbers.
+In their settings, they also evaluate the impact of different input schema styles and various static and dynamic exemplar numbers.
 The impact of cross-domain few-shot examples is also being studied [54].
-When incorporating in-domain and out-of-domain with different numbers of examples, the in-domain demonstration always outperforms zero-shot and out-of-domain examples and gets better as the number of examples rises.
+When incorporating in-domain and out-of-domain examples with different numbers, the in-domain demonstration outperforms zero-shot and out-of-domain examples, which will also perform better as the number of examples rises.
 To explore the detailed construction of input prompt, [53] compare the concise and verbose prompt design approaches.
@@
 $\mathbf{C}_{1}$-Decomposition
-As an intuitive solution, decomposing a challenging user question into simpler sub-questions and using multi-step reasoning for implementation can reduce the complexity of the full text-to-SQL task [51, 8].
-Dealing with less complexity, LLMs have the potential to have better performance.
-The decomposition approaches for LLM-based text-to-SQL are categorized into two paradigms: (1) sub-task decomposing, provides additional parsing to assist the final SQL generation by decomposing the overall text-to-SQL task into smaller effective sub-tasks (e.g., schema linking, domain classification).
-(2) sub-question decomposing, divides the user question into sub-questions to reduce the question’s complexity and difficulty, then generates the sub-SQL by solving these questions to deduce the final SQL query.
+As an intuitive solution, decomposing a challenging user question into simpler sub-questions or using multi-components for implementation can reduce the complexity of the overall text-to-SQL task [51, 8].
+Dealing with less complexity, LLMs have the potential to elicit more accurate SQL generation.
+The decomposition-based approaches for LLM-based text-to-SQL are categorized into two paradigms: (1) sub-task decomposition, provides additional parsing to assist the final SQL generation by decomposing the overall text-to-SQL task into more manageable effective sub-tasks (e.g., schema linking [71], domain classification [54]).
+(2) sub-question decomposition, divides the user question into sub-questions to reduce the question’s complexity and difficulty, then generates the sub-SQL by solving these questions to deduce the final SQL query.
+The technical novelty of the decomposition paradigm
 DIN-SQL [8] proposed a decomposed in-context learning method consisting of four modules: schema linking, classification & decomposition, SQL generation, and self-correction.
-DIN-SQL first finishes the schema linking between the user question and the target database; the following module decomposes the user question into correlated sub-questions and does a difficulty classification.
+DIN-SQL first generates the schema linking between the user question and the target database; the following module decomposes the user question into correlated sub-questions and does a difficulty classification.
 Based on the above information, the SQL generation module generates a corresponding SQL, and the self-correction module identifies and corrects the potential errors in the predicted SQL.
-This approach comprehensively considers the decomposition of both sub-tasks and sub-questions.
+This approach comprehensively considers the sub-questions decomposition as a module of sub-tasks decomposition.
 Coder-Reviewer [56] framework proposed a re-ranking method, combining Coder models for the generation and Reviewer models to evaluate the likelihood of the instruction.
@@
 This method decomposes the overall task, enabling the SQL generation module to have the corresponding prerequisite (information determination, question classification) and subsequent (self-correction, active learning) sub-tasks.
-Through the workflow paradigm, an accurate SQL query is generated.
+The workflow paradigm allows the LLM to generate more accurate SQL queries.
 SGU-SQL [32] is a structure-to-SQL framework, leveraging the inherent structure information to assist SQL generation.
@@
 A meta operator decomposes the user question with a grammar tree and finally designs the input prompt with meta-operation in SQL.
-MetaSQL [59] introduces a three-stage approach for SQL generation, which consists of decomposition, generation, and rank.
+MetaSQL [59] introduces a three-stage approach for SQL generation: decomposition, generation, and ranking.
 The decomposition stage uses semantic decomposition and metadata composition to process the user question.
@@
 As previously introduced, few-shot learning is widely studied for prompting LLMs [85].
-For LLM-based text-to-SQL with in-context learning, trivial few-shot approaches obtained promising results [33, 8, 9], further optimization of few-shot prompting has the potential to achieve higher performance.
-Since the performance of SQL generation in off-the-shelf LLMs largely depends on the quality of the corresponding input prompt [106], many decisive factors that can influence the quality of the prompt have become focuses of the research [9] (e.g., quality and quantity in the few-shot organization, the similarity between user questions and few-shots instances, external knowledge/hints).
-The process of prompt quality improvement is actually the prompt’s optimization, including few-shot sampling strategies, schema augmentation, and external knowledge generation.
+For LLM-based text-to-SQL with in-context learning, trivial few-shot approaches obtained promising results [33, 8, 9], further optimization of few-shot prompting has the potential to elicit better performance.
+Since the accuracy of SQL generation in off-the-shelf LLMs largely depends on the quality of the corresponding input prompt [107], many decisive factors that can influence the quality of the prompt have become points of focus of the current research [9] (e.g., quality and quantity in the few-shot organization, the similarity between user questions and few-shots instances, external knowledge/hints) The process of improving prompt quality can be summarized as the prompt’s optimization, including advanced few-shot sampling strategies, schema information augmentation, and external knowledge integration.
 DESEM [62] is a prompt engineering framework with de-semanticization and skeleton retrieval.
-The framework first employs domain-specific words masking module to remove the semantic tokens in questions that preserve the question’s intentions.
-And then utilizes an adjustable prompting module that retrieves the few-shot examples with identical question intentions and incorporates schema-relevance filtering to guide the LLM’s SQL generation.
+The framework first employs domain-specific words masking module to remove the semantic tokens in user questions that preserve the intentions.
+It then utilizes an adjustable prompting module that retrieves the few-shot examples with identical question intentions and incorporates schema-relevance filtering to guide the LLM’s SQL generation.
 The QDecomp [51] framework introduces the InterCOL mechanism to incrementally incorporate the decomposed sub-questions with correlative table and column names.
@@
 LLMs have exhibited promising capabilities in tasks involving commonsense reasoning, symbolic reasoning, and arithmetic reasoning [108].
-Since for the text-to-SQL tasks, numeric and synonym reasoning frequently occur in realistic scenarios [33, 41], the prompting strategies for the LLMs reasoning possess the potential to enhance their SQL generation capabilities.
-Recent studies primarily focus on incorporating well-designed reasoning-enhanced methods for text-to-SQL adaptation, improving LLMs to address the challenge about complex questions that require multi-step reasoning and the issue of self-consistency [108] in SQL generation.
+For text-to-SQL tasks, numeric and synonym reasoning frequently occur in realistic scenarios [33, 41].
+The prompting strategies for LLMs’ reasoning have the potential to enhance SQL generation capabilities.
+Recent studies primarily focus on integrating well-designed reasoning-enhanced methods for text-to-SQL adaptation, improving LLMs to address the challenge about complex questions that require sophisticated reasoning [^3] The difference between multi-step reasoning (e.g., Chain of Thought) and the decomposition paradigm is that the former research focuses on advancing inherent reasoning within a single-turn generation, whereas the latter study involves using different components to assist the final generation through multiple calls of LLMs.
+and the self-consistency in SQL generation.
 Chain-of-Thoughts (CoT) prompting technique [103] involves a comprehensive reasoning process that guides LLMs towards accurate deduction, eliciting reasoning in LLMs.
@@
 $\mathbf{C}_{4}$-Execution Refinement
-To design criteria for accurate SQL generation, whether a generated SQL can be successfully executed and elicit a correct answer for the user question is always the priority [13].
-As a complex programming task, generating the correct SQL in one go becomes challenging.
+When designing criteria for accurate SQL generation, the priority is always whether a generated SQL can be successfully executed and retrieve content to correctly answer the user question [13].
+As a complex programming task, generating the correct SQL in one go is challenging.
 Intuitively, considering the execution feedback/results in SQL generation assists the alignment to the corresponding database environment, which allows the LLMs to gather the potential executed errors and results to refine the generated SQL or hold a majority vote [30].
-The execution-aware methods in text-to-SQL incorporate the execution feedback in two main approaches: 1) Incorporating the feedback through second round prompting for regeneration, for every SQL query generated in the initial response, it will be executed in the corresponding database, thus obtaining feedback from the database.
+The execution-aware methods in text-to-SQL incorporate the execution feedback in two main approaches: 1) Incorporating the feedback through second round prompting for re-generation, for every SQL query generated in the initial response, it will be executed in the corresponding database, thus obtaining the database’s feedback.
 This feedback might be an error, or it might yield results that will be appended to the second round prompt.
-Through in-context learning of this feedback, LLMs are able to refine or regenerate the original SQL, thereby enhancing accuracy.
-2) Utilize execution-based selection strategies for generated SQL, sample multiple generated SQL queries from LLM), and execute each in the database.
-Based on the results of each SQL query, use selection strategies (e.g., self-consistency, majority vote [60]) to define a query from the SQL set that satisfies the criteria as the final predicted SQL.
+Through in-context learning of this feedback, LLMs are able to refine or re-generate the original SQL, thereby enhancing accuracy.
+2) Utilize execution-based selection strategies for the generated SQL, sample multiple generated SQL queries from LLM, and execute each in the database.
+Based on the execution results of each SQL query, use selection strategies (e.g., self-consistency, majority vote [60]) to define one SQL query from the SQL set that satisfies the criteria as the final predicted SQL.
 MRC-EXEC [67] introduced a natural language to code (NL2Code) translation framework with execution, which executes each sampled SQL query and selects the example with the minimal execution result–based Bayes risk [111].
@@
 Decomposition DTS-SQL [71] Mistral* [13, 40] Feb-2024 arXiv’24
-Since supervised fine-tuning (SFT) is the mainstream approach in the LLMs training [29], for open-source LLMs (e.g., LLaMA-2 [93], Gemma [112]), the most straightforward method to enable the model to adapt a specific domain quickly is to use collected domain label to perform SFT on the model.
+Since supervised fine-tuning (SFT) is the mainstream approach in the LLMs training [29, 91], for open-source LLMs (e.g., LLaMA-2 [94], Gemma [113]), the most straightforward method to enable the model to adapt a specific domain quickly is to use collected domain label to perform SFT on the model.
 The SFT phase is typically the preliminary phase of the well-designed training framework [112, 114], as well as the fine-tuning of text-to-SQL.
@@
 where $\hat{y}_{k}$ is the $k^{th}$ token of the generated SQL query $\hat{Y}$, and $Y$ is the corresponding ground-truth label.
-The SFT approach for text-to-SQL has been widely adopted in text-to-SQL research for various open-source LLMs [10, 46, 9].
+The SFT approach, which is also a vanilla fine-tuning approach for text-to-SQL has been widely adopted in text-to-SQL research for various open-source LLMs [10, 46, 9].
 Compared to in-context learning (ICL) approaches, fine-tuning paradigms are more inclined to be at a starting point in LLM-based text-to-SQL.
@@
 The pre-training phase of the CodeS [10] model consists of three stages of incremental pre-training.
-Starting from a basic code-specific LLM [122], CodeS are further pre-trained on a hybrid training corpus, including SQL-related data, NL-to-Code data, and NL-related data.
+Starting from a basic code-specific LLM [122], CodeS are incre pre-trained on a hybrid training corpus, including SQL-related data, NL-to-Code data, and NL-related data.
 The text-to-SQL understanding and performance are significantly improved.
@@
 From another perspective, we previously discussed that text-to-SQL can enhance LLM-based question-answering (QA) by providing factual information.
-The database can store relational knowledge as structural information, and the structure-based QA can potentially benefit from text-to-SQL (e.g., knowledge-based question-answering, KBQA [133]).
+The database can store relational knowledge as structural information, and the structure-based QA can potentially benefit from text-to-SQL (e.g., knowledge-based question-answering, KBQA [133, 134]).
 Construct the factual knowledge with database structure, and then incorporate the text-to-SQL system to achieve information retrieval, which can potentially assist further QA with more accurate factual knowledge [135].

### v4

@@
 Abstract
-Generating accurate SQL from natural language questions (text-to-SQL) is a long-standing challenge due to the complexities in user question understanding, database schema comprehension, and SQL generation.
-Conventional text-to-SQL systems, comprising human engineering and deep neural networks, have made substantial progress.
-Subsequently, pre-trained language models (PLMs) have been developed and utilized for text-to-SQL tasks, achieving promising performance.
-As modern databases become more complex, the corresponding user questions also grow more challenging, causing PLMs with parameter constraints to produce incorrect SQL.
-This necessitates more sophisticated and tailored optimization methods, which, in turn, restricts the applications of PLM-based systems.
-Recently, large language models (LLMs) have demonstrated significant capabilities in natural language understanding as the model scale increases.
-Therefore, integrating LLM-based implementation can bring unique opportunities, improvements, and solutions to text-to-SQL research.
-In this survey, we present a comprehensive review of LLM-based text-to-SQL.
-Specifically, we propose a brief overview of the technical challenges and the evolutionary process of text-to-SQL.
-Then, we provide a detailed introduction to the datasets and metrics designed to evaluate text-to-SQL systems.
-After that, we present a systematic analysis of recent advances in LLM-based text-to-SQL.
-Finally, we discuss the remaining challenges in this field and propose expectations for future research directions.
+Generating accurate SQL from users’ natural language questions (text-to-SQL) remains a long-standing challenge due to the complexities involved in user question understanding, database schema comprehension, and SQL generation.
+Traditional text-to-SQL systems, which combine human engineering and deep neural networks, have made significant progress.
+Subsequently, pre-trained language models (PLMs) have been developed for text-to-SQL tasks, achieving promising results.
+However, as modern databases and user questions grow more complex, PLMs with a limited parameter size often produce incorrect SQL.
+This necessitates more sophisticated and tailored optimization methods, which restricts the application of PLM-based systems.
+Recently, large language models (LLMs) have shown significant capabilities in natural language understanding as model scale increases.
+Thus, integrating LLM-based solutions can bring unique opportunities, improvements, and solutions to text-to-SQL research.
+In this survey, we provide a comprehensive review of existing LLM-based text-to-SQL studies.
+Specifically, we offer a brief overview of the technical challenges and evolutionary process of text-to-SQL.
+Next, we introduce the datasets and metrics designed to evaluate text-to-SQL systems.
+Subsequently, we present a systematic analysis of recent advances in LLM-based text-to-SQL.
+Finally, we make a summarization and discuss the remaining challenges in this field and suggest expectations for future research directions.
 Index Terms:
@@
 I Introduction
-Figure 1: An example for LLM-based text-to-SQL selected from the BIRD dataset.
-A user proposes a question about football leagues.
-The LLM takes the question and the schema of its corresponding database as the input and then generates an SQL query as the output.
-The SQL query can be executed in the database and retrieve the content “The 5 leagues with the highest matches” to answer the user question.
-Text-to-SQL is a long-standing task in natural language processing research.
-It aims to convert (translate) natural language questions into database-executable SQL queries.
+Figure 1: An example of LLM-based text-to-SQL is selected from the BIRD [1] dataset.
+A user asks a question about football leagues.
+The LLM takes this question along with the schema of the corresponding database as input and generates an SQL query as output.
+The generated SQL query can be executed in the database, retrieves the content “The 5 leagues with the highest matches”, providing the answer to the user’s question.
+Text-to-SQL is a long-standing task in natural language processing, aiming to convert natural language (NL) questions into database-executable SQL queries.
 Fig.
-1 provides an example of a large language model-based (LLM-based) text-to-SQL system.
-Given a user question such as “Could you tell me the names of the 5 leagues with the highest matches of all time and how many matches were played in the said league?”, the LLM takes the question and its corresponding database schema as input and then generates an SQL query as output.
-This SQL query can be executed in the database to retrieve the relevant content to answer the user’s question.
-The above system builds a natural language interface to the database (NLIDB) with LLMs.
-Since SQL remains one of the most widely used programming languages, with over half (51.52%) of professional developers using SQL in their work, it is notable that only around a third (35.29%) of those developers are systematically trained [^1] https://survey.stackoverflow.co/2023, the NLIDB enables non-skilled users to access structured databases like professional database engineers [1, 2] and also accelerates human-computer interaction [3].
-Furthermore, amid the research hotspot of LLMs, text-to-SQL can provide a potential solution to the prevalent hallucination [4, 5] issue by incorporating realistic content from the database to fill the knowledge gaps of LLMs [6].
-The significant value and potential for text-to-SQL have triggered a range of studies on its integration and optimization with LLMs [7, 8, 9, 10]; consequently, LLM-based text-to-SQL remains a highly discussed research field within the NLP and database communities.
-Previous studies have made notable progress in implementing text-to-SQL and have undergone a long evolutionary process.
+1 illustrates an example of a large language model-based (LLM-based) text-to-SQL system.
+Given a user question such as “Could you tell me the names of the 5 leagues with the highest matches of all time and how many matches were played in the said league?”, the LLM takes the question and its corresponding database schema as input to generate an SQL query.
+Executing this SQL query in the database retrieves the relevant content needed to answer the user’s question.
+This system serves as a natural language interface to the database (NLIDB) using LLMs.
+Since SQL is one of the most widely used programming languages, with over half (51.52%) of professional developers employing it, but only around a third (35.29%) are systematically trained, [^1] https://survey.stackoverflow.co/2023 the NLIDB allows non-skilled users to access structured databases like professional database engineers [2, 3] and enhances human-computer interaction [4].
+Furthermore, amid the research hotspot of LLMs, text-to-SQL can potentially mitigate prevalent hallucination issues by integrating realistic content from databases to fill LLMs’ knowledge gaps [5].
+The significant value of text-to-SQL have triggered a range of studies on its integration and optimization with LLMs [6, 7, 8]; consequently, LLM-based text-to-SQL remains a highly discussed research field in the NLP and database communities.
+Previous traditional methods have made notable progress in implementing text-to-SQL.
+As shown in Fig.
+2, these implementations have undergone a long evolutionary process.
 Early efforts were mostly based on well-designed rules and templates [9], specifically suitable for simple database scenarios.
-In recent years, with the heavy labor costs [12] brought by rule-based methods and the growing complexity of database environments [13, 14, 15], designing a rule or template for each scenario has become increasingly difficult and impractical.
+In recent years, due to the heavy labor costs [10] associated with rule-based methods and the growing complexity of database environments [11, 12, 13], designing a rule or template for each scenario has become increasingly difficult and impractical.
 The development of deep neural networks has advanced the progress of text-to-SQL [14, 15], which can automatically learn a mapping from the user question to its corresponding SQL [16, 17].
-Subsequently, pre-trained language models (PLMs) with strong semantic parsing capacity have become the new paradigm for text-to-SQL systems [20], boosting their performance to a new level [21, 22, 23].
+Subsequently, pre-trained language models (PLMs) with strong semantic parsing capabilities have become the new paradigm for text-to-SQL systems [18], boosting their performance to a new level [19, 20, 21].
 Incremental research on PLM-based optimization, such as table content encoding [18, 17, 22] and pre-training [18, 23], has further advanced this field.
-Most recently, the LLM-based approaches implementing text-to-SQL through in-context learning (ICL) [8] and fine-tuning (FT) [10] paradigm, reaching state-of-the-art accuracy with the well-designed framework and stronger comprehension capability compared to PLMs.
-The overall implementation details of LLM-based text-to-SQL can be divided into 3 aspects: 1.
+Most recently, LLM-based approaches have implemented text-to-SQL through the in-context learning [7] and fine-tuning [24] paradigms, reaching state-of-the-art accuracy with well-designed frameworks and stronger comprehension capabilities compared to PLM-based methods.
+The overall implementation details of LLM-based text-to-SQL can be divided into three aspects: 1.
 Question understanding: The NL question is a semantic representation of the user’s intention, which the corresponding generated SQL query is expected to align with; 2.
@@
 SQL generation: This involves incorporating the above parsing and then predicting the correct syntax to generate executable SQL queries that can retrieve the required answer.
-The LLMs have proven to perform a good vanilla implementation [7, 27], benefiting from the more powerful semantic parsing capacity enabled by the richer training corpus [28, 29].
-Further studies on enhancing the LLMs for question understanding [8, 9], schema comprehension [30, 31], and SQL generation [32] are being increasingly released.
-Figure 2: A sketch of the evolutionary process for text-to-SQL research from the perspective of implementation paradigm.
-Each stage is presented with two representative implementation techniques.
-The timestamps for the stages are not exactly accurate; we set each timestamp according to the release time of the representative works of each implementation paradigm, with a margin of error of about one year before and after.
-The format is inspired from [29].
+LLMs have proven to perform a successful vanilla implementation [6, 25], benefiting from the powerful semantic parsing capacity enabled by rich training corpus [26, 27].
+Further studies on enhancing LLMs for question understanding [7, 8], schema comprehension [28, 29], and SQL generation [30] are being increasingly released.
+Figure 2: A sketch of text-to-SQL research evolution outlines the advancement of implementation paradigms, each represented with different color backgrounds.
+Before 2023, the focus is on a selection of representative traditional studies.
+However, from 2023 onward, the emphasis shifts to the rapid advancements driven by LLMs, marking a significant acceleration in the field.
 {forest}
@@
 SParC [40], modelnode-datasets, text width=43.5em] ] [ Knowledge-
-augmented, leaf-datasets, text width=5.5em [BIRD [33], SQUALL [44], Spider-DK [39], modelnode-datasets, text width=43.5em] ] [ Cross-
-lingual, leaf-datasets, text width=5.5em [DuSQL [34], CSpider [42], modelnode-datasets, text width=43.5em] ] [ Context-
-dependent, leaf-datasets, text width=5.5em [CoSQL [35], Spider-SS [38], Spider-CG [38], SparC [43], modelnode-datasets, text width=43.5em] ] [ Robustness, leaf-datasets, text width=5.5em [ADVETA [37], Spider-SYN [40], Spider-Realistic [41], modelnode-datasets, text width=43.5em] ] ] ] [ Evaluation
+augmented, leaf-datasets, text width=5.5em [BIRD [1], SQUALL [41], Spider-DK [36], BIRD-CRITIC [1], Spider 2.0 [42], modelnode-datasets, text width=43.5em] ] [ Cross-
+lingual, leaf-datasets, text width=5.5em [DuSQL [31], CSpider [39], Spider-Vietnamese [43], modelnode-datasets, text width=43.5em] ] [ Context-
+dependent, leaf-datasets, text width=5.5em [CoSQL [32], Spider-SS [35], Spider-CG [35], SparC [40], modelnode-datasets, text width=43.5em] ] [ Robustness, leaf-datasets, text width=5.5em [ADVETA [34], Spider-SYN [37], Spider-Realistic [38], Dr.
+Spider [44], modelnode-datasets, text width=43.5em] ] [ Long-context, leaf-datasets, text width=5.5em [BIRD-CRITIC [1], Spider 2.0 [42], modelnode-datasets, text width=43.5em] ] [ Specialized-
+domain, leaf-datasets, text width=5.5em [BULL [45], modelnode-datasets, text width=43.5em] ] ] ] [ Evaluation
 Metrics
@@
 Paradigm
-IV-A), leaf-methods, text width=6em [ Trivial
-Prompt, leaf-methods, text width=6.5em [ Zero-shot, modelnode-methods, text width=4.5em [[7], [33], [27], [45], [46], [47], [48], [49], [9], [50], modelnode-methods,text width=36.2em]] [ Few-shot, modelnode-methods, text width=4.5em [[33], [8], [51], [52], [53], [54], [49], [9], [32], [55] , modelnode-methods,text width=36.2em]] ] [ Decomposition, leaf-methods, text width=6.5em [Coder-Reviewer [56], DIN-SQL [8], QDecomp [51], C3 [30], MAC-SQL [57], DEA-SQL [58],
-SGU-SQL [32], MetaSQL [59], PET-SQL [60], PURPLE [61], modelnode-methods,text width=42.5em] ] [ Prompt
+IV-A), leaf-methods, text width=6em [ Decomposition, leaf-methods, text width=6.5em [Coder-Reviewer [46], DIN-SQL [7], QDecomp [47], C3 [28], MAC-SQL [48], DEA-SQL [49],
+SGU-SQL [30], MetaSQL [50], $\textit{{R}}^{3}$ [51], PET-SQL [52], PURPLE [53], TA-SQL [54], MCS-SQL [55],
+CHESS [56], SuperSQL [57], Distillery [58], RSL-SQL [59], Gen-SQL [60], ReFoRCE [61], modelnode-methods,text width=42.5em] ] [ Prompt
 Optimization, leaf-methods, text width=6.5em [DESEM+P [62], StructGPT [63], SD+SA+Voting [64], RAG+SP&DRC [65], C3 [28], DAIL-SQL [8],
-ODIS [54], ACT-SQL [49], FUSED [65], DELLM [31], modelnode-methods,text width=42.5em] ] [ Reasoning
+ODIS [66], ACT-SQL [67], FUSED [68], DELLM [69], Dubo-SQL [70], TA-SQL [54],
+MCS-SQL [55], SuperSQL [57], Distillery [58], RSL-SQL [59], ICRL [71], Gen-SQL [60],
+SAFE-SQL [72], ReFoRCE [61], modelnode-methods,text width=42.5em] ] [ Reasoning
 Enhancement, leaf-methods, text width=6.5em[CoT [47, 1, 8, 30], QDecomp [47], Least-to-Most [47], SQL-PaLM [73], ACT-SQL [67], POT [74],
-SQL-CRAFT [55], FUXI [66], modelnode-methods,text width=42.5em] ] [ Execution
+SQL-CRAFT [74], FUXI [75], SuperSQL [57], CHASE-SQL [76], ReFoRCE [61], modelnode-methods,text width=42.5em] ] [ Execution
 Refinement, leaf-methods, text width=6.5em[MBR-Exec [77], Coder-Reviewer [46], LEVER [78], SELF-DEBUGGING [79], DESEM+P [62],
 DIN-SQL [7], SD+SA+Voting [64], SQL-PaLM [73], RAG+SP&DRC [65], C3 [28], MAC-SQL [48],
-DELLM [31], SQL-CRAFT [55], FUXI [66], PET-SQL [60], PURPLE [61], modelnode-methods,text width=42.5em] ] ] [ Fine-tuning
+DELLM [69], SQL-CRAFT [74], FUXI [75], $\textit{{R}}^{3}$ [51], PET-SQL [52], PURPLE [53], Dubo-SQL [70],
+MCS-SQL [55], CHESS [56], Distillery [58], E-SQL [80], CHASE-SQL [76], RSL-SQL [59], modelnode-methods,text width=42.5em] ] ] [ Fine-tuning
 Paradigm
-IV-B), leaf-methods,text width=6em [ Vanilla
-Fine-tuning, leaf-methods, text width=6.5em [[45], [9], [50], [53], modelnode-methods,text width=42.5em] ] [ Enhanced
+IV-B), leaf-methods,text width=6em [ Enhanced
 Architecture, leaf-methods, text width=6.5em [CLLMs [81], modelnode-methods,text width=42.5em] ] [ Pre-training, leaf-methods, text width=6.5em [CodeS [24], modelnode-methods,text width=42.5em] ] [ Data
-Augmentation, leaf-methods, text width=6.5em [DAIL-SQL [9], Symbol-LLM [50], CodeS [10], StructLM [70], modelnode-methods,text width=42.5em] ] [ Decomposition, leaf-methods, text width=6.5em [DTS-SQL [71], modelnode-methods,text width=42.5em] ] ] ] ]
-Figure 3: Taxonomy tree of the research in LLM-based text-to-SQL. The display order in each node is organized by the released time. The format is adapted from [72].
+Augmentation, leaf-methods, text width=6.5em [DAIL-SQL [8], Symbol-LLM [82], CodeS [24], StructLM [83], Dubo-SQL [70], Distillery [58],
+XiYan-SQL [84], modelnode-methods,text width=42.5em] ] [ Multi-task
+Tuning, leaf-methods, text width=6.5em [SQL-LLaMA [85], DTS-SQL [86], KaSLA [29], MSc-SQL [87], XiYan-SQL [84], ROUTE [88], modelnode-methods,text width=42.5em] ] ] ] ]
+Figure 3: Taxonomy tree of research in LLM-based text-to-SQL, taking inspiration from [89]; nodes are ordered by release time.
 Despite the significant progress made in text-to-SQL research, several challenges remain that hinder the development of robust and generalized text-to-SQL systems [90].
-Related works in recent years have surveyed the text-to-SQL systems in deep learning approaches and provided insights into the previous deep neural network and PLM-based research [74, 2, 29].
-In this survey, we aim to catch up with the recent advances and provide a comprehensive review of the current state-of-the-art models and approaches in LLM-based text-to-SQL.
+Recent related works have surveyed text-to-SQL systems using deep learning approaches and provided insights into previous deep neural network and PLM-based research [91, 3, 27, 92].
+In this survey, we aim to catch up with recent advances and provide a comprehensive review of the current state-of-the-art models and approaches in LLM-based text-to-SQL.
 We begin by introducing the fundamental concepts and challenges associated with text-to-SQL, highlighting the importance of this task in various domains.
@@
 After the overview, we provide a detailed introduction and analysis of the recent advances in text-to-SQL integrating LLMs.
-Specifically, the body of our survey covers a range of contents related to LLM-based text-to-SQL, including:
-  • Datasets and Benchmarks: We provide a detailed introduction to the commonly used datasets and benchmarks for evaluating LLM-based text-to-SQL systems.
-We discuss their characteristics, complexity, and the challenges they pose for text-to-SQL development and evaluation.
-  • Evaluation Metrics: We present the evaluation metrics used to assess the performance of LLM-based text-to-SQL systems, including both content matching-based and execution-based paradigms.
-We then briefly introduce the characteristics of each metric.
-  • Methods and Models: We present a systematic analysis of the different methods and models employed for LLM-based text-to-SQL, including in-context learning and fine-tuning-based paradigms.
-We discuss their implementation details, strengths, and adaptations specific to the text-to-SQL task from various implementation perspectives.
-  • Expectations and Future Directions: We discuss the remaining challenges and limitations of LLM-based text-to-SQL, such as real-world robustness, computational efficiency, data privacy, and extensions.
-We also outline potential future research directions and opportunities for improvement and optimization.
-We hope this survey provides a clear overview of recent studies and inspires future research. Fig. 3 shows a taxonomy tree that summarizes the structure and contents of our survey.
+Specifically, the body of our survey covers a range of topics related to LLM-based text-to-SQL, including:
+  • Benchmarks and Metrics: We introduce the commonly used datasets and benchmarks for evaluating LLM-based text-to-SQL systems.
+We discuss their characteristics, complexity, and the challenges they pose for development and evaluation, along with the evaluation metrics.
+  • Methods: We systematically analyze different methods proposed for LLM-based text-to-SQL, including in-context learning and fine-tuning paradigms.
+We explore their implementation details, strengths, and specific adaptations, organized through well-designed categorization.
+  • Summarization and Expectations: We summarize the advantages and limitations of both LLM-based and traditional methods, as well as within LLM-based methods themselves.
+We discuss remaining challenges such as real-world robustness, computational efficiency, data privacy, and possible extensions.
+Additionally, we outline future research directions and opportunities for improvement.
+To summarize our technical contributions, at the time of this paper’s release, we present the first LLM-based text-to-SQL survey in the community, establishing a systematic taxonomy for LLM-based text-to-SQL methods.
+This categorization provides analysis from technical implementation perspectives, clearly indicating their contributions to generating better SQL.
+Then, we further integrate insights while analyzing the trade-offs between LLM-based and traditional methods, as well as between in-context learning and fine-tuning paradigms for future developments.
+Finally, we highlight the remaining challenges in robustness, real-world deployment, efficiency, and domain adaptation, offering actionable directions for future research.
+Fig.
+3 shows a taxonomy tree that summarizes the structure and contents of our survey.
 II Overview
-Text-to-SQL is a task that aims to convert natural language questions into corresponding SQL queries that can be executed in a relational database.
-Formally, given a user question $\mathcal{Q}$ (also known as a user query, natural language question, etc.) and a database schema $\mathcal{S}$, the goal of the task is to generate an SQL query $Y$ that retrieves the required content from the database to answer the user question.
-Text-to-SQL has the potential to democratize access to data by allowing users to interact with databases using natural language without the need for specialized knowledge of SQL programming [75].
-This can benefit various domains, such as business intelligence, customer support, and scientific research, by enabling non-skilled users to easily retrieve target content from databases and facilitating more efficient data analysis.
+Text-to-SQL is a task that aims to convert NL questions into SQL queries that can be executed in a relational database.
+It has the potential to democratize data access by allowing users to interact with databases using natural language, thus eliminating the need for specialized SQL programming knowledge [93].
+This capability can greatly benefit domains such as business intelligence, customer support, and scientific research, by enabling non-specialized users to easily retrieve target content from databases and facilitating more efficient data analysis.
 II-A Challenges in Text-to-SQL
-The technical challenges for text-to-SQL implementations can be summarized as follows:
+The technical challenges are summarized as follows:
 II-A1 Linguistic Complexity and Ambiguity
-Natural language questions often contain complex linguistic representations, such as nested clauses, coreferences, and ellipses, which make it challenging to map them accurately to the corresponding part of SQL queries [41].
+Natural language questions often contain complex linguistic representations, such as nested clauses, coreferences, and ellipses, which make it challenging to map them accurately to the corresponding parts of SQL queries [38].
 Additionally, natural language is inherently ambiguous, with multiple possible representations for a given user question [94, 95].
@@
 II-A2 Schema Understanding and Representation
-To generate accurate SQL queries, text-to-SQL systems need to have a comprehensive understanding of the database schema, including table names, column names, and relationships between various tables.
-However, the database schema can be complex and vary significantly across different domains [13].
-Representing and encoding the schema information in a way that can be effectively utilized by the text-to-SQL model is a challenging task.
+To generate accurate SQL queries, text-to-SQL systems must have a comprehensive understanding of the database schema, including table names, column names, and relationships between various tables.
+Representing and encoding schema information is challenging because the database schema is often complex and varies significantly across different domains [11].
 II-A3 Rare and Complex SQL Operations
@@
 These operations are less frequent in the training data and pose challenges for text-to-SQL systems to generate accurately.
-Designing models that can be generalized to a wide range of SQL operations, including rare and complex scenarios, is an essential consideration.
+Designing models that can generalize to a wide range of SQL operations, including rare and complex scenarios, is an essential consideration.
 II-A4 Cross-Domain Generalization
 Text-to-SQL systems often struggle to generalize across various database scenarios and domains.
-Models trained on a specific domain may not perform well on the proposed questions from other domains due to the variety in vocabulary, database schema structure, and question patterns.
-Developing systems that can effectively generalized to new domains with minimal domain-specific training data or fine-tuning adaptation is a significant challenge [78].
+Models trained on a specific domain may not perform well on questions from other domains due to variations in vocabulary, database schema structure, and question patterns.
+Developing systems that can effectively generalize to new domains with minimal domain-specific training data or fine-tuning adaptation is a significant challenge [96].
 II-B Evolutionary Process
@@
 Early text-to-SQL systems relied heavily on rule-based methods [9, 10, 23], where manually crafted rules and heuristics were used to map natural language questions to SQL queries.
-These approaches often involved extensive feature engineering and domain-specific knowledge.
-While rule-based methods achieved success in specific simple domains, they lacked the flexibility and generalization capabilities needed to handle diverse and complex questions.
+Rule-based systems excel at generating syntactically correct SQL queries by relying on predefined templates and strict grammatical rules, ensuring adherence to SQL syntax and structure.
+Their deterministic nature guarantees consistent outputs for specific input patterns, minimizing variability in query generation.
+However, these methods struggle to interpret linguistically complex or ambiguous natural language questions, such as those involving nested clauses, coreferences, or ellipses, and often fail to map these elements to database schema components like tables or columns.
+Their reliance on manual rules limits generalization to unseen question patterns or schema structures, requiring frequent updates for new domains.
+Additionally, encoding complex schemas with multi-table relationships into usable rules is labor-intensive and prone to errors.
 II-B2 Deep Learning-based Approaches
@@
 Recently, researchers introduced graph neural networks (GNNs) for text-to-SQL tasks by leveraging schema dependency graphs to capture the relationships between database elements [16, 99].
+Despite their flexibility, these methods often generate incorrect or incomplete SQL syntax, such as missing clauses or improper nesting, due to limited structural awareness.
+They also struggle with rare or complex SQL operations, like nested subqueries and window functions, because such constructs are underrepresented in training data.
 II-B3 PLM-based Implementation
-Pre-trained language models (PLMs) have emerged as a powerful solution for text-to-SQL, leveraging the vast amounts of linguistic knowledge and semantic understanding captured during the pre-training process.
+PLMs have emerged as a powerful solution for text-to-SQL, leveraging the vast amounts of linguistic knowledge and semantic understanding captured during the pre-training process.
 The early adoption of PLMs in text-to-SQL primarily focused on fine-tuning off-the-shelf PLMs, such as BERT [100] and RoBERTa [101], on standard text-to-SQL datasets [11, 12].
@@
 Schema-aware PLMs are designed to capture the relationships and constraints present in the database structure [19].
+While PLMs generate more accurate SQL than deep learning-based methods, they still face difficulties with complex operations like outer joins or aggregations, often producing syntactically flawed queries.
+Besides that, cross-domain performance drops significantly when schema structures or vocabulary differ from training data, necessitating resource-intensive fine-tuning.
 II-B4 LLM-based Implementation
-Large language models (LLMs), such as GPT series [84, 85, 86], have gained significant attention in recent years due to their capability to generate coherent and fluent text.
-Researchers have started exploring the potential of LLMs for text-to-SQL by leveraging their extensive knowledge reserves and superior generation capabilities [7, 9].
-These approaches often involve prompt engineering to guide the proprietary LLMs in SQL generation [47] or fine-tuning the open-source LLMs on text-to-SQL datasets [9].
+Large language models (LLMs), such as the GPT series [103, 104, 105], have gained significant attention in recent years due to their ability to generate coherent and fluent text.
+Researchers have begun exploring the potential of LLMs for text-to-SQL by leveraging their extensive knowledge reserves and superior generation capabilities [6, 8].
+These approaches often involve prompt engineering to guide proprietary LLMs in SQL generation [106] or fine-tuning open-source LLMs on text-to-SQL datasets [8].
 The integration of LLMs in text-to-SQL is still an emerging research area with significant potential for further exploration and improvement.
-Researchers are investigating approaches to better leverage the knowledge and reasoning capabilities of LLMs, incorporate domain-specific knowledge [33, 31], and develop more efficient fine-tuning strategies [10].
-As the field continues to evolve, we anticipate the development of more advanced and superior LLM-based implementations that will elevate the performance and generalization capabilities of text-to-SQL to new heights.
+Researchers are investigating ways to better leverage the knowledge and reasoning capabilities of LLMs, incorporate domain-specific knowledge [1, 69], and develop more efficient fine-tuning strategies [24].
+As the field continues to evolve, we anticipate the development of more advanced and superior LLM-based implementations that will elevate the performance and generalization of text-to-SQL to new heights.
 III Benchmarks & evalution
@@
 The original dataset indicates that the dataset is designed with a corresponding database, while post-annotated datasets involve annotating new components within existing datasets and databases rather than releasing a new database.
-Original Dataset Release Time #Example #DB #Table/DB #Row/DB Characteristics
-BIRD [33] May-2023 12,751 95 7.3 549K Cross-domain, Knowledge-augmented
-KaggleDBQA [36] Jun-2021 272 8 2.3 280K Cross-domain
-DuSQL [34] Nov-2020 23,797 200 4.1 - Cross-domain, Cross-lingual
-SQUALL [44] Oct-2020 11,468 1,679 1 - Knowledge-augmented
-CoSQL [35] Sep-2019 15,598 200 - - Cross-domain, Context-dependent
-Spider [13] Sep-2018 10,181 200 5.1 2K Cross-domain
-WikiSQL [14] Aug-2017 80,654 26,521 1 17 Cross-domain
-Post-annotated Dataset Release Time Source Dataset Special Setting Characteristics
-ADVETA [37] Dec-2022 Spider, etc. Adversarial table perturbation Robustness
-Spider-SS&CG [38] May-2022 Spider Splitting example into sub-examples Context-dependent
-Spider-DK [39] Sep-2021 Spider Adding domain knowledge Knowledge-augmented
-Spider-SYN [40] Jun-2021 Spider Manual synonym replacement Robustness
-Spider-Realistic [41] Oct-2020 Spider Removing column names in question Robustness
-CSpider [42] Sep-2019 Spider Chinese version of Spider Cross-lingual
-SParC [43] Jun-2019 Spider Annotate conversational contents Context-dependent
+For each dataset, the languages utilized are denoted by the following abbreviations: EN for English, ZH for Chinese, and VI for Vietnamese.
+Original Dataset Release Time #Example #DB #Table/DB #Row/DB Language Characteristics
+BIRD-CRITIC [1] Feb-2025 600 - - - EN Knowledge-augmented, Long-context
+Spider 2.0 [42] Nov-2024 632 213 - - EN Knowledge-augmented, Long-context
+BULL [45] Jan-2024 4,966 3 26 - EN Specialized-domain
+BIRD [1] May-2023 12,751 95 7.3 549K EN Cross-domain, Knowledge-augmented
+KaggleDBQA [33] Jun-2021 272 8 2.3 280K EN Cross-domain
+DuSQL [31] Nov-2020 23,797 200 4.1 - EN Cross-domain, Cross-lingual
+SQUALL [41] Oct-2020 11,468 1,679 1 - EN Knowledge-augmented
+CoSQL [32] Sep-2019 15,598 200 - - EN Cross-domain, Context-dependent
+Spider [11] Sep-2018 10,181 200 5.1 2K EN Cross-domain
+WikiSQL [12] Aug-2017 80,654 26,521 1 17 EN Cross-domain
+Post-annotated Dataset Release Time Source Dataset Special Setting Language Characteristics
+Dr. Spider [44] Jan-2023 Spider Perturbations in DB, query and SQL EN Robustness
+ADVETA [34] Dec-2022 Spider, etc. Adversarial table perturbation EN Robustness
+Spider-SS&CG [35] May-2022 Spider Splitting example into sub-examples EN Context-dependent
+Spider-DK [36] Sep-2021 Spider Adding domain knowledge EN Knowledge-augmented
+Spider-SYN [37] Jun-2021 Spider Manual synonym replacement EN Robustness
+Spider-Vietnamese [43] Nov-2020 Spider Vietnamese version of Spider VI Cross-lingual
+Spider-Realistic [38] Oct-2020 Spider Removing column names in question EN Robustness
+CSpider [39] Sep-2019 Spider Chinese version of Spider ZH Cross-lingual
+SParC [40] Jun-2019 Spider Annotate conversational contents EN Context-dependent
 III-A Datasets
 As shown in Tab.
-I, we classify the datasets into ’Original Datasets’ and ’Post-annotated Datasets’ based on whether they were released with the original dataset and databases or created by adapting existing datasets and databases with special settings.
-For the original datasets, we provide a detailed analysis, including the number of examples, the number of databases, the number of tables per database, and the number of rows per database.
-For the post-annotated datasets, we identify their source dataset and describe the special setting applied to them.
-To illustrate the potential opportunities of each dataset, we annotate them based on their characteristics.
+I, we categorize the datasets into “Original Datasets” and “Post-annotated Datasets” based on whether they were released with the original dataset and databases or developed by adapting existing datasets and databases with special settings.
+For the original datasets, we offer a detailed analysis, including the number of examples, databases, tables per database, and rows per database.
+For post-annotated datasets, we identify their source dataset and describe the special settings applied to them.
+To highlight the potential opportunities of each dataset, we annotate them according to their characteristics.
 The annotations are listed in the rightmost column of Tab.
@@
 ADVETA [34] introduces adversarial table perturbation (ATP), which perturbs tables by replacing original column names with misleading alternatives and inserting new columns with high semantic associations but low semantic equivalency.
+Based on the generation ability of large-scale pre-trained language models, Dr.
+Spider [44] provided 17 different perturbation strategies on databases, natural language queries and SQL statements to measure the robustness of text-to-SQL system.
 These perturbations lead to significant drops in accuracy, as a text-to-SQL system with low robustness may be misled by incorrect matches between tokens in NL questions and database entities.
 III-A5 Cross-lingual Dataset
-SQL keywords, function names, table names, and column names are typically written in English, posing challenges for applications in other languages.
-CSpider [42] translates the Spider dataset into Chinese, identifying new challenges in word segmentation and cross-lingual matching between Chinese questions and English database contents.
+The SQL keywords, function names, table and column names are typically written in English, posing challenges for applications in other languages.
+CSpider [39] addresses this by translating the Spider dataset into Chinese, identifying new challenges in word segmentation and cross-lingual matching between Chinese questions and English database contents.
 DuSQL [31] introduces a practical text-to-SQL dataset with Chinese questions and database contents provided in both English and Chinese.
+Spider-Vietnamese [43] translates the Spider dataset into Vietnamese, empirically revealing several strategies to enhance text-to-SQL methods within the Vietnamese context.
+III-A6 Long-context dataset
+In real-world database applications, numerous challenging scenarios demand complex SQL queries with significant token lengths, often exceeding 100 tokens, involving multiple intricate operations.
+Spider 2.0 [42] exemplifies this complexity by presenting enterprise-level text-to-SQL problems that require advanced reasoning across various SQL queries and dialects.
+This benchmark reveals the limitations of current models, as demonstrated by lower success rates—with advanced LLMs solving only 17.1% of tasks—underscoring its unique challenges.
+In parallel, BIRD-CRITIC [^2] https://bird-critic.github.io/ [1] introduces a diagnostic dimension, comprising 600 tasks designed to assess the efficacy of LLMs in resolving real-world user issues across diverse SQL dialects.
+Together, these datasets push the boundaries of text-to-SQL research, emphasizing the challenges of complex, long-context SQL environments.
+III-A7 Specialized-domain dataset
+Despite advancements in text-to-SQL research, there is still a shortage of specialized datasets tailored for domain-specific analysis.
+To address this gap, BULL [45] introduces a dedicated text-to-SQL benchmark for the financial sector.
+The dataset encompasses databases related to funds, stocks, and macroeconomic data, offering a comprehensive platform to evaluate and enhance text-to-SQL methods in financial applications, effectively tackling the financial domain unique text-to-SQL challenges.
 III-B Evaluation Metrics
@@
 III-B1 Content Matching-based Metrics
-SQL content matching metrics focus on comparing the predicted SQL query with the ground truth SQL query based on their structural and syntactic similarities.
+SQL content matching metrics focus on comparing the predicted SQL with the ground truth based on their structural and syntactic similarities.
   • Component Matching (CM) [11] evaluates the performance of text-to-SQL system by measuring the exact match between predicted and ground truth SQL components—SELECT, WHERE, GROUP BY, ORDER BY, and KEYWORDS—using the F1 score.
@@
   • Valid Efficiency Score (VES) [1] is defined to measure the efficiency of valid SQL queries.
-A valid SQL query is a predicted SQL query whose executed results exactly match the ground truth results.
+A valid SQL query is a predicted SQL whose executed results exactly match the ground truth results.
 Specifically, VES evaluates both the efficiency and accuracy of predicted SQL queries.
-For a text dataset with $N$ examples, VES can be computed by:
+For a text dataset with $N$ examples, VES is computed by:
     $\text{VES}=\frac{1}{N}\sum_{n=1}^{N}\mathbbm{1}(V_{n},\hat{V}_{n})\cdot\textbf{R}(Y_{n},\hat{Y}_{n}),$ (1)
@@
 BIRD benchmark [1] ensures the stability of this metric by computing the average of $\textbf{R}(Y_{n},\hat{Y}_{n})$ over 100 runs for each example.
-Most of the recent LLM-based text-to-SQL studies focus on these four datasets: Spider [13], Spider-Realistic [41], Spider-SYN [40], and BIRD [33]; and these three evaluation methods: EM, EX, and VES, we will primarily focus on them in the following analysis.
+Most of the recent LLM-based text-to-SQL studies focus on BIRD and Spider-related datasets and execution-based metrics, the corresponding details can be found at Tab. III.
 IV Methods
-The implementation of current LLM-based applications mostly relies on in-context learning (ICL) (prompt engineering) [87, 88, 89] and fine-tuning (FT) [90, 91] paradigms since the powerful proprietary models and well-architected open-source models are being released in large quantities [86, 45, 92, 93, 94, 95].
-LLM-based text-to-SQL systems follow these paradigms for implementation.
-In this survey, we will discuss them accordingly.
+The implementation of the most current LLM-based text-to-SQL methods relies on in-context learning (ICL) [107, 108, 109] and fine-tuning (FT) [110, 111] paradigms due to the release of powerful proprietary and well-architected open-source LLMs in large quantities [105, 112, 113, 114, 115, 116].
+In this section, we will discuss these paradigms with detailed categorization accordingly.
 IV-A In-context Learning
-Through extensive and widely recognized research, prompt engineering has been proven to play a decisive role in the performance of LLMs [96, 28], also impacting the SQL generation in different prompt styles [9, 46].
-Necessarily, developing text-to-SQL methods in the in-context learning (ICL) paradigm is valuable for achieving promising improvement.
-The implementation of LLM-based text-to-SQL process to generate executable SQL query $Y$ can be formulated as:
-$Y=f(\mathcal{Q},\mathcal{S},\mathcal{I}\mid\theta),$ (3)
-where $\mathcal{Q}$ represents the user question.
-$\mathcal{S}$ is the database schema/content, which can be decomposed as $\mathcal{S}=\langle\mathcal{C},\mathcal{T},\mathcal{K}\rangle$, where $\mathcal{C}=\{c_{1},c_{2},...\}$ and tables $\mathcal{T}=\{t_{1},t_{2},...\}$ represent the collection of various columns and tables, $\mathcal{K}$ is the potentially external knowledge (e.g.
-foreign key relationships [49], schema linking [30] and domain knowledge [33, 31]).
-$\mathcal{I}$ represents the instruction for the text-to-SQL task, which provides indicative guidance to trigger the LLMs for generating an accurate SQL query.
-$f(\cdot\mid\theta)$ is a LLM with parameter $\theta$.
-In the in-context learning (ICL) paradigm, we utilize an off-the-shelf text-to-SQL model (i.e., parameter $\theta$ of the model is frozen) for generating the predicted SQL query.
-Various well-designed methods within the ICL paradigm have been adopted for LLM-based text-to-SQL tasks.
-We group them into five categories $\mathbf{C}_{0:4}$, including $\mathbf{C}_{0}$-Trivial Prompt, $\mathbf{C}_{1}$-Decomposition, $\mathbf{C}_{2}$-Prompt Optimization, $\mathbf{C}_{3}$-Reasoning Enhancement, and $\mathbf{C}_{4}$-Execution Refinement, the representative methods of each category are given in Tab.
+Through extensive and widely recognized research, ICL (also called prompt engineering) has been proven to play a decisive role in the performance of LLMs across various tasks [117, 26, 118].
+Consequently, it also impacts SQL generation across different prompt styles [8, 119].
+The ICL implementation of the LLM-based text-to-SQL process to generate a predicted SQL query $\hat{Y}$, can be formulated as:
+$\hat{Y}=\pi\left(I,Q,\mathcal{S}\mid\theta\right),$ (3)
+where $I$ represents the instruction for the text-to-SQL task, $Q$ represents the user question.
+$\mathcal{S}$ is the database schema/content, which can be decomposed as $\mathcal{S}=\langle\mathcal{C},\mathcal{T},\mathcal{K}\rangle$.
+Here, $\mathcal{C}=\{c_{1},c_{2},...\}$ and $\mathcal{T}=\{t_{1},t_{2},...\}$ represent the collection of various columns $c$ and tables $t$, respectively.
+$\mathcal{K}$ is the potentially supplementary knowledge, such as foreign key relationships [67], schema linking [28, 29], and external knowledge [1, 69].
+$\pi(\cdot\mid\theta)$ is a LLM $\pi$ with parameter $\theta$.
+In the ICL paradigm, LLM-based methods primarily utilize an off-the-shelf text-to-SQL model, meaning the parameter $\theta$ of the model is frozen, for generating the predicted SQL query.
+We further specify the prompt engineering strategies in the ICL paradigm as zero-shot and few-shot prompting.
+Trained through massive data, LLMs have a strong overall proficiency in different downstream tasks with zero-shot and few-shot prompting [120, 110, 121], which is widely recognized and utilized in real-world applications.
+Formally, the overall input $P_{0}$ of zero-shot prompting for text-to-SQL can be obtained by concatenating $I$, $\mathcal{S}$, and $Q$ as introduced above:
+$P_{0}=I\oplus\mathcal{S}\oplus Q.$ (4)
+Subsequently, the overall input $P_{k}$ for few-shot prompting (specifically, $k$-shot) can be formulated as an extension of $P_{0}$:
+$P_{k}=\{\mathcal{E}_{1},\mathcal{E}_{2},\ldots,\mathcal{E}_{k}\}\oplus P_{0},$ (5)
+where $k$ is the number of provided instances (examples) in few-shot prompting.
+$\mathcal{E}_{i}$ represents the $i$-th few-shot instance, each of which can be decomposed as $\mathcal{E}_{i}=\left(\mathcal{S}_{i},Q_{i},Y_{i}\right)$, where $Y_{i}$ is the gold SQL query of $Q_{i}$.
+The index $i\in\{1,\ldots,k\}$ also serves as the corresponding serial number for $\mathcal{S}_{i}$, $Q_{i}$, and $Y_{i}$.
+The OpenAI demonstration [^3] The prompt format follows the official document provided by the OpenAI platform: https://platform.openai.com/examples/default-sql-translate provides an indicative example and regulation for the prompting process in LLM-based text-to-SQL [28].
+The prompt designs and formats utilized by a substantial portion of the ICL methods discussed in this survey are developed based on the OpenAI regulation.
+Furthermore, most of these ICL methods incorporate zero-shot and few-shot prompting into their text-to-SQL frameworks.
+As the mainstream research approach in LLM-based text-to-SQL, various well-designed methods within the ICL paradigm have been adopted to enhance LLM’s SQL generation.
+In this section, we group these existing methods into five categories, $\mathbf{C}_{0:4}$, including $\mathbf{C}_{0}$-Vanilla Prompting, $\mathbf{C}_{1}$-Decomposition, $\mathbf{C}_{2}$-Prompt Optimization, $\mathbf{C}_{3}$-Reasoning Enhancement, and $\mathbf{C}_{4}$-Execution Refinement.
+The representative methods of each category are provided in Tab.
 II.
-$\mathbf{C}_{0}$-Trivial Prompt
-Trained through massive data, LLMs have a strong overall proficiency in different downstream tasks with zero-shot and few-shot prompting [97, 90, 98], which is widely recognized and used in real-world applications.
-In our survey, we categorized the above prompting approaches without the well-designed framework as trivial prompts (vanilla prompt engineering).
-As introduced above, Eq.
-3 formulated the process of LLM-based text-to-SQL, which can also represent zero-shot prompting.
-The overall input $\mathcal{P}_{0}$ can be obtained by concatenating $\mathcal{I}$, $\mathcal{S}$ and $\mathcal{Q}$:
-$\mathcal{P}_{0}=\mathcal{I}\oplus\mathcal{S}\oplus\mathcal{Q}.$ (4)
-To regulate the prompting process, the OpenAI demonstration [^2] The prompt style that follows the official document from OpenAI platform: https://platform.openai.com/examples/default-sql-translate is set as the standard (trivial) prompt [30] for text-to-SQL.
-TABLE II: Typical methods used for in-context learning (ICL) in LLM-based text-to-SQL.
-The full table of existing methods with categorization $\mathbf{C}_{1:4}$ and more details are listed in Tab.
+We will introduce each category accordingly and summarize their characteristics at the beginning of each corresponding paragraph.
+IV-A1 $\mathbf{C}_{0}$-Vanilla Prompting
+In our survey, we identified a category of prompting strategies that do not propose a well-designed ICL framework.
+These strategies are based solely on zero-shot and few-shot prompting.
+Despite their simplicity, they offer valuable insights and findings on the ICL paradigm in LLM-based text-to-SQL.
+We refer to these strategies as vanilla prompting strategies and further divide their studies into zero-shot and few-shot studies.
+TABLE II: A summary of representative methods used in the ICL paradigm for LLM-based text-to-SQL is provided.
+The complete table, along with a detailed categorization of existing well-designed methods denoted as $\mathbf{C}_{1:4}$, is available in Tab.
 III.
+*Note that SimpleDDL is a prompting style that provides only table and column names [119], rather than a method.
 Methods Adopted by Applied LLMs
-$\mathbf{C}_{0}$-Trivial Prompt Zero-shot [7] ChatGPT
-Few-shot [9] ChatGPT
+$\mathbf{C}_{0}$-Vanilla Prompting SimpleDDL [119]* SQLCoder
 $\mathbf{C}_{1}$-Decomposition DIN-SQL [7] GPT-4
@@
 $\mathbf{C}_{4}$-Execution Refinement LEVER [78] Codex
-TABLE III: Well-designed methods used in in-context learning (ICL) paradigm for LLM-based text-to-SQL ordered by release time.
-The methods are grouped in four categories based on their implementation perspective: $\mathbf{C}_{1}$-Decomposition, $\mathbf{C}_{2}$-Prompt Optimization, $\mathbf{C}_{3}$-Reasoning Enhancement, $\mathbf{C}_{4}$-Execution Refinement.
-The method in multiple categories will be introduced respectively.
-*There are multiple applied LLMs in the corresponding method; we present the selected LLM with representative performance.
-CoT method are reported in multiple venues: NeurIPS’23 [33], EMNLP’23 [51], VLDB’24 [9], arXiv’24 [32].
+TABLE III: Well-designed methods in the ICL paradigm for LLM-based text-to-SQL are ordered by release time.
+The methods are grouped into four categories based on their implementation perspective: $\mathbf{C}_{1}$-Decomposition, $\mathbf{C}_{2}$-Prompt Optimization, $\mathbf{C}_{3}$-Reasoning Enhancement, and $\mathbf{C}_{4}$-Execution Refinement.
+Methods that fall into multiple categories will be introduced accordingly.
+The dataset column consists of six text-to-SQL datasets, which can be clarified by names: Spider [11], BIRD [1], Spider-DK [36] Spider-SYN [37], Spider-Realistic [38], and Spider 2.0 [42].
+We use the reference ID in the table due to column space limitations.
+*There are multiple applied LLMs in the corresponding method; the model with representative performance is presented.
+CoT method is introduced in multiple venues: NeurIPS’23 [1], EMNLP’23 [47], VLDB’24 [8], arXiv’24 [30].
 Methods Applied LLMs Dataset Metrics $\mathbf{C}_{1}$ $\mathbf{C}_{2}$ $\mathbf{C}_{3}$ $\mathbf{C}_{4}$ Release Time Publication Venue
@@
 DESEM+P [62] ChatGPT [11, 37] EX Apr-2023 PRICAI’23
-DIN-SQL [8] GPT-4* [33, 13] EX, EM, VES Apr-2023 NeurIPS’23
+DIN-SQL [7] GPT-4* [1, 11, 42] EX, EM, VES Apr-2023 NeurIPS’23
 CoT [47, 1, 8, 30] GPT-4 [11, 38, 1] EX, VES May-2023 Multiple Venues
@@
 Least-to-Most [47] Codex [11] EX May-2023 EMNLP’23
-SQL-PaLM [53] PaLM-2 [13] EX May-2023 arXiv’23
+SQL-PaLM [73] PaLM-2 [11] EX May-2023 TMLR’24
 RAG+SP&DRC [65] ChatGPT [11] EX Jul-2023 ICONIP’23
 C3 [28] ChatGPT [11] EX Jul-2023 arXiv’23
-DAIL-SQL [9] GPT-4* [13, 33, 41] EX, EM, VES Aug-2023 VLDB’24
+DAIL-SQL [8] GPT-4* [11, 1, 38, 42] EX, EM, VES Aug-2023 VLDB’24
 ODIS [66] Codex* [11] EX Oct-2023 EMNLP’23 Findings
 ACT-SQL [67] GPT-4* [11, 37] EX, EM Oct-2023 EMNLP’23 Findings
-MAC-SQL [57] GPT-4* [13, 33] EX, EM, VES Dec-2023 arXiv’23
+MAC-SQL [48] GPT-4* [11, 1] EX, EM, VES Dec-2023 COLING’25
 DEA-SQL [49] GPT-4 [11] EX Feb-2024 ACL’24 Findings
-FUSED [65] ChatGPT* [13] EX Feb-2024 arXiv’24
+FUSED [68] ChatGPT* [11] EX Feb-2024 EMNLP’24 Findings
 DELLM [69] GPT-4* [11, 1] EX, VES Feb-2024 ACL’24 Findings
@@
 SQL-CRAFT [74] GPT-4* [11, 1] EX Feb-2024 arXiv’24
-FUXI [66] GPT-4* [33] EX Feb-2024 arXiv’24
+FUXI [75] GPT-4* [1] EX Feb-2024 EMNLP’24
 MetaSQL [50] GPT-4* [11] EX, EM Feb-2024 ICDE’24
+$\textit{{R}}^{3}$ [51] GPT-4* [11, 1] EX Feb-2024 arXiv’24
 PET-SQL [52] GPT-4 [11] EX Mar-2024 arXiv’24
 PURPLE [53] GPT-4* [11, 37, 38] EX, EM Mar-2024 ICDE’24
-Zero-shot: Many research works [7, 27, 46] utilize zero-shot prompting, studying mainly on the influence of the style of prompt construction and the zero-shot performance of various LLMs for text-to-SQL.
-As an empirical evaluation, [7] evaluates the baseline text-to-SQL capabilities of different early-developed LLMs [99, 85, 100] and the performance for different prompting styles.
-The results indicate that prompt design is critical for the performance, with error analysis, [7] propose more database content can harm the overall accuracy.
-Since ChatGPT emerged with impressive capabilities in conversational scenarios and code generation [101], [27] assesses its performance of text-to-SQL.
-With zero-shot settings, the results demonstrate that ChatGPT has a promising text-to-SQL performance compared to the state-of-the-art PLM-based systems.
-For fair comparability, [47] reveal effective prompt construction for LLM-based text-to-SQL; they study different styles of prompt construction and make conclusions of zero-shot prompt design based on the comparisons.
-Primary keys and foreign keys carry contiguous knowledge of different tables.
-[49] studies their impact by incorporating these keys into various prompt styles with different database content to analyze zero-shot prompting results.
-A benchmark evaluation [9] also studies the influence of foreign keys, with five different prompt representation styles, each style can be considered as the permutation and combination of the instruction, rule implication, and foreign key.
-Apart from the foreign key, this study also explores zero-shot prompting combined with the rule implication “no explanation” to collect concise outputs.
-Empowered by the annotated external knowledge of human experts, [33] follows the standard prompting and achieves improvement by incorporating the provided annotated oracle knowledge.
-With the explosion of open-source LLMs, according to the results of similar evaluation, these models are also capable of zero-shot text-to-SQL task [45, 50, 46], especially code generation models [48, 46].
-For zero-shot prompting optimization, [46] raises a challenge for designing an effective prompt template for LLMs; the former prompt construction lacks structure uniformity, which makes it hard to find out a concrete element within a prompt constructing template influences the performance of LLMs.
-They address this challenge by investigating a more unified series of prompt templates warping with different prefixes, infixes, and postfixes.
-Few-shot: The technique of few-shot prompting is widely used in both practical applications and well-designed research, which has been proven efficient for eliciting better performance of LLMs [28, 102].
-The overall input prompt of the few-shot approach LLM-based text-to-SQL can be formulated as an extension of Eq.
-3:
-$\mathcal{P}_{n}=\{\mathcal{F}_{1},\mathcal{F}_{2},\ldots,\mathcal{F}_{n}\}\oplus\mathcal{P}_{0},$ (5)
-where the $\mathcal{P}_{n}$ represent the input prompt for $n$-shot learning, $n$ is the provided instances (examples) number; $\mathcal{F}$ denote the few-shot instance, which can be decomposed as $\mathcal{F}_{i}=(\mathcal{S}_{i},\mathcal{Q}_{i},Y_{i})$, $i$ is the serial number of instances.
-The study of few-shot prompting focuses on the number of representations and few-shot instance selection.
-As empirical studies, few-shot prompting for text-to-SQL are evaluated in multiple datasets with various LLMs [8, 32], exhibit a solid performance during the comparison with zero-shot prompting.
-[33] provides a 1-shot detailed example to trigger the text-to-SQL model for generating accurate SQL.
-[55] study the effect of the number of few-shot examples.
-[52] focus on the sampling strategies by studying the similarity and the diversity between different demonstrations, setting random sampling as the baseline, and evaluating different strategies and their combination for comparison.
-Furthermore, above the similarity-based selection, [9] evaluated masked question similarity selection and the upper limit of similarity approaches with various numbers of few-shot examples.
-A study of difficult-level samples selection [51] compared the performance of few-shot Codex [100], with random selection and difficulty-based selection for few-shot instances on difficulty categorized dataset [13, 41].
-Three difficulty-based selection strategies are devised based on the number of selected samples at different difficulty levels.
-[49] utilize a hybrid strategy for selecting samples, which combines static examples and similarity-based dynamic examples for few-shot prompting.
-In their settings, they also evaluate the impact of different input schema styles and various static and dynamic exemplar numbers.
-The impact of cross-domain few-shot examples is also being studied [54].
-When incorporating in-domain and out-of-domain examples with different numbers, the in-domain demonstration outperforms zero-shot and out-of-domain examples, which will also perform better as the number of examples rises.
-To explore the detailed construction of input prompt, [53] compare the concise and verbose prompt design approaches.
-The former style splits the schema, the column names, and the primary and foreign keys by bar, and the latter organizes them as natural language descriptions.
-$\mathbf{C}_{1}$-Decomposition
-As an intuitive solution, decomposing a challenging user question into simpler sub-questions or using multi-components for implementation can reduce the complexity of the overall text-to-SQL task [51, 8].
-Dealing with less complexity, LLMs have the potential to elicit more accurate SQL generation.
-The decomposition-based approaches for LLM-based text-to-SQL are categorized into two paradigms: (1) sub-task decomposition, provides additional parsing to assist the final SQL generation by decomposing the overall text-to-SQL task into more manageable effective sub-tasks (e.g., schema linking [71], domain classification [54]).
-(2) sub-question decomposition, divides the user question into sub-questions to reduce the question’s complexity and difficulty, then generates the sub-SQL by solving these questions to deduce the final SQL query.
-The technical novelty of the decomposition paradigm
-DIN-SQL [8] proposed a decomposed in-context learning method consisting of four modules: schema linking, classification & decomposition, SQL generation, and self-correction.
-DIN-SQL first generates the schema linking between the user question and the target database; the following module decomposes the user question into correlated sub-questions and does a difficulty classification.
-Based on the above information, the SQL generation module generates a corresponding SQL, and the self-correction module identifies and corrects the potential errors in the predicted SQL.
-This approach comprehensively considers the sub-questions decomposition as a module of sub-tasks decomposition.
-Coder-Reviewer [56] framework proposed a re-ranking method, combining Coder models for the generation and Reviewer models to evaluate the likelihood of the instruction.
-Refer to the Chain-of-Thought [103] and Least-to-Most prompting [104], QDecomp [51] introduce question decomposition prompting, which follows the question reduction stage in least-to-most prompting and instruct the LLM to decompose the original complex question as the intermediate reasoning steps.
-C3 [30] consists of three key components: clear prompting, calibration bias prompting, and consistency; these components are accomplished by assigning ChatGPT with different tasks.
-Firstly, the clear prompting component generates the schema linking and the distilled question-relevant schema as a clear prompt.
-Then, a multi-turn dialogue about text-to-SQL hints is utilized as a calibration bias prompt, which combines with the clear prompt to guide the SQL generation.
-The generated SQL queries are selected by consistency and execution-based voting to get the final SQL.
-MAC-SQL [57] presents a multi-agent collaborating framework; the text-to-SQL process is finished as the collaboration of the agents: Selector, Decomposer, and Refiner.
-The Selector preserves relevant tables for user questions; the Decomposer breaks down user questions into sub-questions and provides solutions; finally, the Refiner validates and refines the defective SQL.
-DEA-SQL [58] introduces a workflow paradigm aiming to enhance the attention and problem-solving scope of LLM-based text-to-SQL through decomposition.
-This method decomposes the overall task, enabling the SQL generation module to have the corresponding prerequisite (information determination, question classification) and subsequent (self-correction, active learning) sub-tasks.
-The workflow paradigm allows the LLM to generate more accurate SQL queries.
-SGU-SQL [32] is a structure-to-SQL framework, leveraging the inherent structure information to assist SQL generation.
-Specifically, the framework constructs a graph structure for the user question and the corresponding database respectively, then uses the encoded graphs to construct structure linking [105, 106].
-A meta operator decomposes the user question with a grammar tree and finally designs the input prompt with meta-operation in SQL.
-MetaSQL [59] introduces a three-stage approach for SQL generation: decomposition, generation, and ranking.
-The decomposition stage uses semantic decomposition and metadata composition to process the user question.
-Taking the previously processed data as input, a text-to-SQL model using metadata-conditioned generation to generate some candidate SQL queries.
-Finally, a two-stage ranking pipeline is applied to get a global-optimal SQL query.
-PET-SQL [60] proposed a prompt-enhanced two-stage framework.
-Firstly, an elaborated prompt instructs the LLMs to generate preliminary SQL (PreSQL) where some few-shot demonstrations are selected based on similarity.
-Then, schema linking is found based on PreSQL and combined to prompt the LLMs to generate the Final SQL (FinSQL).
-Finally, multiple LLMs are utilized to generate a FinSQL, ensuring consistency based on the execution results.
-$\mathbf{C}_{2}$-Prompt Optimization
-As previously introduced, few-shot learning is widely studied for prompting LLMs [85].
-For LLM-based text-to-SQL with in-context learning, trivial few-shot approaches obtained promising results [33, 8, 9], further optimization of few-shot prompting has the potential to elicit better performance.
-Since the accuracy of SQL generation in off-the-shelf LLMs largely depends on the quality of the corresponding input prompt [107], many decisive factors that can influence the quality of the prompt have become points of focus of the current research [9] (e.g., quality and quantity in the few-shot organization, the similarity between user questions and few-shots instances, external knowledge/hints) The process of improving prompt quality can be summarized as the prompt’s optimization, including advanced few-shot sampling strategies, schema information augmentation, and external knowledge integration.
-DESEM [62] is a prompt engineering framework with de-semanticization and skeleton retrieval.
-The framework first employs domain-specific words masking module to remove the semantic tokens in user questions that preserve the intentions.
-It then utilizes an adjustable prompting module that retrieves the few-shot examples with identical question intentions and incorporates schema-relevance filtering to guide the LLM’s SQL generation.
-The QDecomp [51] framework introduces the InterCOL mechanism to incrementally incorporate the decomposed sub-questions with correlative table and column names.
-With difficulty-based selection, the few-shot examples for QDecomp are difficult-level sampled.
-Besides similarity-diversity sampling, [52] proposed SD+SA+Voting (Similarity-Diversity+Schema Augmentation+Voting) sampling strategy.
-They first employ semantic similarity and $k$-Means cluster diversity for sampling few-shot examples and then enhance the prompt with schema knowledge (semantic or structure augmentation).
-C3 [30] framework comprises a clear prompting component, which takes the question and schema as the LLMs input, generates a clear prompt that includes a schema that removes the redundant information irrelevant to the user question and a schema linking, and also a calibration component providing hints.
-The LLMs take their composition as context-augmented prompts for SQL generation.
-A retrieval-augmented framework is introduced with sample-aware prompting [64], which simplifies the original question and extracts the question skeleton from the simplified question, then finishes the sample retrieval in the repository according to skeleton similarities.
-The retrieved samples are combined with the original question for few-shot prompting.
-ODIS [54] introduces the selection of a sample with out-of-domain demonstrations and in-domain synthetic data, which retrieves few-shot demonstrations from hybrid sources to augment the prompt representations.
-DAIL-SQL [9] proposed a novel approach to address the issues in few-shot sampling and organization, presenting a better balance between the quality and quantity of few-shot examples.
-DAIL Selection first masks domain-specific words in user and few-shot example questions, then ranks the candidate examples based on the embedded Euclidean distance.
-Meanwhile, the similarity between the pre-predicted SQL queries is calculated.
-Finally, the selection mechanism obtains the similarity-sorted candidates according to the pre-set criteria.
-The few-shot examples are guaranteed good similarity with both questions and SQL queries with this method.
-ACT-SQL [49] proposed dynamic examples in few-shot prompting, which is selected according to similarity score.
-FUSED [65] are presented to build a high-diversity demonstrations pool through human-free multiple-iteration synthesis to improve the diversity of the few-shot demonstrations.
-The pipeline of FUSED samples the demonstrations to be fused by clustering, then fuse the sampled demonstrations to construct the pool to enhance few-shot learning.
-Knowledge-to-SQL [31] framework aims to build a Data Expert LLM (DELLM) to provide knowledge for SQL generation.
-The DELLM is trained by supervised fine-tuning using human expert annotations [33] and further refined by preference learning with the database’s feedback.
-DELLM generates four categories of knowledge, the well-designed methods (e.g.
-DAIL-SQL [9], MAC-SQL [57]) incorporating the generated knowledge to achieve better performance for LLM-based text-to-SQL with in-context learning.
-$\mathbf{C}_{3}$-Reasoning Enhancement
-LLMs have exhibited promising capabilities in tasks involving commonsense reasoning, symbolic reasoning, and arithmetic reasoning [108].
-For text-to-SQL tasks, numeric and synonym reasoning frequently occur in realistic scenarios [33, 41].
-The prompting strategies for LLMs’ reasoning have the potential to enhance SQL generation capabilities.
-Recent studies primarily focus on integrating well-designed reasoning-enhanced methods for text-to-SQL adaptation, improving LLMs to address the challenge about complex questions that require sophisticated reasoning [^3] The difference between multi-step reasoning (e.g., Chain of Thought) and the decomposition paradigm is that the former research focuses on advancing inherent reasoning within a single-turn generation, whereas the latter study involves using different components to assist the final generation through multiple calls of LLMs.
-and the self-consistency in SQL generation.
-Chain-of-Thoughts (CoT) prompting technique [103] involves a comprehensive reasoning process that guides LLMs towards accurate deduction, eliciting reasoning in LLMs.
-The study of LLM-based text-to-SQL utilizes CoT prompting as rule implication [9], which setting the instruction “Let’s think step by step” in prompt construction [33, 51, 9, 32].
-However, the straightforward (original) CoT strategy has not demonstrated the potential in text-to-SQL tasks that it has in other reasoning tasks; studying CoT for adaptations is still an ongoing research [51].
-Since CoT prompting always uses static examples with human annotation for demonstrations, which requires empirical judgment for the effective selection of few-shot examples, and manual annotating is also an essential need.
-As a solution, ACT-SQL [49] proposed a method to generate CoT examples automatically.
-Specifically, given a question, ACT-SQL truncates a set of slices of the question and then enumerates every column appearing in the corresponding SQL query.
-Each column will be linked with its most relevant slice through the similarity function and appended to the CoT prompt.
-Through systematical study for enhancing LLMs SQL generation incorporating CoT prompting, QDecomp [51] presents a novel framework to address the challenge for CoT how to come up with the reasoning steps to predict the SQL query.
-The framework utilizes every slice of the SQL query to construct a logical step in CoT reasoning, then employs natural language templates to articulate each slice of the SQL query and arranges them in the logical execution order.
-Least-to-Most [104] is another prompting technique that decomposes questions into sub-questions and then sequentially solves them.
-As iterative prompting, pilot experiments [51] demonstrate that it may be unnecessary for text-to-SQL parsing.
-Using detailed reasoning steps tends to have more error propagation issues.
-As a variant of CoT, Program-of-Thoughts (PoT) prompting strategy [109] are proposed to enhance arithmetic reasoning for LLMs.
-Through evaluation [55], PoT enhances the LLM for SQL generation, especially in complicated datasets [33].
-SQL-CRAFT [55] are proposed to enhance LLM-based SQL generation, which incorporates PoT prompting for Python-enhanced reasoning.
-PoT strategy requires the model to simultaneously generate the Python code and the SQL queries, enforcing the model to incorporate Python code in its reasoning process.
-Self-Consistency [110] is a prompting strategy improving reasoning in LLMs, which leverages the intuition that a complex reasoning problem typically admits multiple different ways of thinking, leading to its unique correct answer.
-In the text-to-SQL task, self-consistency is adapted to sampling a set of different SQL and voting for consistent SQL via execution feedback [30, 53].
-Similarly, the SD+SA+Voting [52] framework eliminates those with execution errors identified by the deterministic database management system (DBMS) and opts for the prediction that garners the majority vote.
-Furthermore, motivated by recent research on extending the capabilities of LLMs with tools, FUXI [66] are proposed to enhance LLMs SQL generation through effectively invoking crafted tools.
-$\mathbf{C}_{4}$-Execution Refinement
-When designing criteria for accurate SQL generation, the priority is always whether a generated SQL can be successfully executed and retrieve content to correctly answer the user question [13].
-As a complex programming task, generating the correct SQL in one go is challenging.
-Intuitively, considering the execution feedback/results in SQL generation assists the alignment to the corresponding database environment, which allows the LLMs to gather the potential executed errors and results to refine the generated SQL or hold a majority vote [30].
-The execution-aware methods in text-to-SQL incorporate the execution feedback in two main approaches: 1) Incorporating the feedback through second round prompting for re-generation, for every SQL query generated in the initial response, it will be executed in the corresponding database, thus obtaining the database’s feedback.
-This feedback might be an error, or it might yield results that will be appended to the second round prompt.
-Through in-context learning of this feedback, LLMs are able to refine or re-generate the original SQL, thereby enhancing accuracy.
-2) Utilize execution-based selection strategies for the generated SQL, sample multiple generated SQL queries from LLM, and execute each in the database.
-Based on the execution results of each SQL query, use selection strategies (e.g., self-consistency, majority vote [60]) to define one SQL query from the SQL set that satisfies the criteria as the final predicted SQL.
-MRC-EXEC [67] introduced a natural language to code (NL2Code) translation framework with execution, which executes each sampled SQL query and selects the example with the minimal execution result–based Bayes risk [111].
-LEVER [68] proposed an approach to verify NL2Code with execution, utilizing a generation and execution module to collect sampled SQL set and their execution results, respectively, then using a learned verifier to output the probability of the correctness.
-Similarly, the SELF-DEBUGGING [48] framework is presented to teach LLMs to debug their predicted SQL via few-shot demonstrations.
-The model is able to refine its mistakes by investigating the execution results and explaining the generated SQL in natural language without human interventions
-As previously introduced, to incorporate the well-designed framework with execution feedback, two-stage implications are widely-used: 1.
-sampling a set of SQL queries.
-2.
-majority vote (self-consistency).
-Specifically, the C3 [30] framework removes the errors and identifies the most consistent SQL; The retrieval-augmented framework [64] introduced a dynamic revision chain, combining fine-grained execution messages with database content to prompt the LLMs to convert the generated SQL query into natural language explanation; the LLMs are requested to identify the semantic gaps and revise their own generated SQL.
-Although schema-filtering approaches enhance SQL generation, the generated SQL could be unexecutable.
-DESEM [62] incorporates a fallback revision to address the issue; it revises and regenerates the SQL base on different kinds of errors and sets termination criteria to avoid the loop.
-DIN-SQL [8] designed a generic and gentle prompt in their self-correction module; the generic prompt requests the LLMs to identify and correct the error, and the gentle prompt asks the model to check the potential issue.
-The multi-agent framework MAC-SQL [57] comprises a refiner agent, which is able to detect and automatically rectify SQL errors, taking SQLite error and exception class to regenerate fixed SQL.
-Since different questions may require different numbers of revisions, SQL-CRAFT [55] framework introduced interactive correction with an automated control determination process to avoid over-correction or insufficient correction.
-FUXI [66] considers the error feedback in tool-based reasoning for SQL generation.
-The Knowledge-to-SQL [31] introduces a preference learning framework incorporating the database execution feedback with a direct preference optimization [112] for refining the proposed DELLM.
-PET-SQL [60] proposed cross consistency, which comprises two variants: 1) naive voting: instruct multiple LLMs to generate the SQL query, then utilizing the majority vote for the final SQL base on different execution results; 2) fine-grained voting: refine the naive voting based on the difficulty level to mitigate the voting bias.
+Dubo-SQL [70] ChatGPT* [1] EX Apr-2024 arXiv’24
+TA-SQL [70] GPT-4* [11, 1, 36, 38] EX Apr-2024 ACL’24 Findings
+MCS-SQL [55] GPT-4 [11, 1] EX, VES May-2024 COLING’25
+CHESS [56] GPT-4o* [11, 1, 42] EX, VES May-2024 arXiv’24
+SuperSQL [57] GPT-4* [11, 1] EX, EM, VES Jun-2024 VLDB’24
+Distillery [58] GPT-4o-Mini* [1] EX Aug-2024 NeurIP’24 Workshop
+E-SQL [80] GPT-4o* [11, 1] EX, VES Sep-2024 arXiv’24
+CHASE-SQL [76] Gemini-1.5* [11, 1] EX, VES Oct-2024 ICLR’25
+RSL-SQL [59] DeepSeek-Chat* [11, 1] EX, VES Nov-2024 arXiv’24
+ICRL [71] ChatGPT [11] EX Jan-2025 COLING’25
+Gen-SQL [60] LLaMA-3* [11, 1] EX, EM, VES Jan-2025 COLING’25
+SAFE-SQL [72] GPT-4o-Mini* [11] EX, EM Feb-2025 arXiv’25
+ReFoRCE [61] o1-preview* [42] EX Feb-2025 arXiv’25
+Zero-shot studies: The zero-shot-based ICL strategies primarily focus on studying the influence of prompt representation styles and the baseline performance of various LLMs for text-to-SQL [6, 25, 119].
+As an empirical evaluation, [6] evaluates the baseline text-to-SQL capabilities of various early-developed LLMs [122, 104, 123] and their performance across different prompting styles.
+The experimental results indicate that prompt design is critical for performance.
+Through error analysis, this study [6] further suggests that including more database content can harm the overall accuracy of text-to-SQL.
+Since the emergence of ChatGPT [104] with its impressive capabilities in conversational scenarios and code generation [124], [25] evaluates its performance on Spider [11] and its variant [37, 36, 38] datasets.
+Under zero-shot settings, the results demonstrate that ChatGPT has promising text-to-SQL performance compared to state-of-the-art PLM-based systems.
+Aside from the proprietary models discussed above, similar evaluations indicate that open-source LLMs can also perform zero-shot text-to-SQL tasks [112, 82, 119], particularly code-specific models [79, 119].
+Primary keys and foreign keys convey continuous knowledge across different tables, making them an effective prompting style in zero-shot scenarios.
+[67] studied their impact by incorporating these keys into various prompt representation styles with different database content to analyze performance outcomes.
+A benchmark evaluation [8] also assessed the influence of foreign keys using five distinct prompt representation styles.
+Each style can be viewed as a permutation and combination of instruction, rule implication, and foreign key elements.
+In addition to using foreign keys, this study explores zero-shot prompting combined with the rule implication “no explanation” to produce concise outputs.
+With the help of annotated external knowledge from human experts, the evaluation in the BIRD benchmark [1] significantly improves by incorporating the provided annotated oracle knowledge.
+In terms of zero-shot prompting optimization, [119] highlights the challenge of designing effective prompt templates for LLMs.
+Previous simple prompt constructions often lack structural uniformity, complicating the identification of specific elements in a prompt template that influence LLM performance.
+This challenge is addressed by investigating a more unified set of prompt templates that include various prefixes, infixes, and postfixes.
+To ensure fair comparability, [106] explores effective prompt construction for LLMs by studying different styles of prompt representation and highlights their findings on the crucial role of table structure, content, prompt length, and optimal representations in effective prompting.
+Few-shot studies: The study of few-shot prompting strategies focuses on the number and selection of few-shot instances.
+Empirical studies have evaluated few-shot prompting for text-to-SQL across multiple datasets using various LLMs [7, 30], demonstrating solid improvements compared to zero-shot prompting.
+[1] provides a detailed 1-shot example to trigger the text-to-SQL model for generating accurate SQL.
+[74] explores the effect of the number of few-shot examples.
+[64] focuses on sampling strategies by examining the similarity and diversity between different demonstrations, setting random sampling as the baseline, and evaluating various strategies and their combinations for comparison.
+Furthermore, beyond similarity-based selection, [8] evaluates masked question similarity selection and the upper limit of similarity approaches with various numbers of few-shot examples.
+A study on difficulty-level sample selection [47] compared the performance of few-shot Codex [123], using random selection and difficulty-based selection for few-shot instances on difficulty-categorized datasets [11, 38].
+Three difficulty-based selection strategies are devised based on samples selected by difficulty level.
+[67] utilizes a hybrid strategy for selecting samples, combining static examples and similarity-based dynamic examples for few-shot prompting.
+In their settings, they also evaluate the impact of different input schema styles and various numbers of static and dynamic exemplars.
+The impact of cross-domain examples is also being studied [66].
+When incorporating in- and out-of-domain examples in different numbers, the in-domain demonstrations outperform zero-shot and out-of-domain examples, with performance improving as the number of examples increases.
+To explore the detailed construction of prompts, [73] compares concise and verbose prompt design strategies.
+The former style separates the schema, column names, primary keys, and foreign keys, while the latter organizes them as natural language descriptions.
+IV-A2 $\mathbf{C}_{1}$-Decomposition
+Decomposition-based methods reduce text-to-SQL complexity by dividing tasks or questions into manageable components, employing two main strategies:
+1) Multi-module collaboration: This strategy breaks the task into specialized stages (e.g., schema linking, SQL generation, and SQL refinement) handled by distinct modules or agents.
+C3 [28] employs a three-module pipeline involving schema linking via clear prompting, calibration bias prompting for SQL guidance, and consistency-based voting for final selection.
+Similarly, MAC-SQL [48] and CHESS [56] use multi-agent frameworks where dedicated agents handle schema selection, decomposition, and refinement.
+Workflow-centric methods, like Distillery [58] and DEA-SQL [49], decompose the task into retrieval, generation, and correction stages.
+Iterative refinement and structural decomposition are also prevalent.
+DIN-SQL [7] integrates a four-stage pipeline, including schema linking, classification, SQL generation, and self-correction.
+Coder-Reviewer [46] separates generation and validation into distinct models.
+PET-SQL [52] refines outputs through multi-stage prompting, while MCS-SQL [55] uses execution-guided selection as the final module after multiple SQL generation.
+2) Question decomposition: Complexity is reduced by splitting user questions into intermediate sub-questions via question decomposition.
+QDecomp [47] uses Least-to-Most prompting, breaking complex questions into step-by-step sub-questions to facilitate SQL generation.
+SGU-SQL [30] employs graph-based structure linking and meta-operators for decomposing grammar-aware sub-questions.
+Metadata-driven approaches, such as MetaSQL [50] and TA-SQL [54], further decompose questions using metadata conditioning and task-aligned schema linking.
+Meanwhile, frameworks like DIN-SQL [7] and MAC-SQL [48] use question decomposition as auxiliary modules but primarily focus on module-level collaboration.
+IV-A3 $\mathbf{C}_{2}$-Prompt Optimization
+Prompt optimization methods for text-to-SQL focus on enhancing prompt engineering quality. These methods are categorized into three key strategies:
+1) Advanced few-shot sampling: This strategy optimizes example selection through semantic similarity, diversity, or domain alignment.
+DESEM [62] and retrieval-augmented frameworks [65, 70, 58] de-semanticize questions to retrieve examples with matching intent skeletons or vector embeddings.
+DAIL-SQL [8] balances quality and quantity by masking domain-specific tokens and ranking examples based on the Euclidean distance between questions and SQL queries, while SAFE-SQL [72] combines skeleton masking with self-augmentation.
+ODIS [66] hybridizes out-of-domain and synthetic in-domain examples for robustness, and FUSED [68] enhances diversity through iterative demonstration synthesis.
+Dynamic selection is explored in ACT-SQL [67] and MCS-SQL [55], which adaptively retrieve examples based on defined similarity scores or masked similarity scores.
+2) Schema augmentation: These methods refine schema representations to reduce redundancy and improve relevance.
+C3 [28] distills question-relevant schemas by removing unrelated tables or columns, while RSL-SQL [59] simplifies schemas using NL descriptions.
+ReFoRCE [61] employs table compression and column exploration to manage long-context schema complexity.
+Gen-SQL [60] bridges NL-SQL gaps by generating pseudo-schemas that align with user questions.
+QDecomp [47] and SD+SA+Voting [64] further augment prompts with schema-aware semantic and structural hints.
+3) External knowledge integration: This group enhances prompts with domain-specific knowledge or feedback mechanisms.
+Knowledge-to-SQL [69] trains a DELLM to inject external knowledge about question-database relationships, while ICRL [71] integrates reinforcement learning with schema-graph-derived knowledge bases.
+MAC-SQL [48] and DIN-SQL [7] implicitly utilize decomposition patterns as procedural knowledge to guide generation.
+C3 [28] and SAFE-SQL [72] incorporate calibration bias and self-augmented feedback loops, respectively, to iteratively refine outputs.
+IV-A4 $\mathbf{C}_{3}$-Reasoning Enhancement
+Reasoning enhancement methods [^4] Unlike decomposition paradigms, these methods retain single-turn generation while enhancing intrinsic reasoning through structured prompting.
+improve LLMs’ ability to handle complex SQL through structured reasoning and consistency mechanisms.
+These approaches focus on three crucial strategies:
+1) Chain-of-thought (CoT)-based reasoning: CoT-based methods guide LLMs to generate intermediate reasoning steps before predicting SQLs.
+While standard CoT prompts (e.g., “Let’s think step by step”) show limited effectiveness in text-to-SQL [8, 1], adaptations help bridge this gap.
+ACT-SQL [67] automates CoT example generation by linking question slices to database columns via similarity matching.
+QDecomp [47] restructures SQL clauses into logical steps ordered by execution flow, reducing error propagation compared to generic CoT.
+CHASE-SQL [76] extends this with multi-path reasoning for candidate generation using “divide and conquer” CoT strategy.
+2) Consistency-driven reasoning: These methods improve reliability through candidate diversity and agreement metrics before execution.
+Self-consistency [125] is employed in frameworks like C3 [28], DAIL-SQL [8], and SuperSQL [57], where majority voting selects the most frequent SQL candidate based on syntax or structural similarity.
+SD+SA+Voting [64] refines this by filtering candidates using schema-aware rules (e.g., valid joins, clause ordering) before voting.
+Unlike execution-dependent methods, these strategies prioritize logical consistency and syntactic validity over runtime feedback, making them lightweight but less robust to semantic errors.
+3) Program-based and tool-augmented reasoning: Hybrid techniques combine SQL generation with external computation or tools.
+SQL-CRAFT [74] employs Program-of-Thoughts [126] (PoT) prompting, requiring LLMs to generate Python code alongside SQL for arithmetic-heavy queries, improving performance on datasets like BIRD [1].
+FUXI [75] integrates dedicated tools for schema linking and syntax checking, dynamically invoking them during generation.
+These approaches differ from pure prompting strategies by explicitly bridging LLM reasoning with external symbolic systems.
+IV-A5 $\mathbf{C}_{4}$-Execution Refinement
+Execution refinement methods improve SQL accuracy by leveraging database feedback through regeneration or preferred SQL selection. These approaches are categorized into two paradigms:
+1) Feedback-driven regeneration: These methods iteratively refine SQL queries using execution errors or intermediate results.
+Self-Debugging [79] enables LLMs to autonomously correct errors via natural language explanations of failed executions, while DESEM [62] uses error-type-specific prompts with termination criteria to prevent infinite loops.
+Multi-agent frameworks like MAC-SQL [48] and CHASE-SQL [76] utilize agents (e.g., Refiner, Query Fixer) to diagnose SQLite errors and regenerate SQL fixes.
+SQL-CRAFT [74] introduces adaptive control to balance correction depth, preventing over-refinement.
+Schema-focused methods like RSL-SQL [59] and DIN-SQL [7] use generic or gentle prompts to help LLMs identify schema misalignments during regeneration.
+Knowledge-to-SQL [69] optimizes its DELLM model through preference learning from execution feedback, while FUXI [75] integrates error feedback into tool-based reasoning.
+E-SQL [80] enhances questions with execution-aware candidate predictions.
+2) Execution-guided selection: This strategy leverages database execution results to select or rank candidates.
+ReFoRCE [61] and PET-SQL [52] prune invalid SQL queries that fail execution, then apply bias-aware voting (e.g., difficulty-weighted or efficiency-prioritized as in MCS-SQL [55]).
+Probabilistic verification distinguishes methods: MRC-EXEC [77] minimizes Bayes risk by comparing execution outputs to expected results, and LEVER [78] trains a verifier to predict correctness likelihoods from execution traces.
+Hybrid approaches like the retrieval-augmented framework [65] use execution feedback for iterative SQL revision through semantic gap analysis, bridging generation and selection.
 IV-B Fine-tuning
-TABLE IV: Well-designed methods used in fine-tuning (FT) for LLM-based text-to-SQL.
+TABLE IV: Well-designed methods used in the FT paradigm for LLM-based text-to-SQL.
 The methods in each category are ordered by release time.
@@
 Symbol-LLM [82] CodeLLaMA [11] Nov-2023 ACL’24
-CodeS [10] StarCoder [13, 33] Feb-2024 SIGMOD’24
-StructLM [70] CodeLLaMA [13] Feb-2024 arXiv’24
-Decomposition DTS-SQL [71] Mistral* [13, 40] Feb-2024 arXiv’24
-Since supervised fine-tuning (SFT) is the mainstream approach in the LLMs training [29, 91], for open-source LLMs (e.g., LLaMA-2 [94], Gemma [113]), the most straightforward method to enable the model to adapt a specific domain quickly is to use collected domain label to perform SFT on the model.
-The SFT phase is typically the preliminary phase of the well-designed training framework [112, 114], as well as the fine-tuning of text-to-SQL.
-The auto-regressive generation process of SQL query $Y$ can be formulated as follows:
-$P_{\pi}(Y\mid\mathcal{P})=\prod_{k=1}^{n}P_{\pi}(y_{k}\mid\mathcal{P},Y_{1:k-1}),$ (6)
-where $Y=\{y_{1},y_{2},\ldots,y_{n}\}$ is an SQL query of length $n$, $y_{k}$ is the corresponding $k^{th}$ token of the SQL query, $Y_{1:k-1}$ is the prefix sequence of $Y$ ahead the token $y_{k}$.
-$P_{\pi}(y_{k}\mid\cdot)$ is a conditional probability of a LLM $\pi$ for generating the $k^{th}$ token of $Y$ base on the input prompt $\mathcal{P}$ and the prefix sequence.
-Given a basic open-source model $\pi^{0}$, the goal of SFT is obtain a model $\pi^{SFT}$ through minimizing the cross-entropy loss:
-$\mathcal{L}_{SFT}=-\sum_{k=1}^{n}\operatorname{log}P_{\pi^{0}}(\hat{y}_{k}=y_{k}\mid\mathcal{P},Y_{1:k-1}),$ (7)
-where $\hat{y}_{k}$ is the $k^{th}$ token of the generated SQL query $\hat{Y}$, and $Y$ is the corresponding ground-truth label.
-The SFT approach, which is also a vanilla fine-tuning approach for text-to-SQL has been widely adopted in text-to-SQL research for various open-source LLMs [10, 46, 9].
-Compared to in-context learning (ICL) approaches, fine-tuning paradigms are more inclined to be at a starting point in LLM-based text-to-SQL.
-Currently, several studies exploring a better fine-tuning method have been released.
-We categorize the well-designed fine-tuning methods in different groups based on their mechanisms, as shown in Tab.
+CodeS [24] StarCoder [11, 1, 42] Feb-2024 SIGMOD’24
+StructLM [83] CodeLLaMA [11] Feb-2024 COLM’24
+Dubo-SQL [70] ChatGPT [1] Apr-2024 arXiv’24
+Distillery [58] LLaMA-3.1* [1] Aug-2024 NeurIPS’24 Workshop
+XiYan-SQL [84] Qwen2.5-Coder [11, 1] Nov-2024 arXiv’24
+Multi-task Tuning SQL-LLaMA [48] CodeLLaMA [11, 1] Dec-2023 COLING’25
+DTS-SQL [86] Mistral* [11, 37] Feb-2024 EMNLP’24 Findings
+KaSLA [29] DeepSeek-Coder* [11, 1] Oct-2024 arXiv’24
+MSc-SQL [87] Gemma-2 [11, 1] Oct-2024 NAACL’25
+XiYan-SQL [84] Qwen2.5-Coder [11, 1] Nov-2024 arXiv’24
+ROUTE [88] Qwen2.5* [11, 1, 37, 36, 38] Feb-2025 ICLR’25
+Supervised Fine-tuning (SFT) is the widely-used approach within the FT paradigm for training LLMs [27, 111].
+For open-source LLMs, like LLaMA [115] and Qwen [116] series, the most straightforward way to enable the model to adapt to a specific domain is to apply SFT over collected domain-specific data.
+The SFT phase typically serves as an initial step in a well-designed training pipeline [127, 128], including LLM-based text-to-SQL training.
+Consider the prompt defined in Eq.
+4 and 5 as $P$, and let $Y=[y_{1},y_{2},\dots,y_{T}]$ represent the target ground-truth SQL query.
+The probability Pr for an LLM $\pi$ to generate $Y$ given $P$ is formally modeled as:
+$\text{Pr}_{\pi}\left(Y\mid P\right)=\prod_{t=1}^{T}\text{Pr}_{\pi}\left(y_{t}\mid P,y_{<t}\right).$ (6)
+Contrary to the LLM in the ICL paradigm, the parameters of $\pi$ are tuned. The SFT process minimizes the cross-entropy loss over the training dataset $\mathcal{D}$ for $\pi$:
+$\mathcal{L}_{\text{SFT}}=-\sum_{\left(P,\,Y\right)\in\mathcal{D}}\sum_{t=1}^{T}\log\text{Pr}_{\pi}\left(y_{t}\mid P,y_{<t}\right).$ (7)
+As a vanilla fine-tuning approach for LLM-based text-to-SQL, SFT is widely adopted in text-to-SQL research across various open-source LLMs [24, 119, 8].
+FT paradigms are often more foundational in LLM-based text-to-SQL than ICL, which is the mainstream approach in the community.
+This is largely due to the comparably limited capabilities of open-source LLMs versus proprietary models in understanding complex database content.
+For the existing well-designed methods in the FT paradigm, we group these methods into different categories based on their training mechanisms, as shown in Tab.
 IV.
-Enhanced Architecture
-The widely-used generative pre-trained transformer (GPT) framework utilizes decoder-only transformer architecture and conventional auto-regressive decoding for text generation.
-Recent studies on the efficiency of LLMs have revealed a common challenge: when generating long sequences with the auto-regressive paradigm, the need to incorporate the attention mechanism results in high latency for LLMs [115, 116].
-In LLM-based text-to-SQL, the speed of generating SQL queries is significantly slower compared to traditional language modeling [21, 28], which has become a challenge in constructing high-efficiency local NLIDB.
-As one of the solutions, CLLMs [69] are designed to address the above challenge with an enhanced model architecture and achieve a speedup for SQL generation.
-Data Augmentation
-During the fine-tuning process, the most straightforward factor affecting the model’s performance is the quality of the training labels [117].
-The fine-tuning under the low quality or lack of the training labels is “making bricks without straw”, using high-quality or augmented data for fine-tuning always surpasses the meticulous design of fine-tuning methods on low-quality or raw data [74, 29].
-Data-augmented fine-tuning in text-to-SQL made substantial progress, focusing on enhancing the data quality during the SFT process.
-DAIL-SQL [9] are designed as an in-context learning framework, utilizing a sampling strategy for better few-shot instances.
-Incorporating the sampled instances in the SFT process improves the performance of open-source LLMs.
-Symbol-LLM [50] propose injection and infusion stage for data augmented instruction tuning.
-CodeS [10] augmented the training data with bi-directional generation with the help of ChatGPT.
-StructLM [70] are trained on multiple struct knowledge tasks for improving overall capability.
-Pre-training
-Pre-training is a fundamental phase of the complete fine-tuning process, aimed at acquiring text generation capabilities through auto-regressive training on extensive data [118].
-Conventionally, the current powerful proprietary LLMs (e.g., ChatGPT [119], GPT-4 [86], Claude [120]) are pre-trained on hybrid corpus, which mostly benefit from the dialogue scenario that exhibits text generation capability [85].
-The code-specific LLMs (e.g., CodeLLaMA [121], StarCoder [122]) are pre-trained on code data [100], and the mixture of various programming languages enables the LLMs to generate code to meet with the user’s instruction [123].
-As a sub-task of code generation, the main challenge of SQL-specific pre-training technique is that the SQL/Database-related content occupies only a small portion of the entire pre-training corpus.
-Then, as a result, the open-source LLMs with comparatively limited comprehensive capacity (compared to ChatGPT, GPT-4) do not acquire a promising understanding of how to convert NL questions to SQL during their pre-training process.
-The pre-training phase of the CodeS [10] model consists of three stages of incremental pre-training.
-Starting from a basic code-specific LLM [122], CodeS are incre pre-trained on a hybrid training corpus, including SQL-related data, NL-to-Code data, and NL-related data.
+IV-B1 Enhanced Architecture
+The widely-used Generative Pre-trained Transformer (GPT) framework [103] employs a decoder-only transformer architecture and conventional auto-regressive decoding for text generation.
+With advancements in the GPT framework, enhancements in LLM-based text-to-SQL involve modifications of the standard transformer backbone.
+These modifications incorporate tailored architecture to efficiently handle the structural and syntactic complexities of SQL.
+In traditional text-to-SQL research, this tailored architecture has effectively enhanced PLM-based performance.
+Techniques include schema-aware encoding [20], optimized attention [129], and targeted decoding [130], often relying on graph structures.
+Most recently, the speed of generating SQL queries by LLMs is notably slower than traditional PLM-based methods [19, 26], posing a challenge to building high-efficiency local NLIDBs.
+As a solution, CLLMs [81] are designed to address this challenge with specialized decoding, offering speedup efficiency for SQL generation and reducing high latency.
+IV-B2 Pre-training
+Pre-training is a fundamental phase of the complete training process, aimed at acquiring text generation capabilities through auto-regressive training on extensive data [131].
+Code-specific LLMs (e.g., CodeLLaMA [85], StarCoder [132]) are pre-trained on code data, with diverse programming languages enabling them to generate code per user instructions [123, 133].
+For SQL-specific pre-training, a key challenge is that SQL/database-related content makes up only a small part of the pre-training corpus.
+As a result, the open-source LLMs do not acquire a solid understanding of converting NL questions to SQL during their pre-training.
+PLM-based methods propose initial attempts in pre-training paradigms, such as STAR [134] and multitask approaches [135], which improve SQL generation by leveraging SQL-guided and multitask pre-training strategies.
+For the most recent LLM-based text-to-SQL, the pre-training phase of the CodeS [24] model consists of three stages of incremental pre-training.
+Starting from a code-specific LLM [132], CodeS are incrementally pre-trained on a hybrid training corpus, including SQL-related data, NL-to-Code data, and NL-related data.
 The text-to-SQL understanding and performance are significantly improved.
-Decomposition
-Decomposing a task into multiple steps or using multiple models to solve the task is an intuitive solution for addressing a complex scenario, as we previously introduced in Sec.
-IV-A, ICL paradigm.
-The proprietary models utilized in ICL-based methods have a massive number of parameters that are not at the same parameter level as the open-source models used in fine-tuning methods.
-These models inherently possess the capability to perform assigned sub-tasks well (through mechanisms such as few-shot learning) [30, 57].
-Thus, to replicate the success of this paradigm in ICL methods, it is necessary to reasonably assign corresponding sub-tasks to open-source models (such as generating external knowledge, schema linking, and distilling the schema) for sub-task-specific fine-tuning and constructing the corresponding data for fine-tuning, thereby assisting in the final SQL generation.
-DTS-SQL [71] proposed a two-stage decomposed text-to-SQL fine-tuning framework and designed a schema-linking pre-generation task ahead of the final SQL generation.
-V Expectations
-Despite the significant advancements made in text-to-SQL research, there are still several challenges that need to be addressed.
+IV-B3 Data Augmentation
+During the fine-tuning process, the quality of training labels is the most straightforward factor affecting the model’s performance [136].
+Training with low-quality or absent labels is “making bricks without straw.” Using high-quality or augmented data for fine-tuning consistently outperforms the meticulously designed methods relying on low-quality or raw data [91, 27].
+Data-augmented fine-tuning in text-to-SQL has made substantial progress by focusing on enhancing data quality and efficiency during the SFT process.
+DAIL-SQL [8] is designed as an ICL framework that uses a sampling strategy to enhance few-shot instances.
+Incorporating these sampled instances in the SFT process boosts the performance of open-source LLMs.
+Symbol-LLM [82] proposes injection and infusion stages for data-augmented instruction tuning.
+CodeS [24] augments the training data with bi-directional generation aided by ChatGPT.
+StructLM [83] trains on multiple structured knowledge tasks to improve overall capability.
+Dubo-SQL [70] created concise prompt templates for fine-tuning, which reduces costs and simultaneously achieves record performance.
+Distillery [58] employs an iterative failure-aware approach, selecting new training samples based on error analysis.
+XiYan-SQL [84] uses basic-syntax SFT and generation-enhanced SFT, employing hybrid augmented training data to fine-tune the SQL generator.
+IV-B4 Multi-task Tuning
+Dividing a complex task into multiple tasks or utilizing multiple models is a straightforward approach to tackle complexity, as outlined in Sec.
+IV-A2 on ICL paradigms.
+Proprietary models in ICL-based methods boast a large number of parameters, surpassing the scale of open-source models used in fine-tuning.
+These models are adept at performing assigned sub-tasks effectively, facilitated by mechanisms like few-shot learning [28, 48].
+To replicate this success, it is essential to assign sub-tasks strategically to open-source models, such as generating external knowledge, schema linking, and execution-based refinement, with appropriate training.
+This involves constructing specific datasets for fine-tuning, ultimately aiding in the final SQL generation.
+SQL-LLaMA [48] is a fine-tuned LLM framework utilizing multi-agent collaboration.
+DTS-SQL [86] introduced a two-stage multi-task text-to-SQL fine-tuning framework, incorporating a schema-linking pre-generation task before the final SQL generation.
+Similarly, KaSLA [29] employs a two-stage framework using knapsack optimization as a fine-tuning strategy for the schema linking task.
+MSc-SQL [87] trained separate models for schema linking, SQL generation, and critiquing the predicted SQL query.
+The ROUTE [88] framework involved the SFT process over noise filtering, data synthesis, and continued writing, validating the effectiveness of the multi-task tuning pipeline through collaboration prompting.
+V Summary
+This section contrasts traditional and LLM-based approaches in addressing text-to-SQL challenges, focusing on ambiguity resolution, schema linking, complex SQL generation, and cross-domain adaptation.
+Additionally, we analyze differences between different categories of LLM-based approaches, while highlighting their strengths and trade-offs.
+V-A LLM-based Methods vs. Traditional Methods.
+V-A1 Linguistic Complexity and Ambiguity
+Traditional rule-based systems, like GRAPPA [23], decompose questions via syntactic templates but fail to resolve ambiguous phrasing.
+Neural models like RyenSQL [17] improve lexical matching but lack contextual reasoning.
+In contrast, LLM-based methods such as DAIL-SQL [8] leverage chain-of-thought prompting to infer dependencies stepwise, while C3-SQL [28] iteratively validates SQL drafts against paraphrased questions.
+While traditional methods offer lightweight disambiguation, their reliance on limited training data restricts their ability to handle unseen phrasings.
+LLM-based methods address this with broader pretraining, but their computational cost grows quadratically with context length, making them impractical for low-resource edge devices.
+A promising direction is a two-stage pipeline: use traditional methods for initial ambiguity detection, then invoke LLMs only for resolving flagged ambiguities.
+This hybrid approach balances efficiency and robustness.
+V-A2 Schema Understanding and Representation
+Early systems rank schema elements via string similarity or schema-agnostic embeddings, truncating complex relationships.
+LLM-based methods enhance schema linking through in-context learning and structured attention.
+For example, DIN-SQL [7] serializes schema into natural language prompts with foreign-key hints, while KaSLA [29] mirrors human prioritization of schema elements, reducing the common issues of omissions and redundancies in schema linking.
+While rule-based methods are interpretable, their rigidity limits adaptability.
+LLM-based methods excel at semantic linking but lack transparency: their attention weights may prioritize incorrect columns due to spurious correlations.
+To mitigate this, future systems could combine symbolic rules with LLM inference: first, apply rule-based filters to exclude irrelevant columns, then use LLMs to rank the remaining candidates.
+A hybrid approach that combines LLMs with domain-specific rules can offer a balanced trade-off between performance and resource consumption.
+V-A3 Rare and Complex SQL Operations
+Handling operations like nested sub-queries or window functions requires labor-intensive template engineering.
+For instance, previous rule-based methods [9] manually define 40+ SQL grammar rules to map question patterns, which could not be generalized to unseen formulations.
+Neural semantic parsers like PICARD [137] constrain decoding via transition-based incremental parsing, ensuring syntactically valid output by rejecting invalid tokens at each step.
+LLM-based methods generate rare SQLs by combining parametric knowledge from code-centric pretraining with few-shot demonstrations.
+For example, DAIL-SQL [8] injects prompts with examples of window functions, enabling the model to generalize to unseen formulations.
+Traditional methods like PICARD [137] guarantee syntactic validity via grammar-constrained decoders but cannot innovate on rare and complex SQL operations.
+LLM-based methods provide novel solutions but may produce invalid SQL.
+A synergistic approach would integrate LLM’s creativity with traditional validation: first, generate candidate queries with an LLM, then apply PICARD-like grammars to reject invalid tokens during beam search.
+This mirrors the “generate-then-verify” paradigm in program synthesis, where LLMs propose solutions and symbolic checkers enforce constraints.
+V-A4 Cross-Domain Generalization
+Rule-based systems require manual template adjustments per domain, while PLM-based models like S2SQL [138] rely on domain-specific fine-tuning.
+LLM-based methods like CodeS [24] adapt to niche domains (e.g., biomedical terminology) via code-centric pretraining.
+Traditional domain adaptation methods require moderate fine-tuning data to align latent spaces but provide precise control over domain-specific mappings.
+LLM-based methods enable few-shot adaptation but risk overgeneralization.
+A viable solution is using LLMs as meta-optimizers: train a small adapter network for LLM using domain-specific rules, allowing the LLM to dynamically adjust its linking strategy.
+This combines the efficiency of rule-based adaptation with the LLM’s capacity to learn cross-domain analogies.
+V-B Comparison within LLM-based Methods
+V-B1 Comparison within ICL paradigm
+The in-context learning (ICL) paradigm’s categories exhibit complementary strengths and limitations.
+Decomposition ($\mathbf{C}_{1}$, Sec.
+IV-A2) is effective at managing complex queries through task modularization but introduces cascading error risks from multi-stage pipelines.
+Prompt optimization ($\mathbf{C}_{2}$, Sec.
+IV-A3) enhances input quality through semantic alignment but faces challenges with inherent reasoning limitations.
+Reasoning enhancement ($\mathbf{C}_{3}$, Sec.
+IV-A4) improves logical coherence through structured thinking but significantly increases computational overhead.
+Execution refinement ($\mathbf{C}_{4}$, Sec.
+IV-A5) ensures runtime validity through feedback loops but relies heavily on database accessibility.
+Recent methods like SuperSQL [57] and ReFoRCE [61] illustrate a trend toward hybrid integration-combining decomposition’s structural rigor with prompt optimization’s efficiency, while layering reasoning enhancement and execution consistency.
+This convergence suggests that future ICL methods will increasingly unify cross-category techniques into cohesive frameworks, balancing accuracy, efficiency, and robustness through systematic collaboration.
+V-B2 Comparison within FT paradigm
+Fine-tuning (FT) approaches highlight the trade-off between specialization and generalization.
+Enhanced architectures such as CLLMs [81] reduce latency, while pre-training approaches like CodeS [24] offer robust SQL foundations.
+Consequently, both approaches are constrained by their demands on computational resources and hardware devices.
+Data augmentation (DAIL-SQL [8], XiYan-SQL [84]) improves data efficiency but risks overfitting to synthetic patterns.
+Multi-task tuning (ROUTE [88]) boosts multi-step reasoning but complicates training convergence.
+While current methods focus on isolated improvements, the field lacks holistic frameworks to combine architectural technical contributions.
+Future FT methods need choreographed pipelines that sequentially optimize pre-training objectives, data diversity strategies and task-specific adaptations.
+This orchestration will be crucial for open-source models to achieve the multifaceted capabilities of proprietary LLMs in text-to-SQL.
+V-B3 Comparing ICL and FT paradigm
+The paradigms present asymmetric value propositions.
+ICL methods leverage proprietary LLMs (e.g., GPT-4 [105]) for rapid adaptation but face constraints in controllability and domain customization.
+FT methods empower open-source models (e.g., CodeLLaMA [85]) with targeted SQL capabilities through fine-tuning but require substantial engineering for data curation and training optimization.
+Importantly, FT methods enhance foundational model competence—CodeS’s [24] incremental pre-training boosts question and schema understanding, while SQL-LLaMA’s [48] multi-agent tuning supports complex compositional reasoning.
+Nonetheless, ICL methods’ prompt-based flexibility remain unmatched for addressing emerging database environments without retraining.
+The future landscape will likely reveal a symbiotic relationship: FT develops specialized models with general SQL grounding, while ICL offers lightweight adaptation for downstream application.
+This division mirrors trends in a retrieval-augmented generation, where FT builds core capabilities, and ICL manages dynamic context integration.
+VI Expectations
+Despite the significant advancements made in LLM-based text-to-SQL research, there are still several challenges that need to be addressed.
 In this section, we discuss the remaining challenges that we expect to overcome in future work.
-V-A Robustness in Real-world Applications
+VI-A Robustness in Real-world Applications
 The text-to-SQL implemented by LLMs is expected to perform generalization and robustness across complex scenarios in real-world applications.
@@
 Since extending a large amount of the data by human annotation incurs high labor costs, designing data-augmentation methods to obtain more question-SQL pairs will support the LLM in data scarcity.
-Also, the adaptation study of fine-tuned open-source LLM to the local small-size dataset can be potentially beneficial.
-Furthermore, the extensions on multi-lingual [42, 124] and multi-modal scenarios [125] should be studied comprehensively in future research, which will benefit more language groups and help build more general database interfaces.
-V-B Computational Efficiency
+Besides that, the adaptation study of fine-tuned open-source LLM to local small-size datasets can be potentially beneficial.
+Furthermore, the extensions on multi-lingual [39, 43] and multi-modal scenarios [139] should be studied in future research, which will benefit more language groups and help build more general database interfaces.
+VI-B Computational Efficiency
 The computational efficiency is determined by the inference speed and the cost of computational resources, which is worth considering in both application and research work [81, 67].
@@
 For local LLMs, from a starting point [81], more speedup strategies can be studied in enhancing the model’s architecture in future exploration.
-V-C Data Privacy and Interpretability
+VI-C Data Privacy and Interpretability
 As a part of the LLMs’ study, LLM-based text-to-SQL also faces some general challenges present in LLM research [140, 141, 142].
@@
 Building on this, combining advanced study in interpretability [146, 147] to enhance text-to-SQL performance and interpreting the local model architecture from the database knowledge aspect remain future directions.
-V-D Extensions
+VI-D Extensions
 As a sub-field of LLMs and natural language understanding research, many studies in these fields have been adopted for text-to-SQL tasks, advancing its development [148, 125].

### v5

@@
 This system serves as a natural language interface to the database (NLIDB) using LLMs.
-Since SQL is one of the most widely used programming languages, with over half (51.52%) of professional developers employing it, but only around a third (35.29%) are systematically trained, [^1] https://survey.stackoverflow.co/2023 the NLIDB allows non-skilled users to access structured databases like professional database engineers [2, 3] and enhances human-computer interaction [4].
+Since SQL is one of the most widely used programming languages, with over half (51.52%) of professional developers employing it, but only around a third (35.29%) are systematically trained [^1] https://survey.stackoverflow.co/2023, the NLIDB allows non-skilled users to access structured databases like professional database engineers [2, 3] and enhances human-computer interaction [4].
 Furthermore, amid the research hotspot of LLMs, text-to-SQL can potentially mitigate prevalent hallucination issues by integrating realistic content from databases to fill LLMs’ knowledge gaps [5].
@@
 ReFoRCE [61] o1-preview* [42] EX Feb-2025 arXiv’25
+1.
 Zero-shot studies: The zero-shot-based ICL strategies primarily focus on studying the influence of prompt representation styles and the baseline performance of various LLMs for text-to-SQL [6, 25, 119].
@@
 Under zero-shot settings, the results demonstrate that ChatGPT has promising text-to-SQL performance compared to state-of-the-art PLM-based systems.
-Aside from the proprietary models discussed above, similar evaluations indicate that open-source LLMs can also perform zero-shot text-to-SQL tasks [112, 82, 119], particularly code-specific models [79, 119].
+Aside from the proprietary models discussed above, similar evaluations indicate that open-source LLMs can also perform zero-shot text-to-SQL tasks [112, 82, 119], particularly for the code-specific models [79, 119].
 Primary keys and foreign keys convey continuous knowledge across different tables, making them an effective prompting style in zero-shot scenarios.
@@
 To ensure fair comparability, [106] explores effective prompt construction for LLMs by studying different styles of prompt representation and highlights their findings on the crucial role of table structure, content, prompt length, and optimal representations in effective prompting.
+2.
 Few-shot studies: The study of few-shot prompting strategies focuses on the number and selection of few-shot instances.
@@
 Decomposition-based methods reduce text-to-SQL complexity by dividing tasks or questions into manageable components, employing two main strategies:
-1) Multi-module collaboration: This strategy breaks the task into specialized stages (e.g., schema linking, SQL generation, and SQL refinement) handled by distinct modules or agents.
+1.
+Multi-module collaboration: This strategy breaks the task into specialized stages (e.g., schema linking, SQL generation, and SQL refinement) handled by distinct modules or agents.
 C3 [28] employs a three-module pipeline involving schema linking via clear prompting, calibration bias prompting for SQL guidance, and consistency-based voting for final selection.
@@
 PET-SQL [52] refines outputs through multi-stage prompting, while MCS-SQL [55] uses execution-guided selection as the final module after multiple SQL generation.
-2) Question decomposition: Complexity is reduced by splitting user questions into intermediate sub-questions via question decomposition.
+2.
+Question decomposition: Complexity is reduced by splitting user questions into intermediate sub-questions via question decomposition.
 QDecomp [47] uses Least-to-Most prompting, breaking complex questions into step-by-step sub-questions to facilitate SQL generation.
@@
 Prompt optimization methods for text-to-SQL focus on enhancing prompt engineering quality. These methods are categorized into three key strategies:
-1) Advanced few-shot sampling: This strategy optimizes example selection through semantic similarity, diversity, or domain alignment.
+1.
+Advanced few-shot sampling: This strategy optimizes example selection through semantic similarity, diversity, or domain alignment.
 DESEM [62] and retrieval-augmented frameworks [65, 70, 58] de-semanticize questions to retrieve examples with matching intent skeletons or vector embeddings.
@@
 Dynamic selection is explored in ACT-SQL [67] and MCS-SQL [55], which adaptively retrieve examples based on defined similarity scores or masked similarity scores.
-2) Schema augmentation: These methods refine schema representations to reduce redundancy and improve relevance.
+2.
+Schema augmentation: These methods refine schema representations to reduce redundancy and improve relevance.
 C3 [28] distills question-relevant schemas by removing unrelated tables or columns, while RSL-SQL [59] simplifies schemas using NL descriptions.
@@
 QDecomp [47] and SD+SA+Voting [64] further augment prompts with schema-aware semantic and structural hints.
-3) External knowledge integration: This group enhances prompts with domain-specific knowledge or feedback mechanisms.
+3.
+External knowledge integration: This group enhances prompts with domain-specific knowledge or feedback mechanisms.
 Knowledge-to-SQL [69] trains a DELLM to inject external knowledge about question-database relationships, while ICRL [71] integrates reinforcement learning with schema-graph-derived knowledge bases.
@@
 These approaches focus on three crucial strategies:
-1) Chain-of-thought (CoT)-based reasoning: CoT-based methods guide LLMs to generate intermediate reasoning steps before predicting SQLs.
+1.
+Chain-of-thought (CoT)-based reasoning: CoT-based methods guide LLMs in generating intermediate reasoning steps before predicting SQLs.
 While standard CoT prompts (e.g., “Let’s think step by step”) show limited effectiveness in text-to-SQL [8, 1], adaptations help bridge this gap.
@@
 CHASE-SQL [76] extends this with multi-path reasoning for candidate generation using “divide and conquer” CoT strategy.
-2) Consistency-driven reasoning: These methods improve reliability through candidate diversity and agreement metrics before execution.
+2.
+Consistency-driven reasoning: These methods improve reliability through candidate diversity and agreement metrics before execution.
 Self-consistency [125] is employed in frameworks like C3 [28], DAIL-SQL [8], and SuperSQL [57], where majority voting selects the most frequent SQL candidate based on syntax or structural similarity.
@@
 Unlike execution-dependent methods, these strategies prioritize logical consistency and syntactic validity over runtime feedback, making them lightweight but less robust to semantic errors.
-3) Program-based and tool-augmented reasoning: Hybrid techniques combine SQL generation with external computation or tools.
+3.
+Program-based and tool-augmented reasoning: Hybrid techniques combine SQL generation with external computation or tools.
 SQL-CRAFT [74] employs Program-of-Thoughts [126] (PoT) prompting, requiring LLMs to generate Python code alongside SQL for arithmetic-heavy queries, improving performance on datasets like BIRD [1].
@@
 Execution refinement methods improve SQL accuracy by leveraging database feedback through regeneration or preferred SQL selection. These approaches are categorized into two paradigms:
-1) Feedback-driven regeneration: These methods iteratively refine SQL queries using execution errors or intermediate results.
+1.
+Feedback-driven regeneration: These methods iteratively refine SQL queries using execution errors or intermediate results.
 Self-Debugging [79] enables LLMs to autonomously correct errors via natural language explanations of failed executions, while DESEM [62] uses error-type-specific prompts with termination criteria to prevent infinite loops.
@@
 E-SQL [80] enhances questions with execution-aware candidate predictions.
-2) Execution-guided selection: This strategy leverages database execution results to select or rank candidates.
+2.
+Execution-guided selection: This strategy leverages database execution results to select or rank candidates.
 ReFoRCE [61] and PET-SQL [52] prune invalid SQL queries that fail execution, then apply bias-aware voting (e.g., difficulty-weighted or efficiency-prioritized as in MCS-SQL [55]).

### v6

@@
 However, as modern databases and user questions grow more complex, PLMs with a limited parameter size often produce incorrect SQL.
-This necessitates more sophisticated and tailored optimization methods, which restricts the application of PLM-based systems.
+This necessitates more sophisticated and tailored optimization methods, which restrict the application of PLM-based systems.
 Recently, large language models (LLMs) have shown significant capabilities in natural language understanding as model scale increases.
@@
 Subsequently, we present a systematic analysis of recent advances in LLM-based text-to-SQL.
-Finally, we make a summarization and discuss the remaining challenges in this field and suggest expectations for future research directions.
+Finally, we make a summary and discuss the remaining challenges in this field and suggest expectations for future research directions.
+All the related resources of LLM-based, including research papers, benchmarks, and open-source projects, are collected for the community in our repository: https://github.com/DEEP-PolyU/Awesome-LLM-based-Text2SQL.
 Index Terms:
@@
 This system serves as a natural language interface to the database (NLIDB) using LLMs.
-Since SQL is one of the most widely used programming languages, with over half (51.52%) of professional developers employing it, but only around a third (35.29%) are systematically trained [^1] https://survey.stackoverflow.co/2023, the NLIDB allows non-skilled users to access structured databases like professional database engineers [2, 3] and enhances human-computer interaction [4].
+According to the 2023 Stack Overflow survey, SQL is one of the most widely used programming languages, with over half of professional developers (51.52%) employing it, while only about a third (35.29%) having received systematic training.
+The NLIDB therefore allows non-skilled users to access structured databases like professional database engineers [2, 3] and enhances human-computer interaction [4].
 Furthermore, amid the research hotspot of LLMs, text-to-SQL can potentially mitigate prevalent hallucination issues by integrating realistic content from databases to fill LLMs’ knowledge gaps [5].
-The significant value of text-to-SQL have triggered a range of studies on its integration and optimization with LLMs [6, 7, 8]; consequently, LLM-based text-to-SQL remains a highly discussed research field in the NLP and database communities.
+The significant value of text-to-SQL has triggered a range of studies on its integration and optimization with LLMs [6, 7, 8]; consequently, LLM-based text-to-SQL remains a highly discussed research field in the NLP and database communities.
 Previous traditional methods have made notable progress in implementing text-to-SQL.
@@
 SQL generation: This involves incorporating the above parsing and then predicting the correct syntax to generate executable SQL queries that can retrieve the required answer.
-LLMs have proven to perform a successful vanilla implementation [6, 25], benefiting from the powerful semantic parsing capacity enabled by rich training corpus [26, 27].
+LLMs have proven to perform a successful vanilla implementation [6, 25], benefiting from the powerful semantic parsing capacity enabled by a rich training corpus [26, 27].
 Further studies on enhancing LLMs for question understanding [7, 8], schema comprehension [28, 29], and SQL generation [30] are being increasingly released.
@@
 SGU-SQL [30], MetaSQL [50], $\textit{{R}}^{3}$ [51], PET-SQL [52], PURPLE [53], TA-SQL [54], MCS-SQL [55],
-CHESS [56], SuperSQL [57], Distillery [58], RSL-SQL [59], Gen-SQL [60], ReFoRCE [61], modelnode-methods,text width=42.5em] ] [ Prompt
+CHESS [56], SuperSQL [57], Distillery [58], RSL-SQL [59], Gen-SQL [60], Spider-Agent [42],
+ReFoRCE [61], LinkAlign [62], modelnode-methods,text width=42.5em] ] [ Prompt
 Optimization, leaf-methods, text width=6.5em [DESEM+P [63], StructGPT [64], SD+SA+Voting [65], RAG+SP&DRC [66], C3 [28], DAIL-SQL [8],
@@
 MCS-SQL [55], SuperSQL [57], Distillery [58], RSL-SQL [59], ICRL [72], Gen-SQL [60],
-SAFE-SQL [72], ReFoRCE [61], modelnode-methods,text width=42.5em] ] [ Reasoning
+SAFE-SQL [73], ReFoRCE [61], LinkAlign [62], modelnode-methods,text width=42.5em] ] [ Reasoning
 Enhancement, leaf-methods, text width=6.5em[CoT [47, 1, 8, 30], QDecomp [47], Least-to-Most [47], SQL-PaLM [74], ACT-SQL [68], POT [75],
-SQL-CRAFT [74], FUXI [75], SuperSQL [57], CHASE-SQL [76], ReFoRCE [61], modelnode-methods,text width=42.5em] ] [ Execution
+SQL-CRAFT [75], FUXI [76], SuperSQL [57], CHASE-SQL [77], ReFoRCE [61], Spider-Agent [42], modelnode-methods,text width=42.5em] ] [ Execution
 Refinement, leaf-methods, text width=6.5em[MBR-Exec [78], Coder-Reviewer [46], LEVER [79], SELF-DEBUGGING [80], DESEM+P [63],
@@
 DELLM [70], SQL-CRAFT [75], FUXI [76], $\textit{{R}}^{3}$ [51], PET-SQL [52], PURPLE [53], Dubo-SQL [71],
-MCS-SQL [55], CHESS [56], Distillery [58], E-SQL [80], CHASE-SQL [76], RSL-SQL [59], modelnode-methods,text width=42.5em] ] ] [ Fine-tuning
+MCS-SQL [55], CHESS [56], Distillery [58], E-SQL [81], CHASE-SQL [77], RSL-SQL [59],
+Spider-Agent [42], ReFoRCE [61], modelnode-methods,text width=42.5em] ] ] [ Fine-tuning
 Paradigm
@@
 Augmentation, leaf-methods, text width=6.5em [DAIL-SQL [8], Symbol-LLM [83], CodeS [24], StructLM [84], Dubo-SQL [71], Distillery [58],
-XiYan-SQL [84], modelnode-methods,text width=42.5em] ] [ Multi-task
-Tuning, leaf-methods, text width=6.5em [SQL-LLaMA [85], DTS-SQL [86], KaSLA [29], MSc-SQL [87], XiYan-SQL [84], ROUTE [88], modelnode-methods,text width=42.5em] ] ] ] ]
+XiYan-SQL [85], SHARE [86], modelnode-methods,text width=42.5em] ] [ Multi-task
+Tuning, leaf-methods, text width=6.5em [SQL-LLaMA [87], DTS-SQL [88], KaSLA [29], MSc-SQL [89], XiYan-SQL [85], ROUTE [90],
+SHARE [86], modelnode-methods,text width=42.5em] ] ] ] ]
 Figure 3: Taxonomy tree of research in LLM-based text-to-SQL, taking inspiration from [91]; nodes are ordered by release time.
@@
 This categorization provides analysis from technical implementation perspectives, clearly indicating their contributions to generating better SQL.
-Then, we further integrate insights while analyzing the trade-offs between LLM-based and traditional methods, as well as between in-context learning and fine-tuning paradigms for future developments.
+Then, we further integrate insights while analyzing the trade-offs between LLM-based and traditional methods, as well as between in-context learning and fine-tuning paradigms for future research.
 Finally, we highlight the remaining challenges in robustness, real-world deployment, efficiency, and domain adaptation, offering actionable directions for future research.
 Fig.
-3 shows a taxonomy tree that summarizes the structure and contents of our survey.
+3 shows a taxonomy for the survey.
 II Overview
@@
 It has the potential to democratize data access by allowing users to interact with databases using natural language, thus eliminating the need for specialized SQL programming knowledge [95].
-This capability can greatly benefit domains such as business intelligence, customer support, and scientific research, by enabling non-specialized users to easily retrieve target content from databases and facilitating more efficient data analysis.
+This capability can greatly benefit domains such as business intelligence, customer support, and scientific research by enabling non-specialized users to easily retrieve target content from databases and facilitating more efficient data analysis.
 II-A Challenges in Text-to-SQL
@@
 BIRD [1] employs human database experts to annotate each text-to-SQL sample with external knowledge, categorized into numeric reasoning knowledge, domain knowledge, synonym knowledge, and value illustration.
-Similarly, Spider-DK [36] defines and adds five types of domain knowledge for a human-curated version of the Spider dataset [11]: SELECT columns mentioned by omission, simple inference required, synonyms substitution in cell value word, one non-cell value word generate a condition, and easy to conflict with other domains.
+Similarly, Spider-DK [36] defines and adds five types of domain knowledge for a human-curated version of the Spider dataset [11]: SELECT columns mentioned by omission, simple inference required, synonym substitution in cell value words, one non-cell value word generates a condition, and easy to conflict with other domains.
 Both studies found that human-annotated knowledge significantly improves SQL generation performance for samples requiring external domain knowledge.
@@
 SParC [40] and CoSQL [32] explore context-dependent SQL generation by constructing a conversational database querying system.
-Unlike traditional text-to-SQL datasets that only have a single question-SQL pair for one example, SParC decomposes the question-SQL examples in the Spider dataset into multiple sub-question-SQL pairs to construct a simulated and meaningful interaction, including inter-related sub-questions that aid SQL generation, and unrelated sub-questions that enhance data diversity.
+Unlike traditional text-to-SQL datasets that only have a single question-SQL pair for one example, SParC decomposes the question-SQL examples in the Spider dataset into multiple sub-question-SQL pairs to construct a simulated and meaningful interaction, including interrelated sub-questions that aid SQL generation, and unrelated sub-questions that enhance data diversity.
 CoSQL, in comparison, involves conversational interactions in natural language, simulating real-world scenarios to increase complexity and diversity.
@@
 Based on the generation ability of large-scale pre-trained language models, Dr.
-Spider [44] provided 17 different perturbation strategies on databases, natural language queries and SQL statements to measure the robustness of text-to-SQL system.
+Spider [44] provided 17 different perturbation strategies on databases, natural language queries and SQL statements to measure the robustness of text-to-SQL systems.
 These perturbations lead to significant drops in accuracy, as a text-to-SQL system with low robustness may be misled by incorrect matches between tokens in NL questions and database entities.
@@
 III-A6 Long-context dataset
-In real-world database applications, numerous challenging scenarios demand complex SQL queries with significant token lengths, often exceeding 100 tokens, involving multiple intricate operations.
+In real-world database applications, numerous challenging scenarios demand complex SQL queries with significant token lengths, often exceeding 100 tokens and involving multiple intricate operations.
 Spider 2.0 [42] exemplifies this complexity by presenting enterprise-level text-to-SQL problems that require advanced reasoning across various SQL queries and dialects.
-This benchmark reveals the limitations of current models, as demonstrated by lower success rates—with advanced LLMs solving only 17.1% of tasks—underscoring its unique challenges.
+Distinct from earlier benchmarks, Spider 2.0 contains numerous real-world tasks sourced from diverse production databases such as BigQuery and Snowflake, featuring large-scale, industrial schemas with hundreds or even thousands of columns, nested structures, and multiple schemas per instance.
+Each task requires navigating not only complex SQL logic, but also integrating information from project codebases, database metadata, and dialect-specific documentation, closely reflecting how data workflows operate in enterprise settings.
+Gold-standard queries in this benchmark average close to 150 tokens, far surpassing previous datasets in both length and complexity.
+Notably, all SQLs are carefully annotated and reviewed for both semantic fidelity and execution correctness, ensuring challenges are both realistic and robust.
+As a result, Spider 2.0 highlighted the substantial gap between present-day LLMs’ capabilities and the demands of authentic, long-context database scenarios.
 In parallel, BIRD-CRITIC [^1] https://bird-critic.github.io/ [1] introduces a diagnostic dimension, comprising 600 tasks designed to assess the efficacy of LLMs in resolving real-world user issues across diverse SQL dialects.
@@
 To address this gap, BULL [45] introduces a dedicated text-to-SQL benchmark for the financial sector.
-The dataset encompasses databases related to funds, stocks, and macroeconomic data, offering a comprehensive platform to evaluate and enhance text-to-SQL methods in financial applications, effectively tackling the financial domain unique text-to-SQL challenges.
+The dataset encompasses databases related to funds, stocks, and macroeconomic data, offering a comprehensive platform to evaluate and enhance text-to-SQL methods in financial applications, effectively tackling the financial domain’s unique text-to-SQL challenges.
 III-B Evaluation Metrics
@@
     $\mathbbm{1}(V_{n},\hat{V}_{n})=\begin{cases}1,V_{n}=\hat{V}_{n}\\ 0,V_{n}\neq\hat{V}_{n}\end{cases}$ (2)
-    Then, $\textbf{R}(Y_{n},\hat{Y}_{n})=\sqrt{E(Y_{n})/E(\hat{Y}_{n})}$ denotes the relative execution efficiency of the predicted SQL query in comparison to ground-truth query, where $E(\cdot)$ is the execution time of each SQL in the database.
+    Then, $\textbf{R}(Y_{n},\hat{Y}_{n})=\sqrt{E(Y_{n})/E(\hat{Y}_{n})}$ denotes the relative execution efficiency of the predicted SQL in comparison to the ground-truth, where $E(\cdot)$ is the execution time of each SQL in the database.
 BIRD benchmark [1] ensures the stability of this metric by computing the average of $\textbf{R}(Y_{n},\hat{Y}_{n})$ over 100 runs for each example.
-Most of the recent LLM-based text-to-SQL studies focus on BIRD and Spider-related datasets and execution-based metrics, the corresponding details can be found at Tab. III.
+Most of the recent LLM-based text-to-SQL studies focus on BIRD and Spider-related datasets and execution-based metrics. The corresponding details can be found at Tab. III.
 IV Methods
@@
 IV-A In-context Learning
-Through extensive and widely recognized research, ICL (also called prompt engineering) has been proven to play a decisive role in the performance of LLMs across various tasks [117, 26, 118].
+Through extensive and widely recognized research, ICL (also called prompt engineering) has been proven to play a decisive role in the performance of LLMs across various tasks [119, 26].
 Consequently, it also impacts SQL generation across different prompt styles [8, 120].
@@
 The index $i\in\{1,\ldots,k\}$ also serves as the corresponding serial number for $\mathcal{S}_{i}$, $Q_{i}$, and $Y_{i}$.
-The OpenAI demonstration [^3] The prompt format follows the official document provided by the OpenAI platform: https://platform.openai.com/examples/default-sql-translate provides an indicative example and regulation for the prompting process in LLM-based text-to-SQL [28].
+The OpenAI demonstration [^2] The prompt format follows the official document provided by the OpenAI platform: https://platform.openai.com/examples/default-sql-translate.
+provides an indicative example and regulation for the prompting process in LLM-based text-to-SQL [28].
 The prompt designs and formats utilized by a substantial portion of the ICL methods discussed in this survey are developed based on the OpenAI regulation.
@@
 III.
-*Note that SimpleDDL is a prompting style that provides only table and column names [119], rather than a method.
+*Note that SimpleDDL is a prompting style that provides only table and column names [120], rather than a tailored method.
 Methods Adopted by Applied LLMs
@@
 Methods that fall into multiple categories will be introduced accordingly.
-The dataset column consists of six text-to-SQL datasets, which can be clarified by names: Spider [11], BIRD [1], Spider-DK [36] Spider-SYN [37], Spider-Realistic [38], and Spider 2.0 [42].
+The dataset column consists of six text-to-SQL datasets, which can be clarified by names: Spider [11], BIRD [1], Spider-DK [36] Spider-SYN [37], Spider-Realistic [38], and Spider 2.0 [42]
+.
 We use the reference ID in the table due to column space limitations.
@@
 CoT method is introduced in multiple venues: NeurIPS’23 [1], EMNLP’23 [47], VLDB’24 [8], arXiv’24 [30].
-Methods Applied LLMs Dataset Metrics $\mathbf{C}_{1}$ $\mathbf{C}_{2}$ $\mathbf{C}_{3}$ $\mathbf{C}_{4}$ Release Time Publication Venue
-MBR-Exec [77] Codex [11] EX Apr-2022 EMNLP’22
-Coder-Reviewer [46] Codex [11] EX Nov-2022 ICML’23
-LEVER [78] Codex [11] EX Feb-2023 ICML’23
-Self-Debugging [79] StarCoder* [11] EX Apr-2023 ICLR’24
-DESEM+P [62] ChatGPT [11, 37] EX Apr-2023 PRICAI’23
-DIN-SQL [7] GPT-4* [1, 11, 42] EX, EM, VES Apr-2023 NeurIPS’23
-CoT [47, 1, 8, 30] GPT-4 [11, 38, 1] EX, VES May-2023 Multiple Venues
-StructGPT [63] ChatGPT* [11, 37, 38] EX May-2023 EMNLP’23
-SD+SA+Voting [64] ChatGPT* [11, 37, 38] EX May-2023 EMNLP’23 Findings
-QDecomp [47] Codex [11, 38] EX May-2023 EMNLP’23
-Least-to-Most [47] Codex [11] EX May-2023 EMNLP’23
-SQL-PaLM [73] PaLM-2 [11] EX May-2023 TMLR’24
-RAG+SP&DRC [65] ChatGPT [11] EX Jul-2023 ICONIP’23
-C3 [28] ChatGPT [11] EX Jul-2023 arXiv’23
-DAIL-SQL [8] GPT-4* [11, 1, 38, 42] EX, EM, VES Aug-2023 VLDB’24
-ODIS [66] Codex* [11] EX Oct-2023 EMNLP’23 Findings
-ACT-SQL [67] GPT-4* [11, 37] EX, EM Oct-2023 EMNLP’23 Findings
-MAC-SQL [48] GPT-4* [11, 1] EX, EM, VES Dec-2023 COLING’25
-DEA-SQL [49] GPT-4 [11] EX Feb-2024 ACL’24 Findings
-FUSED [68] ChatGPT* [11] EX Feb-2024 EMNLP’24 Findings
-DELLM [69] GPT-4* [11, 1] EX, VES Feb-2024 ACL’24 Findings
-SGU-SQL [30] GPT-4* [11, 1] EX, EM Feb-2024 arXiv’24
-PoT [74] GPT-4* [11, 1] EX Feb-2024 arXiv’24
-SQL-CRAFT [74] GPT-4* [11, 1] EX Feb-2024 arXiv’24
-FUXI [75] GPT-4* [1] EX Feb-2024 EMNLP’24
-MetaSQL [50] GPT-4* [11] EX, EM Feb-2024 ICDE’24
-$\textit{{R}}^{3}$ [51] GPT-4* [11, 1] EX Feb-2024 arXiv’24
-PET-SQL [52] GPT-4 [11] EX Mar-2024 arXiv’24
-PURPLE [53] GPT-4* [11, 37, 38] EX, EM Mar-2024 ICDE’24
-Dubo-SQL [70] ChatGPT* [1] EX Apr-2024 arXiv’24
-TA-SQL [70] GPT-4* [11, 1, 36, 38] EX Apr-2024 ACL’24 Findings
-MCS-SQL [55] GPT-4 [11, 1] EX, VES May-2024 COLING’25
-CHESS [56] GPT-4o* [11, 1, 42] EX, VES May-2024 arXiv’24
-SuperSQL [57] GPT-4* [11, 1] EX, EM, VES Jun-2024 VLDB’24
-Distillery [58] GPT-4o-Mini* [1] EX Aug-2024 NeurIP’24 Workshop
-E-SQL [80] GPT-4o* [11, 1] EX, VES Sep-2024 arXiv’24
-CHASE-SQL [76] Gemini-1.5* [11, 1] EX, VES Oct-2024 ICLR’25
-RSL-SQL [59] DeepSeek-Chat* [11, 1] EX, VES Nov-2024 arXiv’24
-ICRL [71] ChatGPT [11] EX Jan-2025 COLING’25
-Gen-SQL [60] LLaMA-3* [11, 1] EX, EM, VES Jan-2025 COLING’25
-SAFE-SQL [72] GPT-4o-Mini* [11] EX, EM Feb-2025 arXiv’25
-ReFoRCE [61] o1-preview* [42] EX Feb-2025 arXiv’25
+Methods Applied LLMs Dataset Metrics $\mathbf{C}_{1}$ $\mathbf{C}_{2}$ $\mathbf{C}_{3}$ $\mathbf{C}_{4}$ Release Time Publication Venue MBR-Exec [78] Codex [11] EX Apr-2022 EMNLP’22 Coder-Reviewer [46] Codex [11] EX Nov-2022 ICML’23 LEVER [79] Codex [11] EX Feb-2023 ICML’23 Self-Debugging [80] StarCoder* [11] EX Apr-2023 ICLR’24 DESEM+P [63] ChatGPT [11, 37] EX Apr-2023 PRICAI’23 DIN-SQL [7] GPT-4* [1, 11, 42] EX, EM, VES Apr-2023 NeurIPS’23 CoT [47, 1, 8, 30] GPT-4 [11, 38, 1] EX, VES May-2023 Multiple Venues StructGPT [64] ChatGPT* [11, 37, 38] EX May-2023 EMNLP’23 SD+SA+Voting [65] ChatGPT* [11, 37, 38] EX May-2023 EMNLP’23 Findings QDecomp [47] Codex [11, 38] EX May-2023 EMNLP’23 Least-to-Most [47] Codex [11] EX May-2023 EMNLP’23 SQL-PaLM [74] PaLM-2 [11] EX May-2023 TMLR’24 RAG+SP&DRC [66] ChatGPT [11] EX Jul-2023 ICONIP’23 C3 [28] ChatGPT [11] EX Jul-2023 arXiv’23 DAIL-SQL [8] GPT-4* [11, 1, 38, 42] EX, EM, VES Aug-2023 VLDB’24 ODIS [67] Codex* [11] EX Oct-2023 EMNLP’23 Findings ACT-SQL [68] GPT-4* [11, 37] EX, EM Oct-2023 EMNLP’23 Findings MAC-SQL [48] GPT-4* [11, 1] EX, EM, VES Dec-2023 COLING’25 DEA-SQL [49] GPT-4 [11] EX Feb-2024 ACL’24 Findings FUSED [69] ChatGPT* [11] EX Feb-2024 EMNLP’24 Findings DELLM [70] GPT-4* [11, 1] EX, VES Feb-2024 ACL’24 Findings SGU-SQL [30] GPT-4* [11, 1] EX, EM Feb-2024 ICML’25 PoT [75] GPT-4* [11, 1] EX Feb-2024 arXiv’24 SQL-CRAFT [75] GPT-4* [11, 1] EX Feb-2024 arXiv’24 FUXI [76] GPT-4* [1] EX Feb-2024 EMNLP’24 MetaSQL [50] GPT-4* [11] EX, EM Feb-2024 ICDE’24 $\textit{{R}}^{3}$ [51] GPT-4* [11, 1] EX Feb-2024 ACL’25 Workshop PET-SQL [52] GPT-4 [11] EX Mar-2024 arXiv’24 PURPLE [53] GPT-4* [11, 37, 38] EX, EM Mar-2024 ICDE’24 Dubo-SQL [71] ChatGPT* [1] EX Apr-2024 arXiv’24 TA-SQL [54] GPT-4* [11, 1, 36, 38] EX Apr-2024 ACL’24 Findings MCS-SQL [55] GPT-4 [11, 1] EX, VES May-2024 COLING’25 CHESS [56] GPT-4o* [11, 1, 42] EX, VES May-2024 arXiv’24 SuperSQL [57] GPT-4* [11, 1] EX, EM, VES Jun-2024 VLDB’24 Distillery [58] GPT-4o-Mini* [1] EX Aug-2024 NeurIP’24 Workshop E-SQL [81] GPT-4o* [11, 1] EX, VES Sep-2024 arXiv’24 CHASE-SQL [77] Gemini-1.5* [11, 1] EX, VES Oct-2024 ICLR’25 RSL-SQL [59] DeepSeek-V3* [11, 1] EX, VES Nov-2024 arXiv’24 Spider-Agent [42] o1-preview* [42] EX Nov-2024 ICLR’25 ICRL [72] ChatGPT [11] EX Jan-2025 COLING’25 Gen-SQL [60] LLaMA-3* [11, 1] EX, EM, VES Jan-2025 COLING’25 SAFE-SQL [73] GPT-4o-Mini* [11] EX, EM Feb-2025 arXiv’25 ReFoRCE [61] o1-preview* [42] EX Feb-2025 ICLR’25 Workshop LinkAlign [62] DeepSeek-R1* [1, 42] EX, EM Mar-2025 arXiv’25
 1.
@@
 PET-SQL [52] refines outputs through multi-stage prompting, while MCS-SQL [55] uses execution-guided selection as the final module after multiple SQL generation.
+With the increasing complexity of recently proposed benchmarks, agentic frameworks like ReFoRCE [61], LinkAlign [62] and Spider-Agent [42] leveraged well-designed step-by-step action plans for LLMs to systematically reason toward the final SQL.
 2.
@@
 QDecomp [47] and SD+SA+Voting [65] further augment prompts with schema-aware semantic and structural hints.
+For more complex databases, LinkAlign [62] introduces a schema-based question rewriting strategy, combined with a retrieve-then-filter schema selection mechanism for accurate linking.
 3.
@@
 3.
-Program-based and tool-augmented reasoning: Hybrid techniques combine SQL generation with external computation or tools.
+Tool-augmented and agentic reasoning: Hybrid techniques combine SQL generation with external computation or tools.
 SQL-CRAFT [75] employs Program-of-Thoughts [127] (PoT) prompting, requiring LLMs to generate Python code alongside SQL for arithmetic-heavy queries, improving performance on datasets like BIRD [1].
@@
 These approaches differ from pure prompting strategies by explicitly bridging LLM reasoning with external symbolic systems.
+Also, recent frameworks such as ReFoRCE [61] and Spider-Agent [42] adopt step-by-step agentic reasoning, where LLMs act as specialized agents that iteratively plan actions to handle complex database scenarios.
 IV-A5 $\mathbf{C}_{4}$-Execution Refinement
@@
 E-SQL [81] enhances questions with execution-aware candidate predictions.
+Agentic frameworks [42] also incorporate feedback-driven regeneration, where specialized agents iteratively analyze execution feedback to plan corrective actions to refine SQLs.
 2.
@@
 XiYan-SQL [85] Qwen2.5-Coder [11, 1] Nov-2024 arXiv’24
+SHARE [86] LLaMA-3.1* [11, 1, 36, 38] Jun-2025 ACL’25
 Multi-task Tuning SQL-LLaMA [48] CodeLLaMA [11, 1] Dec-2023 COLING’25
@@
 ROUTE [90] Qwen2.5* [11, 1, 37, 36, 38] Feb-2025 ICLR’25
+SHARE [86] LLaMA-3.1* [11, 1, 36, 38] Jun-2025 ACL’25
 Supervised Fine-tuning (SFT) is the widely-used approach within the FT paradigm for training LLMs [27, 113].
@@
 XiYan-SQL [85] uses basic-syntax SFT and generation-enhanced SFT, employing hybrid augmented training data to fine-tune the SQL generator.
+SHARE [86] designed a rule-based error perturbation to augment schema linking and final SQL inference based on the decomposed trajectories.
 IV-B4 Multi-task Tuning
@@
 DTS-SQL [88] introduced a two-stage multi-task text-to-SQL fine-tuning framework, incorporating a schema-linking pre-generation task before the final SQL generation.
-Similarly, KaSLA [29] employs a two-stage framework using knapsack optimization as a fine-tuning strategy for the schema linking task.
+Similarly, KaSLA [29] employed a two-stage framework using knapsack optimization as a fine-tuning strategy for the schema linking task, and SHARE [86] adopted three sequential models for SQL correction.
 MSc-SQL [89] trained separate models for schema linking, SQL generation, and critiquing the predicted SQL query.
@@
 LLM-based methods provide novel solutions but may produce invalid SQL.
+The most current agentic reasoning [42, 61] mitigates invalid outputs through multi-agent actions, but still suffers from error accumulation.
 A synergistic approach would integrate LLM’s creativity with traditional validation: first, generate candidate queries with an LLM, then apply PICARD-like grammars to reject invalid tokens during beam search.
@@
 IV-A5) ensures runtime validity through feedback loops but relies heavily on database accessibility.
-Recent methods like SuperSQL [57] and ReFoRCE [61] illustrate a trend toward hybrid integration-combining decomposition’s structural rigor with prompt optimization’s efficiency, while layering reasoning enhancement and execution consistency.
+Recent methods like SuperSQL [57], Spider-Agent [42], and ReFoRCE [61] illustrate a trend toward hybrid integration-combining decomposition’s structural rigor with prompt optimization’s efficiency, while layering reasoning enhancement and execution consistency, which are implemented through well-designed agentic reasoning actions.
 This convergence suggests that future ICL methods will increasingly unify cross-category techniques into cohesive frameworks, balancing accuracy, efficiency, and robustness through systematic collaboration.
@@
 FT methods empower open-source models (e.g., CodeLLaMA [87]) with targeted SQL capabilities through fine-tuning but require substantial engineering for data curation and training optimization.
-Importantly, FT methods enhance foundational model competence—CodeS’s [24] incremental pre-training boosts question and schema understanding, while SQL-LLaMA’s [48] multi-agent tuning supports complex compositional reasoning.
+Importantly, FT methods enhance foundational model competence: CodeS’s [24] incremental pre-training boosts question and schema understanding, while SQL-LLaMA’s [48] multi-agent tuning supports complex compositional reasoning.
 Nonetheless, ICL methods’ prompt-based flexibility remain unmatched for addressing emerging database environments without retraining.
@@
 As reported in the corresponding evaluations of the dataset with synonym and incomplete instruction [6, 47], the SQL queries generated by ChatGPT contain around 40% incorrect execution, which is 10% lower than the original evaluation [47].
+For the evaluation in real-world enterprise tasks [42], current state-of-the-art models solve only about 21.3% of these tasks, far below their performance on earlier benchmarks [11, 1].
 Simultaneously, the fine-tuning with local text-to-SQL datasets may contain non-standardized samples and labels.
@@
 The computational efficiency is determined by the inference speed and the cost of computational resources, which is worth considering in both application and research work [82, 68].
-With the increasing complexity of databases in up-to-date text-to-SQL benchmarks [13, 1], databases will carry more information (including more tables and columns), and the token length of the database schema will correspondingly increase, raising a series of challenges.
+With the increasing complexity of databases in up-to-date text-to-SQL benchmarks [13, 1, 42], databases will carry more information (including more tables and columns), and the token length of the database schema will correspondingly increase, raising a series of challenges.
 Dealing with an ultra-complex database, taking the corresponding schema as input may encounter the challenge that the cost of calling proprietary LLMs will significantly increase, potentially exceeding the model’s maximum token length, especially with the implementation of open-source models that have shorter context lengths.
@@
 Although the in-context learning paradigm achieves promising accuracy, as a computational efficiency concern, the well-designed methods with the multi-stage framework or extended context increasing the number of API calls to enhance performance has simultaneously led to a substantial rise in costs [7].
-As reported in related approaches [67], a trade-off between performance and computational efficiency should be considered carefully, and designing a comparable (even better) in-context learning method with less API cost will be a practical implementation and is still under exploration.
+Recent agentic reasoning approaches, relying on iterative multi-step actions, typically multiply API calls severalfold.
+While enabling complex task handling and improved robustness, optimizing action planning to reduce redundant interactions across diverse database scenarios remains an important direction for future research.
+As reported in the related study [68], a trade-off between performance and computational efficiency should be considered carefully, and designing a comparable (even better) in-context learning method with less API cost will be a practical implementation and is still under exploration.
 Compared to PLM-based methods, the inference speed of LLM-based methods is observably slower [26, 19].
@@
 The approaches with a decomposition phase explain the text-to-SQL implementation process from the perspective of step-by-step generation [7, 47].
-Building on this, combining advanced study in interpretability [146, 147] to enhance text-to-SQL performance and interpreting the local model architecture from the database knowledge aspect remain future directions.
+Building on this, combining advanced study in interpretability [147, 148] to enhance text-to-SQL interpreting the model architecture from the database knowledge aspect remains a future direction.
 VI-D Extensions

### v7

@@
 All the related resources of LLM-based, including research papers, benchmarks, and open-source projects, are collected for the community in our repository: https://github.com/DEEP-PolyU/Awesome-LLM-based-Text2SQL.
-Index Terms:
-text-to-SQL, large language models, database, natural language understanding
 I Introduction
-Figure 1: An example of LLM-based text-to-SQL is selected from the BIRD [1] dataset.
+Figure 1: An example of LLM-based text-to-SQL is selected from the BIRD [li2023BIRD] dataset.
 A user asks a question about football leagues.
@@
 According to the 2023 Stack Overflow survey, SQL is one of the most widely used programming languages, with over half of professional developers (51.52%) employing it, while only about a third (35.29%) having received systematic training.
-The NLIDB therefore allows non-skilled users to access structured databases like professional database engineers [2, 3] and enhances human-computer interaction [4].
-Furthermore, amid the research hotspot of LLMs, text-to-SQL can potentially mitigate prevalent hallucination issues by integrating realistic content from databases to fill LLMs’ knowledge gaps [5].
-The significant value of text-to-SQL has triggered a range of studies on its integration and optimization with LLMs [6, 7, 8]; consequently, LLM-based text-to-SQL remains a highly discussed research field in the NLP and database communities.
+The NLIDB therefore allows non-skilled users to access structured databases like professional database engineers [wang2022proton, qin2022survey] and enhances human-computer interaction [xu2020autoqa].
+Furthermore, amid the research hotspot of LLMs, text-to-SQL can potentially mitigate prevalent hallucination issues by integrating realistic content from databases to fill LLMs’ knowledge gaps [lin2021truthfulqa].
+The significant value of text-to-SQL has triggered a range of studies on its integration and optimization with LLMs [rajkumar2022evaluating, pourreza2023dinsql, gao2023dailsql]; consequently, LLM-based text-to-SQL remains a highly discussed research field in the NLP and database communities.
 Previous traditional methods have made notable progress in implementing text-to-SQL.
@@
 2, these implementations have undergone a long evolutionary process.
-Early efforts were mostly based on well-designed rules and templates [9], specifically suitable for simple database scenarios.
-In recent years, due to the heavy labor costs [10] associated with rule-based methods and the growing complexity of database environments [11, 12, 13], designing a rule or template for each scenario has become increasingly difficult and impractical.
-The development of deep neural networks has advanced the progress of text-to-SQL [14, 15], which can automatically learn a mapping from the user question to its corresponding SQL [16, 17].
-Subsequently, pre-trained language models (PLMs) with strong semantic parsing capabilities have become the new paradigm for text-to-SQL systems [18], boosting their performance to a new level [19, 20, 21].
-Incremental research on PLM-based optimization, such as table content encoding [18, 17, 22] and pre-training [18, 23], has further advanced this field.
-Most recently, LLM-based approaches have implemented text-to-SQL through the in-context learning [7] and fine-tuning [24] paradigms, reaching state-of-the-art accuracy with well-designed frameworks and stronger comprehension capabilities compared to PLM-based methods.
+Early efforts were mostly based on well-designed rules and templates [li2014constructing], specifically suitable for simple database scenarios.
+In recent years, due to the heavy labor costs [mahmud2015rule] associated with rule-based methods and the growing complexity of database environments [yu2018Spider, zhong2017WikiSQL, pourreza2023evaluating], designing a rule or template for each scenario has become increasingly difficult and impractical.
+The development of deep neural networks has advanced the progress of text-to-SQL [sutskever2014sequence, vaswani2017attention], which can automatically learn a mapping from the user question to its corresponding SQL [hui2021improving, choi2021ryansql].
+Subsequently, pre-trained language models (PLMs) with strong semantic parsing capabilities have become the new paradigm for text-to-SQL systems [yin2020tabert], boosting their performance to a new level [li2023resdsql, li2023graphix, rai2023improving].
+Incremental research on PLM-based optimization, such as table content encoding [yin2020tabert, choi2021ryansql, lyu2020hybrid] and pre-training [yin2020tabert, yu2021grappa], has further advanced this field.
+Most recently, LLM-based approaches have implemented text-to-SQL through the in-context learning [pourreza2023dinsql] and fine-tuning [li2024codes] paradigms, reaching state-of-the-art accuracy with well-designed frameworks and stronger comprehension capabilities compared to PLM-based methods.
 The overall implementation details of LLM-based text-to-SQL can be divided into three aspects: 1.
@@
 SQL generation: This involves incorporating the above parsing and then predicting the correct syntax to generate executable SQL queries that can retrieve the required answer.
-LLMs have proven to perform a successful vanilla implementation [6, 25], benefiting from the powerful semantic parsing capacity enabled by a rich training corpus [26, 27].
-Further studies on enhancing LLMs for question understanding [7, 8], schema comprehension [28, 29], and SQL generation [30] are being increasingly released.
+LLMs have proven to perform a successful vanilla implementation [rajkumar2022evaluating, liu2023comprehensive], benefiting from the powerful semantic parsing capacity enabled by a rich training corpus [yang2024harnessing, zhao2023survey].
+Further studies on enhancing LLMs for question understanding [pourreza2023dinsql, gao2023dailsql], schema comprehension [dong2023c3, yuan2025knapsack], and SQL generation [zhang2024sgusql] are being increasingly released.
 Figure 2: A sketch of text-to-SQL research evolution outlines the advancement of implementation paradigms, each represented with different color backgrounds.
@@
 However, from 2023 onward, the emphasis shifts to the rapid advancements driven by LLMs, marking a significant acceleration in the field.
-{forest}
-forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-draw, rounded corners, align=left, minimum width=1em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=0pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , [ LLM-based Text-to-SQL, leaf-head, ver [ Datasets
-III-A), leaf-datasets, text width=5em [ Original
-Datasets
-Post-annotated
-Datasets, leaf-datasets, text width=6em [ Cross-
-domain, leaf-datasets, text width=5.5em [BIRD [1], DuSQL [31], CoSQL [32], Spider [11], WikiSQL [12], KaggleDBQA [33], ADVETA [34],
-Spider-SS [35], Spider-CG [35], Spider-DK [36], Spider-SYN [37], Spider-Realistic [38], CSpider [39],
-SParC [40], modelnode-datasets, text width=43.5em] ] [ Knowledge-
-augmented, leaf-datasets, text width=5.5em [BIRD [1], SQUALL [41], Spider-DK [36], BIRD-CRITIC [1], Spider 2.0 [42], modelnode-datasets, text width=43.5em] ] [ Cross-
-lingual, leaf-datasets, text width=5.5em [DuSQL [31], CSpider [39], Spider-Vietnamese [43], modelnode-datasets, text width=43.5em] ] [ Context-
-dependent, leaf-datasets, text width=5.5em [CoSQL [32], Spider-SS [35], Spider-CG [35], SparC [40], modelnode-datasets, text width=43.5em] ] [ Robustness, leaf-datasets, text width=5.5em [ADVETA [34], Spider-SYN [37], Spider-Realistic [38], Dr.
-Spider [44], modelnode-datasets, text width=43.5em] ] [ Long-context, leaf-datasets, text width=5.5em [BIRD-CRITIC [1], Spider 2.0 [42], modelnode-datasets, text width=43.5em] ] [ Specialized-
-domain, leaf-datasets, text width=5.5em [BULL [45], modelnode-datasets, text width=43.5em] ] ] ] [ Evaluation
-Metrics
-III-B), leaf-metrics,text width=5em [ Content-
-Matching
-based, leaf-metrics, text width=6em [Component Matching (CM) [11], Exact Matching (EM) [11], modelnode-metrics, text width=50.7em] ] [ Execution
-based, leaf-metrics, text width=6em [Execution Accuracy (EX) [11], Valid Efficiency Score (VES) [1], modelnode-metrics, text width=50.7em] ] ] [ Methods
-IV), leaf-methods,text width=5em [ In-context
-Learning
-Paradigm
-IV-A), leaf-methods, text width=6em [ Decomposition, leaf-methods, text width=6.5em [Coder-Reviewer [46], DIN-SQL [7], QDecomp [47], C3 [28], MAC-SQL [48], DEA-SQL [49],
-SGU-SQL [30], MetaSQL [50], $\textit{{R}}^{3}$ [51], PET-SQL [52], PURPLE [53], TA-SQL [54], MCS-SQL [55],
-CHESS [56], SuperSQL [57], Distillery [58], RSL-SQL [59], Gen-SQL [60], Spider-Agent [42],
-ReFoRCE [61], LinkAlign [62], modelnode-methods,text width=42.5em] ] [ Prompt
-Optimization, leaf-methods, text width=6.5em [DESEM+P [63], StructGPT [64], SD+SA+Voting [65], RAG+SP&DRC [66], C3 [28], DAIL-SQL [8],
-ODIS [67], ACT-SQL [68], FUSED [69], DELLM [70], Dubo-SQL [71], TA-SQL [54],
-MCS-SQL [55], SuperSQL [57], Distillery [58], RSL-SQL [59], ICRL [72], Gen-SQL [60],
-SAFE-SQL [73], ReFoRCE [61], LinkAlign [62], modelnode-methods,text width=42.5em] ] [ Reasoning
-Enhancement, leaf-methods, text width=6.5em[CoT [47, 1, 8, 30], QDecomp [47], Least-to-Most [47], SQL-PaLM [74], ACT-SQL [68], POT [75],
-SQL-CRAFT [75], FUXI [76], SuperSQL [57], CHASE-SQL [77], ReFoRCE [61], Spider-Agent [42], modelnode-methods,text width=42.5em] ] [ Execution
-Refinement, leaf-methods, text width=6.5em[MBR-Exec [78], Coder-Reviewer [46], LEVER [79], SELF-DEBUGGING [80], DESEM+P [63],
-DIN-SQL [7], SD+SA+Voting [65], SQL-PaLM [74], RAG+SP&DRC [66], C3 [28], MAC-SQL [48],
-DELLM [70], SQL-CRAFT [75], FUXI [76], $\textit{{R}}^{3}$ [51], PET-SQL [52], PURPLE [53], Dubo-SQL [71],
-MCS-SQL [55], CHESS [56], Distillery [58], E-SQL [81], CHASE-SQL [77], RSL-SQL [59],
-Spider-Agent [42], ReFoRCE [61], modelnode-methods,text width=42.5em] ] ] [ Fine-tuning
-Paradigm
-IV-B), leaf-methods,text width=6em [ Enhanced
-Architecture, leaf-methods, text width=6.5em [CLLMs [82], modelnode-methods,text width=42.5em] ] [ Pre-training, leaf-methods, text width=6.5em [CodeS [24], modelnode-methods,text width=42.5em] ] [ Data
-Augmentation, leaf-methods, text width=6.5em [DAIL-SQL [8], Symbol-LLM [83], CodeS [24], StructLM [84], Dubo-SQL [71], Distillery [58],
-XiYan-SQL [85], SHARE [86], modelnode-methods,text width=42.5em] ] [ Multi-task
-Tuning, leaf-methods, text width=6.5em [SQL-LLaMA [87], DTS-SQL [88], KaSLA [29], MSc-SQL [89], XiYan-SQL [85], ROUTE [90],
-SHARE [86], modelnode-methods,text width=42.5em] ] ] ] ]
-Figure 3: Taxonomy tree of research in LLM-based text-to-SQL, taking inspiration from [91]; nodes are ordered by release time.
-Despite the significant progress made in text-to-SQL research, several challenges remain that hinder the development of robust and generalized text-to-SQL systems [92].
-Recent related works have surveyed text-to-SQL systems using deep learning approaches and provided insights into previous deep neural network and PLM-based research [93, 3, 27, 94].
+[…]
+LLM-based Text-to-SQL
+Datasets III-A) Original Datasets & Post-annotated Datasets Cross- domain BIRD [li2023BIRD], DuSQL [wang2020DuSQL], CoSQL [yu2019CoSQL], Spider [yu2018Spider], WikiSQL [zhong2017WikiSQL], KaggleDBQA [lee2021KaggleDBQA], ADVETA [pi2022ADVETA], Spider-$S$ [gan2022Spider-CG], Spider-CG [gan2022Spider-CG], Spider-DK [gan2021Spider-DK], Spider-SYN [gan2021Spider-SYN], Spider-Realistic [deng2021Spider-Realistic], CSpider [min2019CSpider], SParC [yu2019SParC] Knowledge- augmented BIRD [li2023BIRD], SQUALL [shi2020SQUALL], Spider-DK [gan2021Spider-DK], BIRD-CRITIC [li2023BIRD], Spider 2.0 [lei2025spider2] Cross- lingual DuSQL [wang2020DuSQL], CSpider [min2019CSpider], Spider-Vietnamese [tuan2020pilot] Context- dependent CoSQL [yu2019CoSQL], Spider-$S$ [gan2022Spider-CG], Spider-CG [gan2022Spider-CG], SparC [yu2019SParC] Robustness ADVETA [pi2022ADVETA], Spider-SYN [gan2021Spider-SYN], Spider-Realistic [deng2021Spider-Realistic], Dr.
+Spider [chang2023dr] Long-context BIRD-CRITIC [li2023BIRD], Spider 2.0 [lei2025spider2] Specialized- domain BULL [zhang2024finsql] Evaluation Metrics III-B) Content- Matching based Component Matching (CM) [yu2018Spider], Exact Matching (EM) [yu2018Spider] Execution based Execution Accuracy (EX) [yu2018Spider], Valid Efficiency Score (VES) [li2023BIRD] Methods IV) In-context Learning Paradigm IV-A) Decomposition Coder-Reviewer [zhang2022coder], DIN-SQL [pourreza2023dinsql], QDecomp [tai2023exploring], C3 [dong2023c3], MAC-SQL [wang2024macsql], DEA-SQL [xie2024deasql], SGU-SQL [zhang2024sgusql], MetaSQL [fan2024metasql], $\textit{{R}}^{3}$ [xia2024r3], PET-SQL [li2024petsql], PURPLE [ren2024purple], TA-SQL [qu2024generation], MCS-SQL [lee2025mcs], CHESS [talaei2024chess], SuperSQL [li2024dawn], Distillery [maamari2024the], RSL-SQL [cao2024rsl], Gen-SQL [shi2025gen], Spider-Agent [lei2025spider2], ReFoRCE [deng2025reforce], LinkAlign [wang2025linkalign] Prompt Optimization DESEM+P [guo2023desem], StructGPT [jiang2023structgpt], SD+SA+Voting [nan2023enhancing], RAG+SP&DRC [guo2023ragsql], C3 [dong2023c3], DAIL-SQL [gao2023dailsql], ODIS [chang2023selective], ACT-SQL [zhang2023actsql], FUSED [wang2024fused], DELLM [hong2024knowledgetosql], Dubo-SQL [thorpe2024dubo], TA-SQL [qu2024generation], MCS-SQL [lee2025mcs], SuperSQL [li2024dawn], Distillery [maamari2024the], RSL-SQL [cao2024rsl], ICRL [toteja2025context], Gen-SQL [shi2025gen], SAFE-SQL [lee2025safe], ReFoRCE [deng2025reforce], LinkAlign [wang2025linkalign] Reasoning Enhancement CoT [tai2023exploring, li2023BIRD, gao2023dailsql, zhang2024sgusql], QDecomp [tai2023exploring], Least-to-Most [tai2023exploring], SQL-PaLM [sun2023sqlpalm], ACT-SQL [zhang2023actsql], POT [xia2024sqlcraft], SQL-CRAFT [xia2024sqlcraft], FUXI [gu2024fuxi], SuperSQL [li2024dawn], CHASE-SQL [pourreza2025chasesql], ReFoRCE [deng2025reforce], Spider-Agent [lei2025spider2] Execution Refinement MBR-Exec [shi2022natural], Coder-Reviewer [zhang2022coder], LEVER [ni2023lever], SELF-DEBUGGING [chen2024selfdebugging], DESEM+P [guo2023desem], DIN-SQL [pourreza2023dinsql], SD+SA+Voting [nan2023enhancing], SQL-PaLM [sun2023sqlpalm], RAG+SP&DRC [guo2023ragsql], C3 [dong2023c3], MAC-SQL [wang2024macsql], DELLM [hong2024knowledgetosql], SQL-CRAFT [xia2024sqlcraft], FUXI [gu2024fuxi], $\textit{{R}}^{3}$ [xia2024r3], PET-SQL [li2024petsql], PURPLE [ren2024purple], Dubo-SQL [thorpe2024dubo], MCS-SQL [lee2025mcs], CHESS [talaei2024chess], Distillery [maamari2024the], E-SQL [caferouglu2024esql], CHASE-SQL [pourreza2025chasesql], RSL-SQL [cao2024rsl], Spider-Agent [lei2025spider2], ReFoRCE [deng2025reforce] Fine-tuning Paradigm IV-B) Enhanced Architecture CLLMs [kou2024cllms] Pre-training CodeS [li2024codes] Data Augmentation DAIL-SQL [gao2023dailsql], Symbol-LLM [xu2024symbolllm], CodeS [li2024codes], StructLM [zhuang2024structlm], Dubo-SQL [thorpe2024dubo], Distillery [maamari2024the], XiYan-SQL [gao2024xiyan], SHARE [qu2025share] Multi-task Tuning SQL-LLaMA [roziere2023codellama], DTS-SQL [pourreza2024dtssql], KaSLA [yuan2025knapsack], MSc-SQL [gorti2024mscsql], XiYan-SQL [gao2024xiyan], ROUTE [qin2025route], SHARE [qu2025share]
+Figure 3: Taxonomy tree of research in LLM-based text-to-SQL, taking inspiration from [xu2023large]; nodes are ordered by release time.
+Despite the significant progress made in text-to-SQL research, several challenges remain that hinder the development of robust and generalized text-to-SQL systems [katsogiannis2023survey].
+Recent related works have surveyed text-to-SQL systems using deep learning approaches and provided insights into previous deep neural network and PLM-based research [deng2022recent, qin2022survey, zhao2023survey, zhang2024tabluarsurvey].
 In this survey, we aim to catch up with recent advances and provide a comprehensive review of the current state-of-the-art models and approaches in LLM-based text-to-SQL.
@@
 Text-to-SQL is a task that aims to convert NL questions into SQL queries that can be executed in a relational database.
-It has the potential to democratize data access by allowing users to interact with databases using natural language, thus eliminating the need for specialized SQL programming knowledge [95].
+It has the potential to democratize data access by allowing users to interact with databases using natural language, thus eliminating the need for specialized SQL programming knowledge [ma2021mt].
 This capability can greatly benefit domains such as business intelligence, customer support, and scientific research by enabling non-specialized users to easily retrieve target content from databases and facilitating more efficient data analysis.
@@
 II-A1 Linguistic Complexity and Ambiguity
-Natural language questions often contain complex linguistic representations, such as nested clauses, coreferences, and ellipses, which make it challenging to map them accurately to the corresponding parts of SQL queries [38].
-Additionally, natural language is inherently ambiguous, with multiple possible representations for a given user question [96, 97].
-Resolving these ambiguities and understanding the intent behind the user question requires deep natural language understanding and the capability to incorporate context and domain knowledge [1].
+Natural language questions often contain complex linguistic representations, such as nested clauses, coreferences, and ellipses, which make it challenging to map them accurately to the corresponding parts of SQL queries [deng2021Spider-Realistic].
+Additionally, natural language is inherently ambiguous, with multiple possible representations for a given user question [rajpurkar2016squad, rajpurkar2018know].
+Resolving these ambiguities and understanding the intent behind the user question requires deep natural language understanding and the capability to incorporate context and domain knowledge [li2023BIRD].
 II-A2 Schema Understanding and Representation
 To generate accurate SQL queries, text-to-SQL systems must have a comprehensive understanding of the database schema, including table names, column names, and relationships between various tables.
-Representing and encoding schema information is challenging because the database schema is often complex and varies significantly across different domains [11].
+Representing and encoding schema information is challenging because the database schema is often complex and varies significantly across different domains [yu2018Spider].
 II-A3 Rare and Complex SQL Operations
@@
 Models trained on a specific domain may not perform well on questions from other domains due to variations in vocabulary, database schema structure, and question patterns.
-Developing systems that can effectively generalize to new domains with minimal domain-specific training data or fine-tuning adaptation is a significant challenge [98].
+Developing systems that can effectively generalize to new domains with minimal domain-specific training data or fine-tuning adaptation is a significant challenge [yang2024unveiling].
 II-B Evolutionary Process
@@
 II-B1 Rule-based Methods
-Early text-to-SQL systems relied heavily on rule-based methods [9, 10, 23], where manually crafted rules and heuristics were used to map natural language questions to SQL queries.
+Early text-to-SQL systems relied heavily on rule-based methods [li2014constructing, mahmud2015rule, yu2021grappa], where manually crafted rules and heuristics were used to map natural language questions to SQL queries.
 Rule-based systems excel at generating syntactically correct SQL queries by relying on predefined templates and strict grammatical rules, ensuring adherence to SQL syntax and structure.
@@
 II-B2 Deep Learning-based Approaches
-With the rise of deep neural networks, sequence-to-sequence models and encoder-decoder structures, such as LSTMs [99] and transformers [15], were adapted to generate SQL queries from natural language input [100, 17].
-Typically, RYANSQL [17] introduced techniques like intermediate representations and sketch-based slot filling to handle complex questions and improve cross-domain generalization.
-Recently, researchers introduced graph neural networks (GNNs) for text-to-SQL tasks by leveraging schema dependency graphs to capture the relationships between database elements [16, 101].
+With the rise of deep neural networks, sequence-to-sequence models and encoder-decoder structures, such as LSTMs [hochreiter1997long] and transformers [vaswani2017attention], were adapted to generate SQL queries from natural language input [guo2019towards, choi2021ryansql].
+Typically, RYANSQL [choi2021ryansql] introduced techniques like intermediate representations and sketch-based slot filling to handle complex questions and improve cross-domain generalization.
+Recently, researchers introduced graph neural networks (GNNs) for text-to-SQL tasks by leveraging schema dependency graphs to capture the relationships between database elements [hui2021improving, xu2017sqlnet].
 Despite their flexibility, these methods often generate incorrect or incomplete SQL syntax, such as missing clauses or improper nesting, due to limited structural awareness.
@@
 PLMs have emerged as a powerful solution for text-to-SQL, leveraging the vast amounts of linguistic knowledge and semantic understanding captured during the pre-training process.
-The early adoption of PLMs in text-to-SQL primarily focused on fine-tuning off-the-shelf PLMs, such as BERT [102] and RoBERTa [103], on standard text-to-SQL datasets [11, 12].
+The early adoption of PLMs in text-to-SQL primarily focused on fine-tuning off-the-shelf PLMs, such as BERT [devlin2019bert] and RoBERTa [liu2019roberta], on standard text-to-SQL datasets [yu2018Spider, zhong2017WikiSQL].
 These PLMs, pre-trained on large amounts of training corpus, captured rich semantic representations and language understanding capabilities.
-By fine-tuning them on text-to-SQL tasks, researchers aimed to leverage the semantic and linguistic understanding of PLMs to generate accurate SQL queries [100, 18, 104].
+By fine-tuning them on text-to-SQL tasks, researchers aimed to leverage the semantic and linguistic understanding of PLMs to generate accurate SQL queries [guo2019towards, yin2020tabert, dou2022towards].
 Another line of research focuses on incorporating schema information into PLMs to improve their understanding of database structures and enable them to generate more executable SQL queries.
-Schema-aware PLMs are designed to capture the relationships and constraints present in the database structure [19].
+Schema-aware PLMs are designed to capture the relationships and constraints present in the database structure [li2023resdsql].
 While PLMs generate more accurate SQL than deep learning-based methods, they still face difficulties with complex operations like outer joins or aggregations, often producing syntactically flawed queries.
@@
 II-B4 LLM-based Implementation
-Large language models (LLMs), such as the GPT series [105, 106, 107], have gained significant attention in recent years due to their ability to generate coherent and fluent text.
-Researchers have begun exploring the potential of LLMs for text-to-SQL by leveraging their extensive knowledge reserves and superior generation capabilities [6, 8].
-These approaches often involve prompt engineering to guide proprietary LLMs in SQL generation [108] or fine-tuning open-source LLMs on text-to-SQL datasets [8].
+Large language models (LLMs), such as the GPT series [radford2018improving, brown2020gpt3, achiam2023gpt4], have gained significant attention in recent years due to their ability to generate coherent and fluent text.
+Researchers have begun exploring the potential of LLMs for text-to-SQL by leveraging their extensive knowledge reserves and superior generation capabilities [rajkumar2022evaluating, gao2023dailsql].
+These approaches often involve prompt engineering to guide proprietary LLMs in SQL generation [chang2023how] or fine-tuning open-source LLMs on text-to-SQL datasets [gao2023dailsql].
 The integration of LLMs in text-to-SQL is still an emerging research area with significant potential for further exploration and improvement.
-Researchers are investigating ways to better leverage the knowledge and reasoning capabilities of LLMs, incorporate domain-specific knowledge [1, 70], and develop more efficient fine-tuning strategies [24].
+Researchers are investigating ways to better leverage the knowledge and reasoning capabilities of LLMs, incorporate domain-specific knowledge [li2023BIRD, hong2024knowledgetosql], and develop more efficient fine-tuning strategies [li2024codes].
 As the field continues to evolve, we anticipate the development of more advanced and superior LLM-based implementations that will elevate the performance and generalization of text-to-SQL to new heights.
@@
 Original Dataset Release Time #Example #DB #Table/DB #Row/DB Language Characteristics
-BIRD-CRITIC [1] Feb-2025 600 - - - EN Knowledge-augmented, Long-context
-Spider 2.0 [42] Nov-2024 632 213 - - EN Knowledge-augmented, Long-context
-BULL [45] Jan-2024 4,966 3 26 - EN Specialized-domain
-BIRD [1] May-2023 12,751 95 7.3 549K EN Cross-domain, Knowledge-augmented
-KaggleDBQA [33] Jun-2021 272 8 2.3 280K EN Cross-domain
-DuSQL [31] Nov-2020 23,797 200 4.1 - EN Cross-domain, Cross-lingual
-SQUALL [41] Oct-2020 11,468 1,679 1 - EN Knowledge-augmented
-CoSQL [32] Sep-2019 15,598 200 - - EN Cross-domain, Context-dependent
-Spider [11] Sep-2018 10,181 200 5.1 2K EN Cross-domain
-WikiSQL [12] Aug-2017 80,654 26,521 1 17 EN Cross-domain
+BIRD-CRITIC [li2023BIRD] Feb-2025 600 - - - EN Knowledge-augmented, Long-context
+Spider 2.0 [lei2025spider2] Nov-2024 632 213 - - EN Knowledge-augmented, Long-context
+BULL [zhang2024finsql] Jan-2024 4,966 3 26 - EN Specialized-domain
+BIRD [li2023BIRD] May-2023 12,751 95 7.3 549K EN Cross-domain, Knowledge-augmented
+KaggleDBQA [lee2021KaggleDBQA] Jun-2021 272 8 2.3 280K EN Cross-domain
+DuSQL [wang2020DuSQL] Nov-2020 23,797 200 4.1 - EN Cross-domain, Cross-lingual
+SQUALL [shi2020SQUALL] Oct-2020 11,468 1,679 1 - EN Knowledge-augmented
+CoSQL [yu2019CoSQL] Sep-2019 15,598 200 - - EN Cross-domain, Context-dependent
+Spider [yu2018Spider] Sep-2018 10,181 200 5.1 2K EN Cross-domain
+WikiSQL [zhong2017WikiSQL] Aug-2017 80,654 26,521 1 17 EN Cross-domain
 Post-annotated Dataset Release Time Source Dataset Special Setting Language Characteristics
-Dr. Spider [44] Jan-2023 Spider Perturbations in DB, query and SQL EN Robustness
-ADVETA [34] Dec-2022 Spider, etc. Adversarial table perturbation EN Robustness
-Spider-SS&CG [35] May-2022 Spider Splitting example into sub-examples EN Context-dependent
-Spider-DK [36] Sep-2021 Spider Adding domain knowledge EN Knowledge-augmented
-Spider-SYN [37] Jun-2021 Spider Manual synonym replacement EN Robustness
-Spider-Vietnamese [43] Nov-2020 Spider Vietnamese version of Spider VI Cross-lingual
-Spider-Realistic [38] Oct-2020 Spider Removing column names in question EN Robustness
-CSpider [39] Sep-2019 Spider Chinese version of Spider ZH Cross-lingual
-SParC [40] Jun-2019 Spider Annotate conversational contents EN Context-dependent
+Dr. Spider [chang2023dr] Jan-2023 Spider Perturbations in DB, query and SQL EN Robustness
+ADVETA [pi2022ADVETA] Dec-2022 Spider, etc. Adversarial table perturbation EN Robustness
+Spider-SS&CG [gan2022Spider-CG] May-2022 Spider Splitting example into sub-examples EN Context-dependent
+Spider-DK [gan2021Spider-DK] Sep-2021 Spider Adding domain knowledge EN Knowledge-augmented
+Spider-SYN [gan2021Spider-SYN] Jun-2021 Spider Manual synonym replacement EN Robustness
+Spider-Vietnamese [tuan2020pilot] Nov-2020 Spider Vietnamese version of Spider VI Cross-lingual
+Spider-Realistic [deng2021Spider-Realistic] Oct-2020 Spider Removing column names in question EN Robustness
+CSpider [min2019CSpider] Sep-2019 Spider Chinese version of Spider ZH Cross-lingual
+SParC [yu2019SParC] Jun-2019 Spider Annotate conversational contents EN Context-dependent
 III-A Datasets
@@
 This refers to datasets where the background information of different databases comes from various domains.
-Since real-world text-to-SQL applications often involve databases from multiple domains, most original text-to-SQL datasets [12, 11, 32, 31, 33, 1] and post-annotated datasets [40, 39, 36, 37, 38, 34, 35] are in the cross-domain setting to fit well with the requirements of cross-domain applications.
+Since real-world text-to-SQL applications often involve databases from multiple domains, most original text-to-SQL datasets [zhong2017WikiSQL, yu2018Spider, yu2019CoSQL, wang2020DuSQL, lee2021KaggleDBQA, li2023BIRD] and post-annotated datasets [yu2019SParC, min2019CSpider, gan2021Spider-DK, gan2021Spider-SYN, deng2021Spider-Realistic, pi2022ADVETA, gan2022Spider-CG] are in the cross-domain setting to fit well with the requirements of cross-domain applications.
 III-A2 Knowledge-augmented Dataset
 Interest in incorporating domain-specific knowledge into text-to-SQL tasks has increased significantly in recent years.
-BIRD [1] employs human database experts to annotate each text-to-SQL sample with external knowledge, categorized into numeric reasoning knowledge, domain knowledge, synonym knowledge, and value illustration.
-Similarly, Spider-DK [36] defines and adds five types of domain knowledge for a human-curated version of the Spider dataset [11]: SELECT columns mentioned by omission, simple inference required, synonym substitution in cell value words, one non-cell value word generates a condition, and easy to conflict with other domains.
+BIRD [li2023BIRD] employs human database experts to annotate each text-to-SQL sample with external knowledge, categorized into numeric reasoning knowledge, domain knowledge, synonym knowledge, and value illustration.
+Similarly, Spider-DK [gan2021Spider-DK] defines and adds five types of domain knowledge for a human-curated version of the Spider dataset [yu2018Spider]: SELECT columns mentioned by omission, simple inference required, synonym substitution in cell value words, one non-cell value word generates a condition, and easy to conflict with other domains.
 Both studies found that human-annotated knowledge significantly improves SQL generation performance for samples requiring external domain knowledge.
-Additionally, SQUALL [41] manually annotates alignments between the words in NL questions and the entities in SQL, providing finer-grained supervision than other datasets.
+Additionally, SQUALL [shi2020SQUALL] manually annotates alignments between the words in NL questions and the entities in SQL, providing finer-grained supervision than other datasets.
 III-A3 Context-dependent Dataset
-SParC [40] and CoSQL [32] explore context-dependent SQL generation by constructing a conversational database querying system.
+SParC [yu2019SParC] and CoSQL [yu2019CoSQL] explore context-dependent SQL generation by constructing a conversational database querying system.
 Unlike traditional text-to-SQL datasets that only have a single question-SQL pair for one example, SParC decomposes the question-SQL examples in the Spider dataset into multiple sub-question-SQL pairs to construct a simulated and meaningful interaction, including interrelated sub-questions that aid SQL generation, and unrelated sub-questions that enhance data diversity.
 CoSQL, in comparison, involves conversational interactions in natural language, simulating real-world scenarios to increase complexity and diversity.
-Additionally, Spider-SS&CG [35] splits the NL question in the Spider dataset [11] into multiple sub-questions and sub-SQLs, demonstrating that training on these sub-examples can improve a text-to-SQL system’s generalization ability on out-of-distribution samples.
+Additionally, Spider-SS&CG [gan2022Spider-CG] splits the NL question in the Spider dataset [yu2018Spider] into multiple sub-questions and sub-SQLs, demonstrating that training on these sub-examples can improve a text-to-SQL system’s generalization ability on out-of-distribution samples.
 III-A4 Robustness Dataset
 Evaluating the accuracy of text-to-SQL systems with polluted or perturbed database contents (e.g., schema and tables) is crucial for assessing robustness.
-Spider-Realistic [38] removes explicit schema-related words from the NL questions, while Spider-SYN [37] replaces them with manually selected synonyms.
-ADVETA [34] introduces adversarial table perturbation (ATP), which perturbs tables by replacing original column names with misleading alternatives and inserting new columns with high semantic associations but low semantic equivalency.
+Spider-Realistic [deng2021Spider-Realistic] removes explicit schema-related words from the NL questions, while Spider-SYN [gan2021Spider-SYN] replaces them with manually selected synonyms.
+ADVETA [pi2022ADVETA] introduces adversarial table perturbation (ATP), which perturbs tables by replacing original column names with misleading alternatives and inserting new columns with high semantic associations but low semantic equivalency.
 Based on the generation ability of large-scale pre-trained language models, Dr.
-Spider [44] provided 17 different perturbation strategies on databases, natural language queries and SQL statements to measure the robustness of text-to-SQL systems.
+Spider [chang2023dr] provided 17 different perturbation strategies on databases, natural language queries and SQL statements to measure the robustness of text-to-SQL systems.
 These perturbations lead to significant drops in accuracy, as a text-to-SQL system with low robustness may be misled by incorrect matches between tokens in NL questions and database entities.
@@
 The SQL keywords, function names, table and column names are typically written in English, posing challenges for applications in other languages.
-CSpider [39] addresses this by translating the Spider dataset into Chinese, identifying new challenges in word segmentation and cross-lingual matching between Chinese questions and English database contents.
-DuSQL [31] introduces a practical text-to-SQL dataset with Chinese questions and database contents provided in both English and Chinese.
-Spider-Vietnamese [43] translates the Spider dataset into Vietnamese, empirically revealing several strategies to enhance text-to-SQL methods within the Vietnamese context.
+CSpider [min2019CSpider] addresses this by translating the Spider dataset into Chinese, identifying new challenges in word segmentation and cross-lingual matching between Chinese questions and English database contents.
+DuSQL [wang2020DuSQL] introduces a practical text-to-SQL dataset with Chinese questions and database contents provided in both English and Chinese.
+Spider-Vietnamese [tuan2020pilot] translates the Spider dataset into Vietnamese, empirically revealing several strategies to enhance text-to-SQL methods within the Vietnamese context.
 III-A6 Long-context dataset
 In real-world database applications, numerous challenging scenarios demand complex SQL queries with significant token lengths, often exceeding 100 tokens and involving multiple intricate operations.
-Spider 2.0 [42] exemplifies this complexity by presenting enterprise-level text-to-SQL problems that require advanced reasoning across various SQL queries and dialects.
+Spider 2.0 [lei2025spider2] exemplifies this complexity by presenting enterprise-level text-to-SQL problems that require advanced reasoning across various SQL queries and dialects.
 Distinct from earlier benchmarks, Spider 2.0 contains numerous real-world tasks sourced from diverse production databases such as BigQuery and Snowflake, featuring large-scale, industrial schemas with hundreds or even thousands of columns, nested structures, and multiple schemas per instance.
@@
 As a result, Spider 2.0 highlighted the substantial gap between present-day LLMs’ capabilities and the demands of authentic, long-context database scenarios.
-In parallel, BIRD-CRITIC [^1] https://bird-critic.github.io/ [1] introduces a diagnostic dimension, comprising 600 tasks designed to assess the efficacy of LLMs in resolving real-world user issues across diverse SQL dialects.
+In parallel, BIRD-CRITIC [^1] https://bird-critic.github.io/ [li2023BIRD] introduces a diagnostic dimension, comprising 600 tasks designed to assess the efficacy of LLMs in resolving real-world user issues across diverse SQL dialects.
 Together, these datasets push the boundaries of text-to-SQL research, emphasizing the challenges of complex, long-context SQL environments.
@@
 Despite advancements in text-to-SQL research, there is still a shortage of specialized datasets tailored for domain-specific analysis.
-To address this gap, BULL [45] introduces a dedicated text-to-SQL benchmark for the financial sector.
+To address this gap, BULL [zhang2024finsql] introduces a dedicated text-to-SQL benchmark for the financial sector.
 The dataset encompasses databases related to funds, stocks, and macroeconomic data, offering a comprehensive platform to evaluate and enhance text-to-SQL methods in financial applications, effectively tackling the financial domain’s unique text-to-SQL challenges.
@@
 SQL content matching metrics focus on comparing the predicted SQL with the ground truth based on their structural and syntactic similarities.
-  • Component Matching (CM) [11] evaluates the performance of text-to-SQL system by measuring the exact match between predicted and ground truth SQL components—SELECT, WHERE, GROUP BY, ORDER BY, and KEYWORDS—using the F1 score.
+  • Component Matching (CM) [yu2018Spider] evaluates the performance of text-to-SQL system by measuring the exact match between predicted and ground truth SQL components—SELECT, WHERE, GROUP BY, ORDER BY, and KEYWORDS—using the F1 score.
 Each component is decomposed into sets of sub-components and compared for an exact match, accounting for SQL components without order constraints.
-  • Exact Matching (EM) [11] measures the percentage of examples whose predicted SQL query is identical to the ground truth SQL query.
+  • Exact Matching (EM) [yu2018Spider] measures the percentage of examples whose predicted SQL query is identical to the ground truth SQL query.
 A predicted SQL is considered correct only if all its components, as described in CM, match exactly with those of the ground truth query.
@@
 Execution result metrics assess the correctness of the generated SQL query by comparing the results obtained from executing the query on the target database with the expected results.
-  • Execution Accuracy (EX) [11] measures the correctness of a predicted SQL query by executing it in the corresponding database and comparing the executed results with the results obtained by the ground truth query.
-  • Valid Efficiency Score (VES) [1] is defined to measure the efficiency of valid SQL queries.
+  • Execution Accuracy (EX) [yu2018Spider] measures the correctness of a predicted SQL query by executing it in the corresponding database and comparing the executed results with the results obtained by the ground truth query.
+  • Valid Efficiency Score (VES) [li2023BIRD] is defined to measure the efficiency of valid SQL queries.
 A valid SQL query is a predicted SQL whose executed results exactly match the ground truth results.
@@
     Then, $\textbf{R}(Y_{n},\hat{Y}_{n})=\sqrt{E(Y_{n})/E(\hat{Y}_{n})}$ denotes the relative execution efficiency of the predicted SQL in comparison to the ground-truth, where $E(\cdot)$ is the execution time of each SQL in the database.
-BIRD benchmark [1] ensures the stability of this metric by computing the average of $\textbf{R}(Y_{n},\hat{Y}_{n})$ over 100 runs for each example.
+BIRD benchmark [li2023BIRD] ensures the stability of this metric by computing the average of $\textbf{R}(Y_{n},\hat{Y}_{n})$ over 100 runs for each example.
 Most of the recent LLM-based text-to-SQL studies focus on BIRD and Spider-related datasets and execution-based metrics. The corresponding details can be found at Tab. III.
 IV Methods
-The implementation of the most current LLM-based text-to-SQL methods relies on in-context learning (ICL) [109, 110, 111] and fine-tuning (FT) [112, 113] paradigms due to the release of powerful proprietary and well-architected open-source LLMs in large quantities [107, 114, 115, 116, 117, 118].
+The implementation of the most current LLM-based text-to-SQL methods relies on in-context learning (ICL) [sahoo2024systematic, wang2023prompt, chen2023unleashing] and fine-tuning (FT) [wei2021finetuned, zheng2024llamafactory] paradigms due to the release of powerful proprietary and well-architected open-source LLMs in large quantities [achiam2023gpt4, xue2024dbgpt, wang2023dbcopilot, touvron2023llama, touvron2023llama2, bai2023qwen].
 In this section, we will discuss these paradigms with detailed categorization accordingly.
 IV-A In-context Learning
-Through extensive and widely recognized research, ICL (also called prompt engineering) has been proven to play a decisive role in the performance of LLMs across various tasks [119, 26].
-Consequently, it also impacts SQL generation across different prompt styles [8, 120].
+Through extensive and widely recognized research, ICL (also called prompt engineering) has been proven to play a decisive role in the performance of LLMs across various tasks [radford2019unsupervised, yang2024harnessing].
+Consequently, it also impacts SQL generation across different prompt styles [gao2023dailsql, zhang2024benchmarking].
 The ICL implementation of the LLM-based text-to-SQL process to generate a predicted SQL query $\hat{Y}$, can be formulated as:
@@
 Here, $\mathcal{C}=\{c_{1},c_{2},$...\} and $\mathcal{T}=\{t_{1},t_{2},$...\} represent the collection of various columns $c$ and tables $t$, respectively.
-$\mathcal{K}$ is the potentially supplementary knowledge, such as foreign key relationships [68], schema linking [28, 29], and external knowledge [1, 70].
+$\mathcal{K}$ is the potentially supplementary knowledge, such as foreign key relationships [zhang2023actsql], schema linking [dong2023c3, yuan2025knapsack], and external knowledge [li2023BIRD, hong2024knowledgetosql].
 $\pi(\cdot\mid\theta)$ is a LLM $\pi$ with parameter $\theta$.
@@
 We further specify the prompt engineering strategies in the ICL paradigm as zero-shot and few-shot prompting.
-Trained through massive data, LLMs have a strong overall proficiency in different downstream tasks with zero-shot and few-shot prompting [121, 112, 122], which is widely recognized and utilized in real-world applications.
+Trained through massive data, LLMs have a strong overall proficiency in different downstream tasks with zero-shot and few-shot prompting [reynolds2021prompt, wei2021finetuned, ye2022unreliability], which is widely recognized and utilized in real-world applications.
 Formally, the overall input $P_{0}$ of zero-shot prompting for text-to-SQL can be obtained by concatenating $I$, $\mathcal{S}$, and $Q$ as introduced above:
@@
 The OpenAI demonstration [^2] The prompt format follows the official document provided by the OpenAI platform: https://platform.openai.com/examples/default-sql-translate.
-provides an indicative example and regulation for the prompting process in LLM-based text-to-SQL [28].
+provides an indicative example and regulation for the prompting process in LLM-based text-to-SQL [dong2023c3].
 The prompt designs and formats utilized by a substantial portion of the ICL methods discussed in this survey are developed based on the OpenAI regulation.
@@
 III.
-*Note that SimpleDDL is a prompting style that provides only table and column names [120], rather than a tailored method.
+*Note that SimpleDDL is a prompting style that provides only table and column names [zhang2024benchmarking], rather than a tailored method.
 Methods Adopted by Applied LLMs
-$\mathbf{C}_{0}$-Vanilla Prompting SimpleDDL [120]* SQLCoder
-$\mathbf{C}_{1}$-Decomposition DIN-SQL [7] GPT-4
-$\mathbf{C}_{2}$-Prompt Optimization DAIL-SQL [8] GPT-4
-$\mathbf{C}_{3}$-Reasoning Enhancement ACT-SQL [68] GPT-4
-$\mathbf{C}_{4}$-Execution Refinement LEVER [79] Codex
+$\mathbf{C}_{0}$-Vanilla Prompting SimpleDDL [zhang2024benchmarking]* SQLCoder
+$\mathbf{C}_{1}$-Decomposition DIN-SQL [pourreza2023dinsql] GPT-4
+$\mathbf{C}_{2}$-Prompt Optimization DAIL-SQL [gao2023dailsql] GPT-4
+$\mathbf{C}_{3}$-Reasoning Enhancement ACT-SQL [zhang2023actsql] GPT-4
+$\mathbf{C}_{4}$-Execution Refinement LEVER [ni2023lever] Codex
 TABLE III: Well-designed methods in the ICL paradigm for LLM-based text-to-SQL are ordered by release time.
@@
 Methods that fall into multiple categories will be introduced accordingly.
-The dataset column consists of six text-to-SQL datasets, which can be clarified by names: Spider [11], BIRD [1], Spider-DK [36] Spider-SYN [37], Spider-Realistic [38], and Spider 2.0 [42]
+The dataset column consists of six text-to-SQL datasets, which can be clarified by names: Spider [yu2018Spider], BIRD [li2023BIRD], Spider-DK [gan2021Spider-DK] Spider-SYN [gan2021Spider-SYN], Spider-Realistic [deng2021Spider-Realistic], and Spider 2.0 [lei2025spider2]
 .
@@
 *There are multiple applied LLMs in the corresponding method; the model with representative performance is presented.
-CoT method is introduced in multiple venues: NeurIPS’23 [1], EMNLP’23 [47], VLDB’24 [8], arXiv’24 [30].
-Methods Applied LLMs Dataset Metrics $\mathbf{C}_{1}$ $\mathbf{C}_{2}$ $\mathbf{C}_{3}$ $\mathbf{C}_{4}$ Release Time Publication Venue MBR-Exec [78] Codex [11] EX Apr-2022 EMNLP’22 Coder-Reviewer [46] Codex [11] EX Nov-2022 ICML’23 LEVER [79] Codex [11] EX Feb-2023 ICML’23 Self-Debugging [80] StarCoder* [11] EX Apr-2023 ICLR’24 DESEM+P [63] ChatGPT [11, 37] EX Apr-2023 PRICAI’23 DIN-SQL [7] GPT-4* [1, 11, 42] EX, EM, VES Apr-2023 NeurIPS’23 CoT [47, 1, 8, 30] GPT-4 [11, 38, 1] EX, VES May-2023 Multiple Venues StructGPT [64] ChatGPT* [11, 37, 38] EX May-2023 EMNLP’23 SD+SA+Voting [65] ChatGPT* [11, 37, 38] EX May-2023 EMNLP’23 Findings QDecomp [47] Codex [11, 38] EX May-2023 EMNLP’23 Least-to-Most [47] Codex [11] EX May-2023 EMNLP’23 SQL-PaLM [74] PaLM-2 [11] EX May-2023 TMLR’24 RAG+SP&DRC [66] ChatGPT [11] EX Jul-2023 ICONIP’23 C3 [28] ChatGPT [11] EX Jul-2023 arXiv’23 DAIL-SQL [8] GPT-4* [11, 1, 38, 42] EX, EM, VES Aug-2023 VLDB’24 ODIS [67] Codex* [11] EX Oct-2023 EMNLP’23 Findings ACT-SQL [68] GPT-4* [11, 37] EX, EM Oct-2023 EMNLP’23 Findings MAC-SQL [48] GPT-4* [11, 1] EX, EM, VES Dec-2023 COLING’25 DEA-SQL [49] GPT-4 [11] EX Feb-2024 ACL’24 Findings FUSED [69] ChatGPT* [11] EX Feb-2024 EMNLP’24 Findings DELLM [70] GPT-4* [11, 1] EX, VES Feb-2024 ACL’24 Findings SGU-SQL [30] GPT-4* [11, 1] EX, EM Feb-2024 ICML’25 PoT [75] GPT-4* [11, 1] EX Feb-2024 arXiv’24 SQL-CRAFT [75] GPT-4* [11, 1] EX Feb-2024 arXiv’24 FUXI [76] GPT-4* [1] EX Feb-2024 EMNLP’24 MetaSQL [50] GPT-4* [11] EX, EM Feb-2024 ICDE’24 $\textit{{R}}^{3}$ [51] GPT-4* [11, 1] EX Feb-2024 ACL’25 Workshop PET-SQL [52] GPT-4 [11] EX Mar-2024 arXiv’24 PURPLE [53] GPT-4* [11, 37, 38] EX, EM Mar-2024 ICDE’24 Dubo-SQL [71] ChatGPT* [1] EX Apr-2024 arXiv’24 TA-SQL [54] GPT-4* [11, 1, 36, 38] EX Apr-2024 ACL’24 Findings MCS-SQL [55] GPT-4 [11, 1] EX, VES May-2024 COLING’25 CHESS [56] GPT-4o* [11, 1, 42] EX, VES May-2024 arXiv’24 SuperSQL [57] GPT-4* [11, 1] EX, EM, VES Jun-2024 VLDB’24 Distillery [58] GPT-4o-Mini* [1] EX Aug-2024 NeurIP’24 Workshop E-SQL [81] GPT-4o* [11, 1] EX, VES Sep-2024 arXiv’24 CHASE-SQL [77] Gemini-1.5* [11, 1] EX, VES Oct-2024 ICLR’25 RSL-SQL [59] DeepSeek-V3* [11, 1] EX, VES Nov-2024 arXiv’24 Spider-Agent [42] o1-preview* [42] EX Nov-2024 ICLR’25 ICRL [72] ChatGPT [11] EX Jan-2025 COLING’25 Gen-SQL [60] LLaMA-3* [11, 1] EX, EM, VES Jan-2025 COLING’25 SAFE-SQL [73] GPT-4o-Mini* [11] EX, EM Feb-2025 arXiv’25 ReFoRCE [61] o1-preview* [42] EX Feb-2025 ICLR’25 Workshop LinkAlign [62] DeepSeek-R1* [1, 42] EX, EM Mar-2025 arXiv’25
+CoT method is introduced in multiple venues: NeurIPS’23 [li2023BIRD], EMNLP’23 [tai2023exploring], VLDB’24 [gao2023dailsql], arXiv’24 [zhang2024sgusql].
+Methods Applied LLMs Dataset Metrics $\mathbf{C}_{1}$ $\mathbf{C}_{2}$ $\mathbf{C}_{3}$ $\mathbf{C}_{4}$ Release Time Publication Venue MBR-Exec [shi2022natural] Codex [yu2018Spider] EX Apr-2022 EMNLP’22 Coder-Reviewer [zhang2022coder] Codex [yu2018Spider] EX Nov-2022 ICML’23 LEVER [ni2023lever] Codex [yu2018Spider] EX Feb-2023 ICML’23 Self-Debugging [chen2024selfdebugging] StarCoder* [yu2018Spider] EX Apr-2023 ICLR’24 DESEM+P [guo2023desem] ChatGPT [yu2018Spider, gan2021Spider-SYN] EX Apr-2023 PRICAI’23 DIN-SQL [pourreza2023dinsql] GPT-4* [li2023BIRD, yu2018Spider, lei2025spider2] EX, EM, VES Apr-2023 NeurIPS’23 CoT [tai2023exploring, li2023BIRD, gao2023dailsql, zhang2024sgusql] GPT-4 [yu2018Spider, deng2021Spider-Realistic, li2023BIRD] EX, VES May-2023 Multiple Venues StructGPT [jiang2023structgpt] ChatGPT* [yu2018Spider, gan2021Spider-SYN, deng2021Spider-Realistic] EX May-2023 EMNLP’23 SD+SA+Voting [nan2023enhancing] ChatGPT* [yu2018Spider, gan2021Spider-SYN, deng2021Spider-Realistic] EX May-2023 EMNLP’23 Findings QDecomp [tai2023exploring] Codex [yu2018Spider, deng2021Spider-Realistic] EX May-2023 EMNLP’23 Least-to-Most [tai2023exploring] Codex [yu2018Spider] EX May-2023 EMNLP’23 SQL-PaLM [sun2023sqlpalm] PaLM-2 [yu2018Spider] EX May-2023 TMLR’24 RAG+SP&DRC [guo2023ragsql] ChatGPT [yu2018Spider] EX Jul-2023 ICONIP’23 C3 [dong2023c3] ChatGPT [yu2018Spider] EX Jul-2023 arXiv’23 DAIL-SQL [gao2023dailsql] GPT-4* [yu2018Spider, li2023BIRD, deng2021Spider-Realistic, lei2025spider2] EX, EM, VES Aug-2023 VLDB’24 ODIS [chang2023selective] Codex* [yu2018Spider] EX Oct-2023 EMNLP’23 Findings ACT-SQL [zhang2023actsql] GPT-4* [yu2018Spider, gan2021Spider-SYN] EX, EM Oct-2023 EMNLP’23 Findings MAC-SQL [wang2024macsql] GPT-4* [yu2018Spider, li2023BIRD] EX, EM, VES Dec-2023 COLING’25 DEA-SQL [xie2024deasql] GPT-4 [yu2018Spider] EX Feb-2024 ACL’24 Findings FUSED [wang2024fused] ChatGPT* [yu2018Spider] EX Feb-2024 EMNLP’24 Findings DELLM [hong2024knowledgetosql] GPT-4* [yu2018Spider, li2023BIRD] EX, VES Feb-2024 ACL’24 Findings SGU-SQL [zhang2024sgusql] GPT-4* [yu2018Spider, li2023BIRD] EX, EM Feb-2024 ICML’25 PoT [xia2024sqlcraft] GPT-4* [yu2018Spider, li2023BIRD] EX Feb-2024 arXiv’24 SQL-CRAFT [xia2024sqlcraft] GPT-4* [yu2018Spider, li2023BIRD] EX Feb-2024 arXiv’24 FUXI [gu2024fuxi] GPT-4* [li2023BIRD] EX Feb-2024 EMNLP’24 MetaSQL [fan2024metasql] GPT-4* [yu2018Spider] EX, EM Feb-2024 ICDE’24 $\textit{{R}}^{3}$ [xia2024r3] GPT-4* [yu2018Spider, li2023BIRD] EX Feb-2024 ACL’25 Workshop PET-SQL [li2024petsql] GPT-4 [yu2018Spider] EX Mar-2024 arXiv’24 PURPLE [ren2024purple] GPT-4* [yu2018Spider, gan2021Spider-SYN, deng2021Spider-Realistic] EX, EM Mar-2024 ICDE’24 Dubo-SQL [thorpe2024dubo] ChatGPT* [li2023BIRD] EX Apr-2024 arXiv’24 TA-SQL [qu2024generation] GPT-4* [yu2018Spider, li2023BIRD, gan2021Spider-DK, deng2021Spider-Realistic] EX Apr-2024 ACL’24 Findings MCS-SQL [lee2025mcs] GPT-4 [yu2018Spider, li2023BIRD] EX, VES May-2024 COLING’25 CHESS [talaei2024chess] GPT-4o* [yu2018Spider, li2023BIRD, lei2025spider2] EX, VES May-2024 arXiv’24 SuperSQL [li2024dawn] GPT-4* [yu2018Spider, li2023BIRD] EX, EM, VES Jun-2024 VLDB’24 Distillery [maamari2024the] GPT-4o-Mini* [li2023BIRD] EX Aug-2024 NeurIP’24 Workshop E-SQL [caferouglu2024esql] GPT-4o* [yu2018Spider, li2023BIRD] EX, VES Sep-2024 arXiv’24 CHASE-SQL [pourreza2025chasesql] Gemini-1.5* [yu2018Spider, li2023BIRD] EX, VES Oct-2024 ICLR’25 RSL-SQL [cao2024rsl] DeepSeek-V3* [yu2018Spider, li2023BIRD] EX, VES Nov-2024 arXiv’24 Spider-Agent [lei2025spider2] o1-preview* [lei2025spider2] EX Nov-2024 ICLR’25 ICRL [toteja2025context] ChatGPT [yu2018Spider] EX Jan-2025 COLING’25 Gen-SQL [shi2025gen] LLaMA-3* [yu2018Spider, li2023BIRD] EX, EM, VES Jan-2025 COLING’25 SAFE-SQL [lee2025safe] GPT-4o-Mini* [yu2018Spider] EX, EM Feb-2025 arXiv’25 ReFoRCE [deng2025reforce] o1-preview* [lei2025spider2] EX Feb-2025 ICLR’25 Workshop LinkAlign [wang2025linkalign] DeepSeek-R1* [li2023BIRD, lei2025spider2] EX, EM Mar-2025 arXiv’25
 1.
-Zero-shot studies: The zero-shot-based ICL strategies primarily focus on studying the influence of prompt representation styles and the baseline performance of various LLMs for text-to-SQL [6, 25, 120].
-As an empirical evaluation, [6] evaluates the baseline text-to-SQL capabilities of various early-developed LLMs [123, 106, 124] and their performance across different prompting styles.
+Zero-shot studies: The zero-shot-based ICL strategies primarily focus on studying the influence of prompt representation styles and the baseline performance of various LLMs for text-to-SQL [rajkumar2022evaluating, liu2023comprehensive, zhang2024benchmarking].
+As an empirical evaluation, [rajkumar2022evaluating] evaluates the baseline text-to-SQL capabilities of various early-developed LLMs [raffel2020exploring, brown2020gpt3, chen2021evaluating] and their performance across different prompting styles.
 The experimental results indicate that prompt design is critical for performance.
-Through error analysis, this study [6] further suggests that including more database content can harm the overall accuracy of text-to-SQL.
-Since the emergence of ChatGPT [106] with its impressive capabilities in conversational scenarios and code generation [125], [25] evaluates its performance on Spider [11] and its variant [37, 36, 38] datasets.
+Through error analysis, this study [rajkumar2022evaluating] further suggests that including more database content can harm the overall accuracy of text-to-SQL.
+Since the emergence of ChatGPT [brown2020gpt3] with its impressive capabilities in conversational scenarios and code generation [ray2023chatgpt], [liu2023comprehensive] evaluates its performance on Spider [yu2018Spider] and its variant [gan2021Spider-SYN, gan2021Spider-DK, deng2021Spider-Realistic] datasets.
 Under zero-shot settings, the results demonstrate that ChatGPT has promising text-to-SQL performance compared to state-of-the-art PLM-based systems.
-Aside from the proprietary models discussed above, similar evaluations indicate that open-source LLMs can also perform zero-shot text-to-SQL tasks [114, 83, 120], particularly for the code-specific models [80, 120].
+Aside from the proprietary models discussed above, similar evaluations indicate that open-source LLMs can also perform zero-shot text-to-SQL tasks [xue2024dbgpt, xu2024symbolllm, zhang2024benchmarking], particularly for the code-specific models [chen2024selfdebugging, zhang2024benchmarking].
 Primary keys and foreign keys convey continuous knowledge across different tables, making them an effective prompting style in zero-shot scenarios.
-[68] studied their impact by incorporating these keys into various prompt representation styles with different database content to analyze performance outcomes.
-A benchmark evaluation [8] also assessed the influence of foreign keys using five distinct prompt representation styles.
+[zhang2023actsql] studied their impact by incorporating these keys into various prompt representation styles with different database content to analyze performance outcomes.
+A benchmark evaluation [gao2023dailsql] also assessed the influence of foreign keys using five distinct prompt representation styles.
 Each style can be viewed as a permutation and combination of instruction, rule implication, and foreign key elements.
 In addition to using foreign keys, this study explores zero-shot prompting combined with the rule implication “no explanation” to produce concise outputs.
-With the help of annotated external knowledge from human experts, the evaluation in the BIRD benchmark [1] significantly improves by incorporating the provided annotated oracle knowledge.
-In terms of zero-shot prompting optimization, [120] highlights the challenge of designing effective prompt templates for LLMs.
+With the help of annotated external knowledge from human experts, the evaluation in the BIRD benchmark [li2023BIRD] significantly improves by incorporating the provided annotated oracle knowledge.
+In terms of zero-shot prompting optimization, [zhang2024benchmarking] highlights the challenge of designing effective prompt templates for LLMs.
 Previous simple prompt constructions often lack structural uniformity, complicating the identification of specific elements in a prompt template that influence LLM performance.
 This challenge is addressed by investigating a more unified set of prompt templates that include various prefixes, infixes, and postfixes.
-To ensure fair comparability, [108] explores effective prompt construction for LLMs by studying different styles of prompt representation and highlights their findings on the crucial role of table structure, content, prompt length, and optimal representations in effective prompting.
+To ensure fair comparability, [chang2023how] explores effective prompt construction for LLMs by studying different styles of prompt representation and highlights their findings on the crucial role of table structure, content, prompt length, and optimal representations in effective prompting.
 2.
 Few-shot studies: The study of few-shot prompting strategies focuses on the number and selection of few-shot instances.
-Empirical studies have evaluated few-shot prompting for text-to-SQL across multiple datasets using various LLMs [7, 30], demonstrating solid improvements compared to zero-shot prompting.
-[1] provides a detailed 1-shot example to trigger the text-to-SQL model for generating accurate SQL.
-[75] explores the effect of the number of few-shot examples.
-[65] focuses on sampling strategies by examining the similarity and diversity between different demonstrations, setting random sampling as the baseline, and evaluating various strategies and their combinations for comparison.
-Furthermore, beyond similarity-based selection, [8] evaluates masked question similarity selection and the upper limit of similarity approaches with various numbers of few-shot examples.
-A study on difficulty-level sample selection [47] compared the performance of few-shot Codex [124], using random selection and difficulty-based selection for few-shot instances on difficulty-categorized datasets [11, 38].
+Empirical studies have evaluated few-shot prompting for text-to-SQL across multiple datasets using various LLMs [pourreza2023dinsql, zhang2024sgusql], demonstrating solid improvements compared to zero-shot prompting.
+[li2023BIRD] provides a detailed 1-shot example to trigger the text-to-SQL model for generating accurate SQL.
+[xia2024sqlcraft] explores the effect of the number of few-shot examples.
+[nan2023enhancing] focuses on sampling strategies by examining the similarity and diversity between different demonstrations, setting random sampling as the baseline, and evaluating various strategies and their combinations for comparison.
+Furthermore, beyond similarity-based selection, [gao2023dailsql] evaluates masked question similarity selection and the upper limit of similarity approaches with various numbers of few-shot examples.
+A study on difficulty-level sample selection [tai2023exploring] compared the performance of few-shot Codex [chen2021evaluating], using random selection and difficulty-based selection for few-shot instances on difficulty-categorized datasets [yu2018Spider, deng2021Spider-Realistic].
 Three difficulty-based selection strategies are devised based on samples selected by difficulty level.
-[68] utilizes a hybrid strategy for selecting samples, combining static examples and similarity-based dynamic examples for few-shot prompting.
+[zhang2023actsql] utilizes a hybrid strategy for selecting samples, combining static examples and similarity-based dynamic examples for few-shot prompting.
 In their settings, they also evaluate the impact of different input schema styles and various numbers of static and dynamic exemplars.
-The impact of cross-domain examples is also being studied [67].
+The impact of cross-domain examples is also being studied [chang2023selective].
 When incorporating in- and out-of-domain examples in different numbers, the in-domain demonstrations outperform zero-shot and out-of-domain examples, with performance improving as the number of examples increases.
-To explore the detailed construction of prompts, [74] compares concise and verbose prompt design strategies.
+To explore the detailed construction of prompts, [sun2023sqlpalm] compares concise and verbose prompt design strategies.
 The former style separates the schema, column names, primary keys, and foreign keys, while the latter organizes them as natural language descriptions.
@@
 Multi-module collaboration: This strategy breaks the task into specialized stages (e.g., schema linking, SQL generation, and SQL refinement) handled by distinct modules or agents.
-C3 [28] employs a three-module pipeline involving schema linking via clear prompting, calibration bias prompting for SQL guidance, and consistency-based voting for final selection.
-Similarly, MAC-SQL [48] and CHESS [56] use multi-agent frameworks where dedicated agents handle schema selection, decomposition, and refinement.
-Workflow-centric methods, like Distillery [58] and DEA-SQL [49], decompose the task into retrieval, generation, and correction stages.
+C3 [dong2023c3] employs a three-module pipeline involving schema linking via clear prompting, calibration bias prompting for SQL guidance, and consistency-based voting for final selection.
+Similarly, MAC-SQL [wang2024macsql] and CHESS [talaei2024chess] use multi-agent frameworks where dedicated agents handle schema selection, decomposition, and refinement.
+Workflow-centric methods, like Distillery [maamari2024the] and DEA-SQL [xie2024deasql], decompose the task into retrieval, generation, and correction stages.
 Iterative refinement and structural decomposition are also prevalent.
-DIN-SQL [7] integrates a four-stage pipeline, including schema linking, classification, SQL generation, and self-correction.
-Coder-Reviewer [46] separates generation and validation into distinct models.
-PET-SQL [52] refines outputs through multi-stage prompting, while MCS-SQL [55] uses execution-guided selection as the final module after multiple SQL generation.
-With the increasing complexity of recently proposed benchmarks, agentic frameworks like ReFoRCE [61], LinkAlign [62] and Spider-Agent [42] leveraged well-designed step-by-step action plans for LLMs to systematically reason toward the final SQL.
+DIN-SQL [pourreza2023dinsql] integrates a four-stage pipeline, including schema linking, classification, SQL generation, and self-correction.
+Coder-Reviewer [zhang2022coder] separates generation and validation into distinct models.
+PET-SQL [li2024petsql] refines outputs through multi-stage prompting, while MCS-SQL [lee2025mcs] uses execution-guided selection as the final module after multiple SQL generation.
+With the increasing complexity of recently proposed benchmarks, agentic frameworks like ReFoRCE [deng2025reforce], LinkAlign [wang2025linkalign] and Spider-Agent [lei2025spider2] leveraged well-designed step-by-step action plans for LLMs to systematically reason toward the final SQL.
 2.
 Question decomposition: Complexity is reduced by splitting user questions into intermediate sub-questions via question decomposition.
-QDecomp [47] uses Least-to-Most prompting, breaking complex questions into step-by-step sub-questions to facilitate SQL generation.
-SGU-SQL [30] employs graph-based structure linking and meta-operators for decomposing grammar-aware sub-questions.
-Metadata-driven approaches, such as MetaSQL [50] and TA-SQL [54], further decompose questions using metadata conditioning and task-aligned schema linking.
-Meanwhile, frameworks like DIN-SQL [7] and MAC-SQL [48] use question decomposition as auxiliary modules but primarily focus on module-level collaboration.
+QDecomp [tai2023exploring] uses Least-to-Most prompting, breaking complex questions into step-by-step sub-questions to facilitate SQL generation.
+SGU-SQL [zhang2024sgusql] employs graph-based structure linking and meta-operators for decomposing grammar-aware sub-questions.
+Metadata-driven approaches, such as MetaSQL [fan2024metasql] and TA-SQL [qu2024generation], further decompose questions using metadata conditioning and task-aligned schema linking.
+Meanwhile, frameworks like DIN-SQL [pourreza2023dinsql] and MAC-SQL [wang2024macsql] use question decomposition as auxiliary modules but primarily focus on module-level collaboration.
 IV-A3 $\mathbf{C}_{2}$-Prompt Optimization
@@
 Advanced few-shot sampling: This strategy optimizes example selection through semantic similarity, diversity, or domain alignment.
-DESEM [63] and retrieval-augmented frameworks [66, 71, 58] de-semanticize questions to retrieve examples with matching intent skeletons or vector embeddings.
-DAIL-SQL [8] balances quality and quantity by masking domain-specific tokens and ranking examples based on the Euclidean distance between questions and SQL queries, while SAFE-SQL [73] combines skeleton masking with self-augmentation.
-ODIS [67] hybridizes out-of-domain and synthetic in-domain examples for robustness, and FUSED [69] enhances diversity through iterative demonstration synthesis.
-Dynamic selection is explored in ACT-SQL [68] and MCS-SQL [55], which adaptively retrieve examples based on defined similarity scores or masked similarity scores.
+DESEM [guo2023desem] and retrieval-augmented frameworks [guo2023ragsql, thorpe2024dubo, maamari2024the] de-semanticize questions to retrieve examples with matching intent skeletons or vector embeddings.
+DAIL-SQL [gao2023dailsql] balances quality and quantity by masking domain-specific tokens and ranking examples based on the Euclidean distance between questions and SQL queries, while SAFE-SQL [lee2025safe] combines skeleton masking with self-augmentation.
+ODIS [chang2023selective] hybridizes out-of-domain and synthetic in-domain examples for robustness, and FUSED [wang2024fused] enhances diversity through iterative demonstration synthesis.
+Dynamic selection is explored in ACT-SQL [zhang2023actsql] and MCS-SQL [lee2025mcs], which adaptively retrieve examples based on defined similarity scores or masked similarity scores.
 2.
 Schema augmentation: These methods refine schema representations to reduce redundancy and improve relevance.
-C3 [28] distills question-relevant schemas by removing unrelated tables or columns, while RSL-SQL [59] simplifies schemas using NL descriptions.
-ReFoRCE [61] employs table compression and column exploration to manage long-context schema complexity.
-Gen-SQL [60] bridges NL-SQL gaps by generating pseudo-schemas that align with user questions.
-QDecomp [47] and SD+SA+Voting [65] further augment prompts with schema-aware semantic and structural hints.
-For more complex databases, LinkAlign [62] introduces a schema-based question rewriting strategy, combined with a retrieve-then-filter schema selection mechanism for accurate linking.
+C3 [dong2023c3] distills question-relevant schemas by removing unrelated tables or columns, while RSL-SQL [cao2024rsl] simplifies schemas using NL descriptions.
+ReFoRCE [deng2025reforce] employs table compression and column exploration to manage long-context schema complexity.
+Gen-SQL [shi2025gen] bridges NL-SQL gaps by generating pseudo-schemas that align with user questions.
+QDecomp [tai2023exploring] and SD+SA+Voting [nan2023enhancing] further augment prompts with schema-aware semantic and structural hints.
+For more complex databases, LinkAlign [wang2025linkalign] introduces a schema-based question rewriting strategy, combined with a retrieve-then-filter schema selection mechanism for accurate linking.
 3.
 External knowledge integration: This group enhances prompts with domain-specific knowledge or feedback mechanisms.
-Knowledge-to-SQL [70] trains a DELLM to inject external knowledge about question-database relationships, while ICRL [72] integrates reinforcement learning with schema-graph-derived knowledge bases.
-MAC-SQL [48] and DIN-SQL [7] implicitly utilize decomposition patterns as procedural knowledge to guide generation.
-C3 [28] and SAFE-SQL [73] incorporate calibration bias and self-augmented feedback loops, respectively, to iteratively refine outputs.
+Knowledge-to-SQL [hong2024knowledgetosql] trains a DELLM to inject external knowledge about question-database relationships, while ICRL [toteja2025context] integrates reinforcement learning with schema-graph-derived knowledge bases.
+MAC-SQL [wang2024macsql] and DIN-SQL [pourreza2023dinsql] implicitly utilize decomposition patterns as procedural knowledge to guide generation.
+C3 [dong2023c3] and SAFE-SQL [lee2025safe] incorporate calibration bias and self-augmented feedback loops, respectively, to iteratively refine outputs.
 IV-A4 $\mathbf{C}_{3}$-Reasoning Enhancement
@@
 Chain-of-thought (CoT)-based reasoning: CoT-based methods guide LLMs in generating intermediate reasoning steps before predicting SQLs.
-While standard CoT prompts (e.g., “Let’s think step by step”) show limited effectiveness in text-to-SQL [8, 1], adaptations help bridge this gap.
-ACT-SQL [68] automates CoT example generation by linking question slices to database columns via similarity matching.
-QDecomp [47] restructures SQL clauses into logical steps ordered by execution flow, reducing error propagation compared to generic CoT.
-CHASE-SQL [77] extends this with multi-path reasoning for candidate generation using “divide and conquer” CoT strategy.
+While standard CoT prompts (e.g., “Let’s think step by step”) show limited effectiveness in text-to-SQL [gao2023dailsql, li2023BIRD], adaptations help bridge this gap.
+ACT-SQL [zhang2023actsql] automates CoT example generation by linking question slices to database columns via similarity matching.
+QDecomp [tai2023exploring] restructures SQL clauses into logical steps ordered by execution flow, reducing error propagation compared to generic CoT.
+CHASE-SQL [pourreza2025chasesql] extends this with multi-path reasoning for candidate generation using “divide and conquer” CoT strategy.
 2.
 Consistency-driven reasoning: These methods improve reliability through candidate diversity and agreement metrics before execution.
-Self-consistency [126] is employed in frameworks like C3 [28], DAIL-SQL [8], and SuperSQL [57], where majority voting selects the most frequent SQL candidate based on syntax or structural similarity.
-SD+SA+Voting [65] refines this by filtering candidates using schema-aware rules (e.g., valid joins, clause ordering) before voting.
+Self-consistency [wang2023selfconsistency] is employed in frameworks like C3 [dong2023c3], DAIL-SQL [gao2023dailsql], and SuperSQL [li2024dawn], where majority voting selects the most frequent SQL candidate based on syntax or structural similarity.
+SD+SA+Voting [nan2023enhancing] refines this by filtering candidates using schema-aware rules (e.g., valid joins, clause ordering) before voting.
 Unlike execution-dependent methods, these strategies prioritize logical consistency and syntactic validity over runtime feedback, making them lightweight but less robust to semantic errors.
@@
 Tool-augmented and agentic reasoning: Hybrid techniques combine SQL generation with external computation or tools.
-SQL-CRAFT [75] employs Program-of-Thoughts [127] (PoT) prompting, requiring LLMs to generate Python code alongside SQL for arithmetic-heavy queries, improving performance on datasets like BIRD [1].
-FUXI [76] integrates dedicated tools for schema linking and syntax checking, dynamically invoking them during generation.
+SQL-CRAFT [xia2024sqlcraft] employs Program-of-Thoughts [chen2023program] (PoT) prompting, requiring LLMs to generate Python code alongside SQL for arithmetic-heavy queries, improving performance on datasets like BIRD [li2023BIRD].
+FUXI [gu2024fuxi] integrates dedicated tools for schema linking and syntax checking, dynamically invoking them during generation.
 These approaches differ from pure prompting strategies by explicitly bridging LLM reasoning with external symbolic systems.
-Also, recent frameworks such as ReFoRCE [61] and Spider-Agent [42] adopt step-by-step agentic reasoning, where LLMs act as specialized agents that iteratively plan actions to handle complex database scenarios.
+Also, recent frameworks such as ReFoRCE [deng2025reforce] and Spider-Agent [lei2025spider2] adopt step-by-step agentic reasoning, where LLMs act as specialized agents that iteratively plan actions to handle complex database scenarios.
 IV-A5 $\mathbf{C}_{4}$-Execution Refinement
@@
 Feedback-driven regeneration: These methods iteratively refine SQL queries using execution errors or intermediate results.
-Self-Debugging [80] enables LLMs to autonomously correct errors via natural language explanations of failed executions, while DESEM [63] uses error-type-specific prompts with termination criteria to prevent infinite loops.
-Multi-agent frameworks like MAC-SQL [48] and CHASE-SQL [77] utilize agents (e.g., Refiner, Query Fixer) to diagnose SQLite errors and regenerate SQL fixes.
-SQL-CRAFT [75] introduces adaptive control to balance correction depth, preventing over-refinement.
-Schema-focused methods like RSL-SQL [59] and DIN-SQL [7] use generic or gentle prompts to help LLMs identify schema misalignments during regeneration.
-Knowledge-to-SQL [70] optimizes its DELLM model through preference learning from execution feedback, while FUXI [76] integrates error feedback into tool-based reasoning.
-E-SQL [81] enhances questions with execution-aware candidate predictions.
-Agentic frameworks [42] also incorporate feedback-driven regeneration, where specialized agents iteratively analyze execution feedback to plan corrective actions to refine SQLs.
+Self-Debugging [chen2024selfdebugging] enables LLMs to autonomously correct errors via natural language explanations of failed executions, while DESEM [guo2023desem] uses error-type-specific prompts with termination criteria to prevent infinite loops.
+Multi-agent frameworks like MAC-SQL [wang2024macsql] and CHASE-SQL [pourreza2025chasesql] utilize agents (e.g., Refiner, Query Fixer) to diagnose SQLite errors and regenerate SQL fixes.
+SQL-CRAFT [xia2024sqlcraft] introduces adaptive control to balance correction depth, preventing over-refinement.
+Schema-focused methods like RSL-SQL [cao2024rsl] and DIN-SQL [pourreza2023dinsql] use generic or gentle prompts to help LLMs identify schema misalignments during regeneration.
+Knowledge-to-SQL [hong2024knowledgetosql] optimizes its DELLM model through preference learning from execution feedback, while FUXI [gu2024fuxi] integrates error feedback into tool-based reasoning.
+E-SQL [caferouglu2024esql] enhances questions with execution-aware candidate predictions.
+Agentic frameworks [lei2025spider2] also incorporate feedback-driven regeneration, where specialized agents iteratively analyze execution feedback to plan corrective actions to refine SQLs.
 2.
 Execution-guided selection: This strategy leverages database execution results to select or rank candidates.
-ReFoRCE [61] and PET-SQL [52] prune invalid SQL queries that fail execution, then apply bias-aware voting (e.g., difficulty-weighted or efficiency-prioritized as in MCS-SQL [55]).
-Probabilistic verification distinguishes methods: MRC-EXEC [78] minimizes Bayes risk by comparing execution outputs to expected results, and LEVER [79] trains a verifier to predict correctness likelihoods from execution traces.
-Hybrid approaches like the retrieval-augmented framework [66] use execution feedback for iterative SQL revision through semantic gap analysis, bridging generation and selection.
+ReFoRCE [deng2025reforce] and PET-SQL [li2024petsql] prune invalid SQL queries that fail execution, then apply bias-aware voting (e.g., difficulty-weighted or efficiency-prioritized as in MCS-SQL [lee2025mcs]).
+Probabilistic verification distinguishes methods: MRC-EXEC [shi2022natural] minimizes Bayes risk by comparing execution outputs to expected results, and LEVER [ni2023lever] trains a verifier to predict correctness likelihoods from execution traces.
+Hybrid approaches like the retrieval-augmented framework [guo2023ragsql] use execution feedback for iterative SQL revision through semantic gap analysis, bridging generation and selection.
 IV-B Fine-tuning
@@
 Category Adopted by Applied LLMs Dataset EX EM VES Release Time Publication Venue
-Enhanced Architecture CLLMs [82] DeepSeek* [11] Mar-2024 ICML’24
-Pre-training CodeS [24] StarCoder [11, 1] Feb-2024 SIGMOD’24
-Data Augmentation DAIL-SQL [8] LLaMA* [11, 38] Aug-2023 VLDB’24
-Symbol-LLM [83] CodeLLaMA [11] Nov-2023 ACL’24
-CodeS [24] StarCoder [11, 1, 42] Feb-2024 SIGMOD’24
-StructLM [84] CodeLLaMA [11] Feb-2024 COLM’24
-Dubo-SQL [71] ChatGPT [1] Apr-2024 arXiv’24
-Distillery [58] LLaMA-3.1* [1] Aug-2024 NeurIPS’24 Workshop
-XiYan-SQL [85] Qwen2.5-Coder [11, 1] Nov-2024 arXiv’24
-SHARE [86] LLaMA-3.1* [11, 1, 36, 38] Jun-2025 ACL’25
-Multi-task Tuning SQL-LLaMA [48] CodeLLaMA [11, 1] Dec-2023 COLING’25
-DTS-SQL [88] Mistral* [11, 37] Feb-2024 EMNLP’24 Findings
-KaSLA [29] DeepSeek-Coder* [11, 1] Oct-2024 arXiv’24
-MSc-SQL [89] Gemma-2 [11, 1] Oct-2024 NAACL’25
-XiYan-SQL [85] Qwen2.5-Coder [11, 1] Nov-2024 arXiv’24
-ROUTE [90] Qwen2.5* [11, 1, 37, 36, 38] Feb-2025 ICLR’25
-SHARE [86] LLaMA-3.1* [11, 1, 36, 38] Jun-2025 ACL’25
-Supervised Fine-tuning (SFT) is the widely-used approach within the FT paradigm for training LLMs [27, 113].
-For open-source LLMs, like LLaMA [117] and Qwen [118] series, the most straightforward way to enable the model to adapt to a specific domain is to apply SFT over collected domain-specific data.
-The SFT phase typically serves as an initial step in a well-designed training pipeline [128, 129], including LLM-based text-to-SQL training.
+Enhanced Architecture CLLMs [kou2024cllms] DeepSeek* [yu2018Spider] Mar-2024 ICML’24
+Pre-training CodeS [li2024codes] StarCoder [yu2018Spider, li2023BIRD] Feb-2024 SIGMOD’24
+Data Augmentation DAIL-SQL [gao2023dailsql] LLaMA* [yu2018Spider, deng2021Spider-Realistic] Aug-2023 VLDB’24
+Symbol-LLM [xu2024symbolllm] CodeLLaMA [yu2018Spider] Nov-2023 ACL’24
+CodeS [li2024codes] StarCoder [yu2018Spider, li2023BIRD, lei2025spider2] Feb-2024 SIGMOD’24
+StructLM [zhuang2024structlm] CodeLLaMA [yu2018Spider] Feb-2024 COLM’24
+Dubo-SQL [thorpe2024dubo] ChatGPT [li2023BIRD] Apr-2024 arXiv’24
+Distillery [maamari2024the] LLaMA-3.1* [li2023BIRD] Aug-2024 NeurIPS’24 Workshop
+XiYan-SQL [gao2024xiyan] Qwen2.5-Coder [yu2018Spider, li2023BIRD] Nov-2024 arXiv’24
+SHARE [qu2025share] LLaMA-3.1* [yu2018Spider, li2023BIRD, gan2021Spider-DK, deng2021Spider-Realistic] Jun-2025 ACL’25
+Multi-task Tuning SQL-LLaMA [wang2024macsql] CodeLLaMA [yu2018Spider, li2023BIRD] Dec-2023 COLING’25
+DTS-SQL [pourreza2024dtssql] Mistral* [yu2018Spider, gan2021Spider-SYN] Feb-2024 EMNLP’24 Findings
+KaSLA [yuan2025knapsack] DeepSeek-Coder* [yu2018Spider, li2023BIRD] Oct-2024 arXiv’24
+MSc-SQL [gorti2024mscsql] Gemma-2 [yu2018Spider, li2023BIRD] Oct-2024 NAACL’25
+XiYan-SQL [gao2024xiyan] Qwen2.5-Coder [yu2018Spider, li2023BIRD] Nov-2024 arXiv’24
+ROUTE [qin2025route] Qwen2.5* [yu2018Spider, li2023BIRD, gan2021Spider-SYN, gan2021Spider-DK, deng2021Spider-Realistic] Feb-2025 ICLR’25
+SHARE [qu2025share] LLaMA-3.1* [yu2018Spider, li2023BIRD, gan2021Spider-DK, deng2021Spider-Realistic] Jun-2025 ACL’25
+Supervised Fine-tuning (SFT) is the widely-used approach within the FT paradigm for training LLMs [zhao2023survey, zheng2024llamafactory].
+For open-source LLMs, like LLaMA [touvron2023llama2] and Qwen [bai2023qwen] series, the most straightforward way to enable the model to adapt to a specific domain is to apply SFT over collected domain-specific data.
+The SFT phase typically serves as an initial step in a well-designed training pipeline [rafailo2023direct, ouyang2022training], including LLM-based text-to-SQL training.
 Consider the prompt defined in Eq.
@@
 $\mathcal{L}_{\text{SFT}}=-\sum_{\left(P,\,Y\right)\in\mathcal{D}}\sum_{t=1}^{T}\log\text{Pr}_{\pi}\left(y_{t}\mid P,y_{<t}\right).$ (7)
-As a vanilla fine-tuning approach for LLM-based text-to-SQL, SFT is widely adopted in text-to-SQL research across various open-source LLMs [24, 120, 8].
+As a vanilla fine-tuning approach for LLM-based text-to-SQL, SFT is widely adopted in text-to-SQL research across various open-source LLMs [li2024codes, zhang2024benchmarking, gao2023dailsql].
 FT paradigms are often more foundational in LLM-based text-to-SQL than ICL, which is the mainstream approach in the community.
@@
 IV-B1 Enhanced Architecture
-The widely-used Generative Pre-trained Transformer (GPT) framework [105] employs a decoder-only transformer architecture and conventional auto-regressive decoding for text generation.
+The widely-used Generative Pre-trained Transformer (GPT) framework [radford2018improving] employs a decoder-only transformer architecture and conventional auto-regressive decoding for text generation.
 With advancements in the GPT framework, enhancements in LLM-based text-to-SQL involve modifications of the standard transformer backbone.
@@
 In traditional text-to-SQL research, this tailored architecture has effectively enhanced PLM-based performance.
-Techniques include schema-aware encoding [20], optimized attention [130], and targeted decoding [131], often relying on graph structures.
-Most recently, the speed of generating SQL queries by LLMs is notably slower than traditional PLM-based methods [19, 26], posing a challenge to building high-efficiency local NLIDBs.
-As a solution, CLLMs [82] are designed to address this challenge with specialized decoding, offering speedup efficiency for SQL generation and reducing high latency.
+Techniques include schema-aware encoding [li2023graphix], optimized attention [gong2024graph], and targeted decoding [cao2023heterogeneous], often relying on graph structures.
+Most recently, the speed of generating SQL queries by LLMs is notably slower than traditional PLM-based methods [li2023resdsql, yang2024harnessing], posing a challenge to building high-efficiency local NLIDBs.
+As a solution, CLLMs [kou2024cllms] are designed to address this challenge with specialized decoding, offering speedup efficiency for SQL generation and reducing high latency.
 IV-B2 Pre-training
-Pre-training is a fundamental phase of the complete training process, aimed at acquiring text generation capabilities through auto-regressive training on extensive data [132].
-Code-specific LLMs (e.g., CodeLLaMA [87], StarCoder [133]) are pre-trained on code data, with diverse programming languages enabling them to generate code per user instructions [124, 134].
+Pre-training is a fundamental phase of the complete training process, aimed at acquiring text generation capabilities through auto-regressive training on extensive data [erhan2010does].
+Code-specific LLMs (e.g., CodeLLaMA [roziere2023codellama], StarCoder [li2023starcoder]) are pre-trained on code data, with diverse programming languages enabling them to generate code per user instructions [chen2021evaluating, yufei2023aligning].
 For SQL-specific pre-training, a key challenge is that SQL/database-related content makes up only a small part of the pre-training corpus.
 As a result, the open-source LLMs do not acquire a solid understanding of converting NL questions to SQL during their pre-training.
-PLM-based methods propose initial attempts in pre-training paradigms, such as STAR [135] and multitask approaches [136], which improve SQL generation by leveraging SQL-guided and multitask pre-training strategies.
-For the most recent LLM-based text-to-SQL, the pre-training phase of the CodeS [24] model consists of three stages of incremental pre-training.
-Starting from a code-specific LLM [133], CodeS are incrementally pre-trained on a hybrid training corpus, including SQL-related data, NL-to-Code data, and NL-related data.
+PLM-based methods propose initial attempts in pre-training paradigms, such as STAR [cai2022star] and multitask approaches [giaquinto2023multitask], which improve SQL generation by leveraging SQL-guided and multitask pre-training strategies.
+For the most recent LLM-based text-to-SQL, the pre-training phase of the CodeS [li2024codes] model consists of three stages of incremental pre-training.
+Starting from a code-specific LLM [li2023starcoder], CodeS are incrementally pre-trained on a hybrid training corpus, including SQL-related data, NL-to-Code data, and NL-related data.
 The text-to-SQL understanding and performance are significantly improved.
 IV-B3 Data Augmentation
-During the fine-tuning process, the quality of training labels is the most straightforward factor affecting the model’s performance [137].
-Training with low-quality or absent labels is “making bricks without straw.” Using high-quality or augmented data for fine-tuning consistently outperforms the meticulously designed methods relying on low-quality or raw data [93, 27].
+During the fine-tuning process, the quality of training labels is the most straightforward factor affecting the model’s performance [song2022learningnoise].
+Training with low-quality or absent labels is “making bricks without straw.” Using high-quality or augmented data for fine-tuning consistently outperforms the meticulously designed methods relying on low-quality or raw data [deng2022recent, zhao2023survey].
 Data-augmented fine-tuning in text-to-SQL has made substantial progress by focusing on enhancing data quality and efficiency during the SFT process.
-DAIL-SQL [8] is designed as an ICL framework that uses a sampling strategy to enhance few-shot instances.
+DAIL-SQL [gao2023dailsql] is designed as an ICL framework that uses a sampling strategy to enhance few-shot instances.
 Incorporating these sampled instances in the SFT process boosts the performance of open-source LLMs.
-Symbol-LLM [83] proposes injection and infusion stages for data-augmented instruction tuning.
-CodeS [24] augments the training data with bi-directional generation aided by ChatGPT.
-StructLM [84] trains on multiple structured knowledge tasks to improve overall capability.
-Dubo-SQL [71] created concise prompt templates for fine-tuning, which reduces costs and simultaneously achieves record performance.
-Distillery [58] employs an iterative failure-aware approach, selecting new training samples based on error analysis.
-XiYan-SQL [85] uses basic-syntax SFT and generation-enhanced SFT, employing hybrid augmented training data to fine-tune the SQL generator.
-SHARE [86] designed a rule-based error perturbation to augment schema linking and final SQL inference based on the decomposed trajectories.
+Symbol-LLM [xu2024symbolllm] proposes injection and infusion stages for data-augmented instruction tuning.
+CodeS [li2024codes] augments the training data with bi-directional generation aided by ChatGPT.
+StructLM [zhuang2024structlm] trains on multiple structured knowledge tasks to improve overall capability.
+Dubo-SQL [thorpe2024dubo] created concise prompt templates for fine-tuning, which reduces costs and simultaneously achieves record performance.
+Distillery [maamari2024the] employs an iterative failure-aware approach, selecting new training samples based on error analysis.
+XiYan-SQL [gao2024xiyan] uses basic-syntax SFT and generation-enhanced SFT, employing hybrid augmented training data to fine-tune the SQL generator.
+SHARE [qu2025share] designed a rule-based error perturbation to augment schema linking and final SQL inference based on the decomposed trajectories.
 IV-B4 Multi-task Tuning
@@
 Proprietary models in ICL-based methods boast a large number of parameters, surpassing the scale of open-source models used in fine-tuning.
-These models are adept at performing assigned sub-tasks effectively, facilitated by mechanisms like few-shot learning [28, 48].
+These models are adept at performing assigned sub-tasks effectively, facilitated by mechanisms like few-shot learning [dong2023c3, wang2024macsql].
 To replicate this success, it is essential to assign sub-tasks strategically to open-source models, such as generating external knowledge, schema linking, and execution-based refinement, with appropriate training.
 This involves constructing specific datasets for fine-tuning, ultimately aiding in the final SQL generation.
-SQL-LLaMA [48] is a fine-tuned LLM framework utilizing multi-agent collaboration.
-DTS-SQL [88] introduced a two-stage multi-task text-to-SQL fine-tuning framework, incorporating a schema-linking pre-generation task before the final SQL generation.
-Similarly, KaSLA [29] employed a two-stage framework using knapsack optimization as a fine-tuning strategy for the schema linking task, and SHARE [86] adopted three sequential models for SQL correction.
-MSc-SQL [89] trained separate models for schema linking, SQL generation, and critiquing the predicted SQL query.
-The ROUTE [90] framework involved the SFT process over noise filtering, data synthesis, and continued writing, validating the effectiveness of the multi-task tuning pipeline through collaboration prompting.
+SQL-LLaMA [wang2024macsql] is a fine-tuned LLM framework utilizing multi-agent collaboration.
+DTS-SQL [pourreza2024dtssql] introduced a two-stage multi-task text-to-SQL fine-tuning framework, incorporating a schema-linking pre-generation task before the final SQL generation.
+Similarly, KaSLA [yuan2025knapsack] employed a two-stage framework using knapsack optimization as a fine-tuning strategy for the schema linking task, and SHARE [qu2025share] adopted three sequential models for SQL correction.
+MSc-SQL [gorti2024mscsql] trained separate models for schema linking, SQL generation, and critiquing the predicted SQL query.
+The ROUTE [qin2025route] framework involved the SFT process over noise filtering, data synthesis, and continued writing, validating the effectiveness of the multi-task tuning pipeline through collaboration prompting.
 V Summary
@@
 V-A1 Linguistic Complexity and Ambiguity
-Traditional rule-based systems, like GRAPPA [23], decompose questions via syntactic templates but fail to resolve ambiguous phrasing.
-Neural models like RyenSQL [17] improve lexical matching but lack contextual reasoning.
-In contrast, LLM-based methods such as DAIL-SQL [8] leverage chain-of-thought prompting to infer dependencies stepwise, while C3-SQL [28] iteratively validates SQL drafts against paraphrased questions.
+Traditional rule-based systems, like GRAPPA [yu2021grappa], decompose questions via syntactic templates but fail to resolve ambiguous phrasing.
+Neural models like RyenSQL [choi2021ryansql] improve lexical matching but lack contextual reasoning.
+In contrast, LLM-based methods such as DAIL-SQL [gao2023dailsql] leverage chain-of-thought prompting to infer dependencies stepwise, while C3-SQL [dong2023c3] iteratively validates SQL drafts against paraphrased questions.
 While traditional methods offer lightweight disambiguation, their reliance on limited training data restricts their ability to handle unseen phrasings.
@@
 LLM-based methods enhance schema linking through in-context learning and structured attention.
-For example, DIN-SQL [7] serializes schema into natural language prompts with foreign-key hints, while KaSLA [29] mirrors human prioritization of schema elements, reducing the common issues of omissions and redundancies in schema linking.
+For example, DIN-SQL [pourreza2023dinsql] serializes schema into natural language prompts with foreign-key hints, while KaSLA [yuan2025knapsack] mirrors human prioritization of schema elements, reducing the common issues of omissions and redundancies in schema linking.
 While rule-based methods are interpretable, their rigidity limits adaptability.
@@
 Handling operations like nested sub-queries or window functions requires labor-intensive template engineering.
-For instance, previous rule-based methods [9] manually define 40+ SQL grammar rules to map question patterns, which could not be generalized to unseen formulations.
-Neural semantic parsers like PICARD [138] constrain decoding via transition-based incremental parsing, ensuring syntactically valid output by rejecting invalid tokens at each step.
+For instance, previous rule-based methods [li2014constructing] manually define 40+ SQL grammar rules to map question patterns, which could not be generalized to unseen formulations.
+Neural semantic parsers like PICARD [scholak2021picard] constrain decoding via transition-based incremental parsing, ensuring syntactically valid output by rejecting invalid tokens at each step.
 LLM-based methods generate rare SQLs by combining parametric knowledge from code-centric pretraining with few-shot demonstrations.
-For example, DAIL-SQL [8] injects prompts with examples of window functions, enabling the model to generalize to unseen formulations.
-Traditional methods like PICARD [138] guarantee syntactic validity via grammar-constrained decoders but cannot innovate on rare and complex SQL operations.
+For example, DAIL-SQL [gao2023dailsql] injects prompts with examples of window functions, enabling the model to generalize to unseen formulations.
+Traditional methods like PICARD [scholak2021picard] guarantee syntactic validity via grammar-constrained decoders but cannot innovate on rare and complex SQL operations.
 LLM-based methods provide novel solutions but may produce invalid SQL.
-The most current agentic reasoning [42, 61] mitigates invalid outputs through multi-agent actions, but still suffers from error accumulation.
+The most current agentic reasoning [lei2025spider2, deng2025reforce] mitigates invalid outputs through multi-agent actions, but still suffers from error accumulation.
 A synergistic approach would integrate LLM’s creativity with traditional validation: first, generate candidate queries with an LLM, then apply PICARD-like grammars to reject invalid tokens during beam search.
@@
 V-A4 Cross-Domain Generalization
-Rule-based systems require manual template adjustments per domain, while PLM-based models like S2SQL [139] rely on domain-specific fine-tuning.
-LLM-based methods like CodeS [24] adapt to niche domains (e.g., biomedical terminology) via code-centric pretraining.
+Rule-based systems require manual template adjustments per domain, while PLM-based models like S2SQL [hui2022s] rely on domain-specific fine-tuning.
+LLM-based methods like CodeS [li2024codes] adapt to niche domains (e.g., biomedical terminology) via code-centric pretraining.
 Traditional domain adaptation methods require moderate fine-tuning data to align latent spaces but provide precise control over domain-specific mappings.
@@
 IV-A5) ensures runtime validity through feedback loops but relies heavily on database accessibility.
-Recent methods like SuperSQL [57], Spider-Agent [42], and ReFoRCE [61] illustrate a trend toward hybrid integration-combining decomposition’s structural rigor with prompt optimization’s efficiency, while layering reasoning enhancement and execution consistency, which are implemented through well-designed agentic reasoning actions.
+Recent methods like SuperSQL [li2024dawn], Spider-Agent [lei2025spider2], and ReFoRCE [deng2025reforce] illustrate a trend toward hybrid integration-combining decomposition’s structural rigor with prompt optimization’s efficiency, while layering reasoning enhancement and execution consistency, which are implemented through well-designed agentic reasoning actions.
 This convergence suggests that future ICL methods will increasingly unify cross-category techniques into cohesive frameworks, balancing accuracy, efficiency, and robustness through systematic collaboration.
@@
 Fine-tuning (FT) approaches highlight the trade-off between specialization and generalization.
-Enhanced architectures such as CLLMs [82] reduce latency, while pre-training approaches like CodeS [24] offer robust SQL foundations.
+Enhanced architectures such as CLLMs [kou2024cllms] reduce latency, while pre-training approaches like CodeS [li2024codes] offer robust SQL foundations.
 Consequently, both approaches are constrained by their demands on computational resources and hardware devices.
-Data augmentation (DAIL-SQL [8], XiYan-SQL [85]) improves data efficiency but risks overfitting to synthetic patterns.
-Multi-task tuning (ROUTE [90]) boosts multi-step reasoning but complicates training convergence.
+Data augmentation (DAIL-SQL [gao2023dailsql], XiYan-SQL [gao2024xiyan]) improves data efficiency but risks overfitting to synthetic patterns.
+Multi-task tuning (ROUTE [qin2025route]) boosts multi-step reasoning but complicates training convergence.
 While current methods focus on isolated improvements, the field lacks holistic frameworks to combine architectural technical contributions.
@@
 The paradigms present asymmetric value propositions.
-ICL methods leverage proprietary LLMs (e.g., GPT-4 [107]) for rapid adaptation but face constraints in controllability and domain customization.
-FT methods empower open-source models (e.g., CodeLLaMA [87]) with targeted SQL capabilities through fine-tuning but require substantial engineering for data curation and training optimization.
-Importantly, FT methods enhance foundational model competence: CodeS’s [24] incremental pre-training boosts question and schema understanding, while SQL-LLaMA’s [48] multi-agent tuning supports complex compositional reasoning.
+ICL methods leverage proprietary LLMs (e.g., GPT-4 [achiam2023gpt4]) for rapid adaptation but face constraints in controllability and domain customization.
+FT methods empower open-source models (e.g., CodeLLaMA [roziere2023codellama]) with targeted SQL capabilities through fine-tuning but require substantial engineering for data curation and training optimization.
+Importantly, FT methods enhance foundational model competence: CodeS’s [li2024codes] incremental pre-training boosts question and schema understanding, while SQL-LLaMA’s [wang2024macsql] multi-agent tuning supports complex compositional reasoning.
 Nonetheless, ICL methods’ prompt-based flexibility remain unmatched for addressing emerging database environments without retraining.
@@
 The text-to-SQL implemented by LLMs is expected to perform generalization and robustness across complex scenarios in real-world applications.
-Despite recent advances having made substantial progress in robustness-specific datasets [34, 38], its performance still falls short of practical application [1].
+Despite recent advances having made substantial progress in robustness-specific datasets [pi2022ADVETA, deng2021Spider-Realistic], its performance still falls short of practical application [li2023BIRD].
 There are still challenges that are expected to be overcome in future studies.
-From the user aspect, there is a phenomenon that the user is not always a clear question proposer, which means the user questions might not have the exact database value and also can be varied from the standard datasets, the synonyms, typos, and vague expressions could be included [37].
+From the user aspect, there is a phenomenon that the user is not always a clear question proposer, which means the user questions might not have the exact database value and also can be varied from the standard datasets, the synonyms, typos, and vague expressions could be included [gan2021Spider-SYN].
 For instance, the models are trained on clear indicative questions with concrete expressions in the fine-tuning paradigm.
-Since the model has not learned the mapping of realistic questions to the corresponding database, this leads to a knowledge gap when applied to real-world scenarios [1].
-As reported in the corresponding evaluations of the dataset with synonym and incomplete instruction [6, 47], the SQL queries generated by ChatGPT contain around 40% incorrect execution, which is 10% lower than the original evaluation [47].
-For the evaluation in real-world enterprise tasks [42], current state-of-the-art models solve only about 21.3% of these tasks, far below their performance on earlier benchmarks [11, 1].
+Since the model has not learned the mapping of realistic questions to the corresponding database, this leads to a knowledge gap when applied to real-world scenarios [li2023BIRD].
+As reported in the corresponding evaluations of the dataset with synonym and incomplete instruction [rajkumar2022evaluating, tai2023exploring], the SQL queries generated by ChatGPT contain around 40% incorrect execution, which is 10% lower than the original evaluation [tai2023exploring].
+For the evaluation in real-world enterprise tasks [lei2025spider2], current state-of-the-art models solve only about 21.3% of these tasks, far below their performance on earlier benchmarks [yu2018Spider, li2023BIRD].
 Simultaneously, the fine-tuning with local text-to-SQL datasets may contain non-standardized samples and labels.
@@
 Besides that, the adaptation study of fine-tuned open-source LLM to local small-size datasets can be potentially beneficial.
-Furthermore, the extensions on multi-lingual [39, 43] and multi-modal scenarios [140] should be studied in future research, which will benefit more language groups and help build more general database interfaces.
+Furthermore, the extensions on multi-lingual [min2019CSpider, tuan2020pilot] and multi-modal scenarios [song2024speech] should be studied in future research, which will benefit more language groups and help build more general database interfaces.
 VI-B Computational Efficiency
-The computational efficiency is determined by the inference speed and the cost of computational resources, which is worth considering in both application and research work [82, 68].
-With the increasing complexity of databases in up-to-date text-to-SQL benchmarks [13, 1, 42], databases will carry more information (including more tables and columns), and the token length of the database schema will correspondingly increase, raising a series of challenges.
+The computational efficiency is determined by the inference speed and the cost of computational resources, which is worth considering in both application and research work [kou2024cllms, zhang2023actsql].
+With the increasing complexity of databases in up-to-date text-to-SQL benchmarks [pourreza2023evaluating, li2023BIRD, lei2025spider2], databases will carry more information (including more tables and columns), and the token length of the database schema will correspondingly increase, raising a series of challenges.
 Dealing with an ultra-complex database, taking the corresponding schema as input may encounter the challenge that the cost of calling proprietary LLMs will significantly increase, potentially exceeding the model’s maximum token length, especially with the implementation of open-source models that have shorter context lengths.
-Meanwhile, another obvious challenge is that most works use the full schema as model input, which introduces significant redundancy [48].
-Providing LLMs with a precise question-related filtered schema directly from the user end to reduce cost and redundancy is a potential solution to improve computational efficiency [28].
+Meanwhile, another obvious challenge is that most works use the full schema as model input, which introduces significant redundancy [wang2024macsql].
+Providing LLMs with a precise question-related filtered schema directly from the user end to reduce cost and redundancy is a potential solution to improve computational efficiency [dong2023c3].
 Designing an accurate method for schema filtering remains a future direction.
-Although the in-context learning paradigm achieves promising accuracy, as a computational efficiency concern, the well-designed methods with the multi-stage framework or extended context increasing the number of API calls to enhance performance has simultaneously led to a substantial rise in costs [7].
+Although the in-context learning paradigm achieves promising accuracy, as a computational efficiency concern, the well-designed methods with the multi-stage framework or extended context increasing the number of API calls to enhance performance has simultaneously led to a substantial rise in costs [pourreza2023dinsql].
 Recent agentic reasoning approaches, relying on iterative multi-step actions, typically multiply API calls severalfold.
 While enabling complex task handling and improved robustness, optimizing action planning to reduce redundant interactions across diverse database scenarios remains an important direction for future research.
-As reported in the related study [68], a trade-off between performance and computational efficiency should be considered carefully, and designing a comparable (even better) in-context learning method with less API cost will be a practical implementation and is still under exploration.
-Compared to PLM-based methods, the inference speed of LLM-based methods is observably slower [26, 19].
+As reported in the related study [zhang2023actsql], a trade-off between performance and computational efficiency should be considered carefully, and designing a comparable (even better) in-context learning method with less API cost will be a practical implementation and is still under exploration.
+Compared to PLM-based methods, the inference speed of LLM-based methods is observably slower [yang2024harnessing, li2023resdsql].
 Accelerating inference by shortening the input length and reducing the number of stages in implementation would be intuitive for the in-context learning paradigm.
-For local LLMs, from a starting point [82], more speedup strategies can be studied in enhancing the model’s architecture in future exploration.
+For local LLMs, from a starting point [kou2024cllms], more speedup strategies can be studied in enhancing the model’s architecture in future exploration.
 VI-C Data Privacy and Interpretability
-As a part of the LLMs’ study, LLM-based text-to-SQL also faces some general challenges present in LLM research [141, 142, 143].
+As a part of the LLMs’ study, LLM-based text-to-SQL also faces some general challenges present in LLM research [yan2024protecting, zhang2023siren, singh2024rethinking].
 Potential improvements from the text-to-SQL perspective are also expected to be seen in these challenges, thereby extensively benefiting the study of LLMs.
 As previously discussed in Sec.
-IV-A, the in-context learning paradigm predominates the number and performance in recent studies, with the majority of work using proprietary models for implementation [7, 8].
+IV-A, the in-context learning paradigm predominates the number and performance in recent studies, with the majority of work using proprietary models for implementation [pourreza2023dinsql, gao2023dailsql].
 A straightforward challenge is proposed regarding data privacy, as calling proprietary APIs to handle local databases with confidentiality can pose a risk of data leakage.
 Using a local fine-tuning paradigm can partially address this issue.
-Still, the current performance of vanilla fine-tuning is not ideal [8], and advanced fine-tuning framework potentially relies on proprietary LLMs for data augmentation [24].
+Still, the current performance of vanilla fine-tuning is not ideal [gao2023dailsql], and advanced fine-tuning framework potentially relies on proprietary LLMs for data augmentation [li2024codes].
 Based on the current status, more tailored frameworks in the local fine-tuning paradigm for text-to-SQL deserve widespread attention.
-Overall, the development of deep learning continually faces challenges regarding interpretability [143, 144].
-As a long-standing challenge, considerable work has already been studied to address this issue [145, 146].
+Overall, the development of deep learning continually faces challenges regarding interpretability [singh2024rethinking, dai2022knowledge].
+As a long-standing challenge, considerable work has already been studied to address this issue [zhang2024comprehensive, meng2023massediting].
 However, in text-to-SQL research, the interpretability of LLM-based implementation is still not being discussed, whether in the in-context learning or fine-tuning paradigm.
-The approaches with a decomposition phase explain the text-to-SQL implementation process from the perspective of step-by-step generation [7, 47].
-Building on this, combining advanced study in interpretability [147, 148] to enhance text-to-SQL interpreting the model architecture from the database knowledge aspect remains a future direction.
+The approaches with a decomposition phase explain the text-to-SQL implementation process from the perspective of step-by-step generation [pourreza2023dinsql, tai2023exploring].
+Building on this, combining advanced study in interpretability [meng2022locating, zheng2023edit] to enhance text-to-SQL interpreting the model architecture from the database knowledge aspect remains a future direction.
 VI-D Extensions
-As a sub-field of LLMs and natural language understanding research, many studies in these fields have been adopted for text-to-SQL tasks, advancing its development [149, 126].
+As a sub-field of LLMs and natural language understanding research, many studies in these fields have been adopted for text-to-SQL tasks, advancing its development [wei2022chain, wang2023selfconsistency].
 However, text-to-SQL research can also be extended to the larger scope studies of these fields at meanwhile.
 For instance, SQL generation is a part of code generation.
-The well-designed approaches in code generation also obtain promising performance in text-to-SQL [79, 80], performing generalization across various programming languages.
+The well-designed approaches in code generation also obtain promising performance in text-to-SQL [ni2023lever, chen2024selfdebugging], performing generalization across various programming languages.
 The potential extension of some tailored text-to-SQL frameworks to NL-to-code studies can also be discussed.
-For instance, frameworks integrating execution output in NL-to-code can also achieve solid performance in SQL generation [7].
-An attempt to extend execution-aware approaches in text-to-SQL with other advancing modules [70, 28] for code generation is worth discussing.
+For instance, frameworks integrating execution output in NL-to-code can also achieve solid performance in SQL generation [pourreza2023dinsql].
+An attempt to extend execution-aware approaches in text-to-SQL with other advancing modules [hong2024knowledgetosql, dong2023c3] for code generation is worth discussing.
 From another perspective, we previously discussed that text-to-SQL can enhance LLM-based question-answering (QA) by providing factual information.
-The database can store relational knowledge as structural information, and the structure-based QA can potentially benefit from text-to-SQL (e.g., knowledge-based question-answering, KBQA [150, 151]).
-Construct the factual knowledge with database structure, and then incorporate the text-to-SQL system to achieve information retrieval, which can potentially assist further QA with more accurate factual knowledge [152].
+The database can store relational knowledge as structural information, and the structure-based QA can potentially benefit from text-to-SQL (e.g., knowledge-based question-answering, KBQA [luo2023chatkbqa, li2024flexkbqa]).
+Construct the factual knowledge with database structure, and then incorporate the text-to-SQL system to achieve information retrieval, which can potentially assist further QA with more accurate factual knowledge [xiong2024interactive].
 More extensions of text-to-SQL studies are expected in future work.
+[Uncaptioned image] Zijin Hong is a first-year Ph.D.
+student in the DEEP Lab, Department of Computing, The Hong Kong Polytechnic University, under the supervision of Prof.
+Xiao Huang.
+He holds a dual bachelor’s degree in Applied Mathematics with Mathematics from Jinan University and the University of Birmingham.
+His research interests include machine learning, large language models, and their applications, particularly LLM-based text-to-SQL.
+He has published papers and served as a reviewer for venues such as ACL, EMNLP, NAACL, AAAI, ICLR, and IEEE TKDE.
+[Uncaptioned image] Zheng Yuan is a second-year Ph.D.
+student at the Department of Computing, The Hong Kong Polytechnic University.
+He received a bachelor degree in Computer Science and Technology from Shenzhen University in 2021.
+He is currently a group member in the DEEP Lab supervised by Prof.
+Xiao Huang at the Hong Kong Polytechnic University.
+His research interests focus on large language models and their applications in data-driven scenarios, including retrieval-augmented generation, text-to-SQL, and recommender systems.
+[Uncaptioned image] Qinggang Zhang is currently a Postdoctoral Fellow at the Deep Lab, Department of Computing, The Hong Kong Polytechnic University, Hong Kong SAR.
+He received his Ph.D.
+and B.Eng.
+degrees from The Hong Kong Polytechnic University and Northwestern Polytechnical University, respectively.
+His research interests include KGs, LLMs, RAG, and Text-to-SQL.
+He has published over 20 papers while serving as a reviewer for NeurIPS, ICML, ICLR, KDD, IEEE TKDE, and IEEE TPAMI.
+[Uncaptioned image] Hao Chen is an assistant professor at the Faculty of Data Science, City University of Macau.
+Before that, he worked as a postdoctoral research fellow at The Hong Kong Polytechnic University.
+He received his Ph.D degree from Beihang University, China in 2022.
+He has published over 40 papers, such as TKDE, NeurIPS, ACL, SIGKDD, WWW, and SIGIR, etc.
+He received the Best Paper Award Honorable Mention at SIGIR 2023.
+His research interests include recommender systems, graph neural networks, and large language models.
+[Uncaptioned image] Junnan Dong is the final-year Ph.D.
+candidate in the department of computing, Hong Kong Polytechnic University.
+He focuses on Large Language Models, Knowledge Graphs and Graph Retrieval Augmented Generation for advanced and knowledgeable domain-specific reasoning.
+He has published more than ten papers on top-tier conferences, including NeurIPS, ACL, WWW, KDD, SIGIR and WSDM.
+He has been continuously serving as a program committee member for NeurIPS, ICML, ICLR, ACL and WWW.
+[Uncaptioned image] Feiran Huang received the Ph.D.
+degree from the School of Computer Science and Engineering, Beihang University, Beijing, China, in 2018.
+He is currently a professor with the College of Information Science and Technology, Jinan University.
+He has over 100 publications appearing in top conferences and journals, such as SIGIR, KDD, WWW, MM, TIP, IJCV, and TKDE.
+He received 5 best paper awards, such as the Best Paper Award Honourable Mention in SIGIR 2023 and Best Paper Award Runner Up in PAKDD 2023.
+He holds over 50 US, Chinese, and international granted patents.
+He is an editorial board member of IEEE Transactions on Affective Computing, ACM Transactions on Recommender Systems, etc.
+His research interests include recommender systems, social networks, sentiment analysis, and large language models.
+[Uncaptioned image] Xiao Huang is an Assistant Professor in the Department of Computing at The Hong Kong Polytechnic University.
+He earned his Ph.D.
+in Computer Engineering from Texas A&M University in 2020, an M.S.
+in Electrical Engineering from the Illinois Institute of Technology in 2015, and a B.S.
+in Engineering from Shanghai Jiao Tong University in 2012.
+His scholarly contributions are highly regarded within the academic community, amassing over 4,500 citations.
+He received the Best Paper Award Honorable Mention at SIGIR 2023.
+He has successfully led or completed seven research projects as Principal Investigator.

### v8

@@
 Methods that fall into multiple categories will be introduced accordingly.
-The dataset column consists of six text-to-SQL datasets, which can be clarified by names: Spider [yu2018Spider], BIRD [li2023BIRD], Spider-DK [gan2021Spider-DK] Spider-SYN [gan2021Spider-SYN], Spider-Realistic [deng2021Spider-Realistic], and Spider 2.0 [lei2025spider2]
-.
+The dataset column consists of six text-to-SQL datasets, which can be clarified by names: Spider [yu2018Spider], BIRD [li2023BIRD], Spider-DK [gan2021Spider-DK] Spider-SYN [gan2021Spider-SYN], Spider-Realistic [deng2021Spider-Realistic], and Spider 2.0 [lei2025spider2].
 We use the reference ID in the table due to column space limitations.
@@
 CoT method is introduced in multiple venues: NeurIPS’23 [li2023BIRD], EMNLP’23 [tai2023exploring], VLDB’24 [gao2023dailsql], arXiv’24 [zhang2024sgusql].
-Methods Applied LLMs Dataset Metrics $\mathbf{C}_{1}$ $\mathbf{C}_{2}$ $\mathbf{C}_{3}$ $\mathbf{C}_{4}$ Release Time Publication Venue MBR-Exec [shi2022natural] Codex [yu2018Spider] EX Apr-2022 EMNLP’22 Coder-Reviewer [zhang2022coder] Codex [yu2018Spider] EX Nov-2022 ICML’23 LEVER [ni2023lever] Codex [yu2018Spider] EX Feb-2023 ICML’23 Self-Debugging [chen2024selfdebugging] StarCoder* [yu2018Spider] EX Apr-2023 ICLR’24 DESEM+P [guo2023desem] ChatGPT [yu2018Spider, gan2021Spider-SYN] EX Apr-2023 PRICAI’23 DIN-SQL [pourreza2023dinsql] GPT-4* [li2023BIRD, yu2018Spider, lei2025spider2] EX, EM, VES Apr-2023 NeurIPS’23 CoT [tai2023exploring, li2023BIRD, gao2023dailsql, zhang2024sgusql] GPT-4 [yu2018Spider, deng2021Spider-Realistic, li2023BIRD] EX, VES May-2023 Multiple Venues StructGPT [jiang2023structgpt] ChatGPT* [yu2018Spider, gan2021Spider-SYN, deng2021Spider-Realistic] EX May-2023 EMNLP’23 SD+SA+Voting [nan2023enhancing] ChatGPT* [yu2018Spider, gan2021Spider-SYN, deng2021Spider-Realistic] EX May-2023 EMNLP’23 Findings QDecomp [tai2023exploring] Codex [yu2018Spider, deng2021Spider-Realistic] EX May-2023 EMNLP’23 Least-to-Most [tai2023exploring] Codex [yu2018Spider] EX May-2023 EMNLP’23 SQL-PaLM [sun2023sqlpalm] PaLM-2 [yu2018Spider] EX May-2023 TMLR’24 RAG+SP&DRC [guo2023ragsql] ChatGPT [yu2018Spider] EX Jul-2023 ICONIP’23 C3 [dong2023c3] ChatGPT [yu2018Spider] EX Jul-2023 arXiv’23 DAIL-SQL [gao2023dailsql] GPT-4* [yu2018Spider, li2023BIRD, deng2021Spider-Realistic, lei2025spider2] EX, EM, VES Aug-2023 VLDB’24 ODIS [chang2023selective] Codex* [yu2018Spider] EX Oct-2023 EMNLP’23 Findings ACT-SQL [zhang2023actsql] GPT-4* [yu2018Spider, gan2021Spider-SYN] EX, EM Oct-2023 EMNLP’23 Findings MAC-SQL [wang2024macsql] GPT-4* [yu2018Spider, li2023BIRD] EX, EM, VES Dec-2023 COLING’25 DEA-SQL [xie2024deasql] GPT-4 [yu2018Spider] EX Feb-2024 ACL’24 Findings FUSED [wang2024fused] ChatGPT* [yu2018Spider] EX Feb-2024 EMNLP’24 Findings DELLM [hong2024knowledgetosql] GPT-4* [yu2018Spider, li2023BIRD] EX, VES Feb-2024 ACL’24 Findings SGU-SQL [zhang2024sgusql] GPT-4* [yu2018Spider, li2023BIRD] EX, EM Feb-2024 ICML’25 PoT [xia2024sqlcraft] GPT-4* [yu2018Spider, li2023BIRD] EX Feb-2024 arXiv’24 SQL-CRAFT [xia2024sqlcraft] GPT-4* [yu2018Spider, li2023BIRD] EX Feb-2024 arXiv’24 FUXI [gu2024fuxi] GPT-4* [li2023BIRD] EX Feb-2024 EMNLP’24 MetaSQL [fan2024metasql] GPT-4* [yu2018Spider] EX, EM Feb-2024 ICDE’24 $\textit{{R}}^{3}$ [xia2024r3] GPT-4* [yu2018Spider, li2023BIRD] EX Feb-2024 ACL’25 Workshop PET-SQL [li2024petsql] GPT-4 [yu2018Spider] EX Mar-2024 arXiv’24 PURPLE [ren2024purple] GPT-4* [yu2018Spider, gan2021Spider-SYN, deng2021Spider-Realistic] EX, EM Mar-2024 ICDE’24 Dubo-SQL [thorpe2024dubo] ChatGPT* [li2023BIRD] EX Apr-2024 arXiv’24 TA-SQL [qu2024generation] GPT-4* [yu2018Spider, li2023BIRD, gan2021Spider-DK, deng2021Spider-Realistic] EX Apr-2024 ACL’24 Findings MCS-SQL [lee2025mcs] GPT-4 [yu2018Spider, li2023BIRD] EX, VES May-2024 COLING’25 CHESS [talaei2024chess] GPT-4o* [yu2018Spider, li2023BIRD, lei2025spider2] EX, VES May-2024 arXiv’24 SuperSQL [li2024dawn] GPT-4* [yu2018Spider, li2023BIRD] EX, EM, VES Jun-2024 VLDB’24 Distillery [maamari2024the] GPT-4o-Mini* [li2023BIRD] EX Aug-2024 NeurIP’24 Workshop E-SQL [caferouglu2024esql] GPT-4o* [yu2018Spider, li2023BIRD] EX, VES Sep-2024 arXiv’24 CHASE-SQL [pourreza2025chasesql] Gemini-1.5* [yu2018Spider, li2023BIRD] EX, VES Oct-2024 ICLR’25 RSL-SQL [cao2024rsl] DeepSeek-V3* [yu2018Spider, li2023BIRD] EX, VES Nov-2024 arXiv’24 Spider-Agent [lei2025spider2] o1-preview* [lei2025spider2] EX Nov-2024 ICLR’25 ICRL [toteja2025context] ChatGPT [yu2018Spider] EX Jan-2025 COLING’25 Gen-SQL [shi2025gen] LLaMA-3* [yu2018Spider, li2023BIRD] EX, EM, VES Jan-2025 COLING’25 SAFE-SQL [lee2025safe] GPT-4o-Mini* [yu2018Spider] EX, EM Feb-2025 arXiv’25 ReFoRCE [deng2025reforce] o1-preview* [lei2025spider2] EX Feb-2025 ICLR’25 Workshop LinkAlign [wang2025linkalign] DeepSeek-R1* [li2023BIRD, lei2025spider2] EX, EM Mar-2025 arXiv’25
+Methods Applied LLMs Dataset Metrics $\mathbf{C}_{1}$ $\mathbf{C}_{2}$ $\mathbf{C}_{3}$ $\mathbf{C}_{4}$ Release Time Publication Venue
+MBR-Exec [shi2022natural] Codex [yu2018Spider] EX Apr-2022 EMNLP’22
+Coder-Reviewer [zhang2022coder] Codex [yu2018Spider] EX Nov-2022 ICML’23
+LEVER [ni2023lever] Codex [yu2018Spider] EX Feb-2023 ICML’23
+Self-Debugging [chen2024selfdebugging] StarCoder* [yu2018Spider] EX Apr-2023 ICLR’24
+DESEM+P [guo2023desem] ChatGPT [yu2018Spider, gan2021Spider-SYN] EX Apr-2023 PRICAI’23
+DIN-SQL [pourreza2023dinsql] GPT-4* [li2023BIRD, yu2018Spider, lei2025spider2] EX, EM, VES Apr-2023 NeurIPS’23
+CoT [tai2023exploring, li2023BIRD, gao2023dailsql, zhang2024sgusql] GPT-4 [yu2018Spider, deng2021Spider-Realistic, li2023BIRD] EX, VES May-2023 Multiple Venues
+StructGPT [jiang2023structgpt] ChatGPT* [yu2018Spider, gan2021Spider-SYN, deng2021Spider-Realistic] EX May-2023 EMNLP’23
+SD+SA+Voting [nan2023enhancing] ChatGPT* [yu2018Spider, gan2021Spider-SYN, deng2021Spider-Realistic] EX May-2023 EMNLP’23 Findings
+QDecomp [tai2023exploring] Codex [yu2018Spider, deng2021Spider-Realistic] EX May-2023 EMNLP’23
+Least-to-Most [tai2023exploring] Codex [yu2018Spider] EX May-2023 EMNLP’23
+SQL-PaLM [sun2023sqlpalm] PaLM-2 [yu2018Spider] EX May-2023 TMLR’24
+RAG+SP&DRC [guo2023ragsql] ChatGPT [yu2018Spider] EX Jul-2023 ICONIP’23
+C3 [dong2023c3] ChatGPT [yu2018Spider] EX Jul-2023 arXiv’23
+DAIL-SQL [gao2023dailsql] GPT-4* [yu2018Spider, li2023BIRD, deng2021Spider-Realistic, lei2025spider2] EX, EM, VES Aug-2023 VLDB’24
+ODIS [chang2023selective] Codex* [yu2018Spider] EX Oct-2023 EMNLP’23 Findings
+ACT-SQL [zhang2023actsql] GPT-4* [yu2018Spider, gan2021Spider-SYN] EX, EM Oct-2023 EMNLP’23 Findings
+MAC-SQL [wang2024macsql] GPT-4* [yu2018Spider, li2023BIRD] EX, EM, VES Dec-2023 COLING’25
+DEA-SQL [xie2024deasql] GPT-4 [yu2018Spider] EX Feb-2024 ACL’24 Findings
+FUSED [wang2024fused] ChatGPT* [yu2018Spider] EX Feb-2024 EMNLP’24 Findings
+DELLM [hong2024knowledgetosql] GPT-4* [yu2018Spider, li2023BIRD] EX, VES Feb-2024 ACL’24 Findings
+SGU-SQL [zhang2024sgusql] GPT-4* [yu2018Spider, li2023BIRD] EX, EM Feb-2024 ICML’25
+PoT [xia2024sqlcraft] GPT-4* [yu2018Spider, li2023BIRD] EX Feb-2024 arXiv’24
+SQL-CRAFT [xia2024sqlcraft] GPT-4* [yu2018Spider, li2023BIRD] EX Feb-2024 arXiv’24
+FUXI [gu2024fuxi] GPT-4* [li2023BIRD] EX Feb-2024 EMNLP’24
+MetaSQL [fan2024metasql] GPT-4* [yu2018Spider] EX, EM Feb-2024 ICDE’24
+$\textit{{R}}^{3}$ [xia2024r3] GPT-4* [yu2018Spider, li2023BIRD] EX Feb-2024 ACL’25 Workshop
+PET-SQL [li2024petsql] GPT-4 [yu2018Spider] EX Mar-2024 arXiv’24
+PURPLE [ren2024purple] GPT-4* [yu2018Spider, gan2021Spider-SYN, deng2021Spider-Realistic] EX, EM Mar-2024 ICDE’24
+Dubo-SQL [thorpe2024dubo] ChatGPT* [li2023BIRD] EX Apr-2024 arXiv’24
+TA-SQL [qu2024generation] GPT-4* [yu2018Spider, li2023BIRD, gan2021Spider-DK, deng2021Spider-Realistic] EX Apr-2024 ACL’24 Findings
+MCS-SQL [lee2025mcs] GPT-4 [yu2018Spider, li2023BIRD] EX, VES May-2024 COLING’25
+CHESS [talaei2024chess] GPT-4o* [yu2018Spider, li2023BIRD, lei2025spider2] EX, VES May-2024 arXiv’24
+SuperSQL [li2024dawn] GPT-4* [yu2018Spider, li2023BIRD] EX, EM, VES Jun-2024 VLDB’24
+Distillery [maamari2024the] GPT-4o-Mini* [li2023BIRD] EX Aug-2024 NeurIP’24 Workshop
+E-SQL [caferouglu2024esql] GPT-4o* [yu2018Spider, li2023BIRD] EX, VES Sep-2024 arXiv’24
+CHASE-SQL [pourreza2025chasesql] Gemini-1.5* [yu2018Spider, li2023BIRD] EX, VES Oct-2024 ICLR’25
+RSL-SQL [cao2024rsl] DeepSeek-V3* [yu2018Spider, li2023BIRD] EX, VES Nov-2024 arXiv’24
+Spider-Agent [lei2025spider2] o1-preview* [lei2025spider2] EX Nov-2024 ICLR’25
+ICRL [toteja2025context] ChatGPT [yu2018Spider] EX Jan-2025 COLING’25
+Gen-SQL [shi2025gen] LLaMA-3* [yu2018Spider, li2023BIRD] EX, EM, VES Jan-2025 COLING’25
+SAFE-SQL [lee2025safe] GPT-4o-Mini* [yu2018Spider] EX, EM Feb-2025 arXiv’25
+ReFoRCE [deng2025reforce] o1-preview* [lei2025spider2] EX Feb-2025 ICLR’25 Workshop
+LinkAlign [wang2025linkalign] DeepSeek-R1* [li2023BIRD, lei2025spider2] EX, EM Mar-2025 arXiv’25
 1.